### v1

# A solution of the problem of standard compact Clifford-Klein forms

Abstract
We solve the long standing problem of classification of standard compact Clifford-Klein forms of homogeneous spaces of simple non-compact real Lie groups. The result is that standard compact Clifford-Klein forms always arise from triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ of real Lie algebras such that $\mathfrak{h}\subset\mathfrak{g},\mathfrak{l}\subset\mathfrak{g}$, $\mathfrak{g}$ is simple and absolutely simple, $\mathfrak{h},\mathfrak{l}$ are (non-compact) reductive, $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$, and the intersection $\mathfrak{h}\cap\mathfrak{l}$ is compact. The consequence of this is the following characterization of proper co-compact actions of reductive Lie subgroups $L\subset G$ on a homogeneous spaces $G/H$ determined by absolutely simple real Lie group $G$ and a closed reductive subgroup $H$: $L$ acts on $G/H$ properly and co-compactly if and only if $G=H\cdot L$ and $H\cap L$ is compact.
Key words and phrases: proper action, homogeneous space, semisimple Lie algebra. 2020 Mathematics Subject Classification: 17B05, 17B10, 53C30, 57S25, 57S30.
1 Introduction
Let $G$ be a connected real semisimple linear Lie group and $H\subset G$ a closed connected reductive subgroup.
Definition 1 (T. Kobayashi, [19],[24]).
A homogeneous space $G/H$ admits a standard compact Clifford-Klein form, if there exists a closed reductive subgroup $L\subset G$ such that $L$ acts properly and co-compactly on $G/H.$
This notion was first introduced in [24]. A list of such spaces was given in [23] and [25]. The term "standard" was first used in [19]. Standard Clifford-Klein forms are determined by triples $(G,H,L)$, where $G$ is a semisimple real linear Lie group, and $H$ and $L$ are reductive subgroups. The general theory [24] shows that we can restrict ourselves to considering semisimple Lie subgroups $H,L$ (see Section 2.1 for details), and we will do that throughout this article.
Theorem 1.
Assume that $\mathfrak{g}$ is absolutely simple, and $\mathfrak{h}$ and $\mathfrak{l}$ are reductive. All triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ listed in Table 1 yield standard compact Clifford-Klein forms.
Standard Clifford-Klein forms
$\mathfrak{g}$ $\mathfrak{h}$ $\mathfrak{l}$
$\mathfrak{su}(2,2n)$ $\mathfrak{sp}(1,n)$ $\mathfrak{su}(1,2n)$
$\mathfrak{su}(2,2n)$ $\mathfrak{sp}(1,n)$ $\mathfrak{u}(1,2n)$
$\mathfrak{so}(2,2n)$ $\mathfrak{so}(1,2n)$ $\mathfrak{su}(1,n)$
$\mathfrak{so}(2,2n)$ $\mathfrak{so}(1,2n)$ $\mathfrak{u}(1,n)$
$\mathfrak{so}(4,4n)$ $\mathfrak{so}(3,4n)$ $\mathfrak{sp}(1,n)$
$\mathfrak{so}(4,4n)$ $\mathfrak{so}(3,4n)$ $\mathfrak{sp}(1,n)\times\mathfrak{sp}(1)$
$\mathfrak{so}(4,4n)$ $\mathfrak{so}(3,4n)$ $\mathfrak{sp}(1,n)\times\mathfrak{so}(2)$
$\mathfrak{so}(4,4n)$ $\mathfrak{so}(3,4n)$ $\mathfrak{sp}(1,n)$
$\mathfrak{so}(3,4)$ $\mathfrak{g}_{2(2)}$ $\mathfrak{so}(1,4)\times\mathfrak{so}(2)$
$\mathfrak{so}(3,4)$ $\mathfrak{g}_{2(2)}$ $\mathfrak{so}(1,4)$
$\mathfrak{so}(8,8)$ $\mathfrak{so}(7,8)$ $\mathfrak{so}(1,8)$
$\mathfrak{so}(4,4)$ $\mathfrak{so}(3,4)$ $\mathfrak{so}(1,4)\times\mathfrak{so}(3)$
$\mathfrak{so}(4,4)$ $\mathfrak{so}(3,4)$ $\mathfrak{so}(1,4)\times\mathfrak{so}(2)$
$\mathfrak{so}(4,4)$ $\mathfrak{so}(3,4)$ $\mathfrak{so}(1,4)$
Table 1:
The proof of this result is well known: all triples in Table 1 are Lie algebra decompositions in the sense of Onishchik: $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$ [29]. It follows from this work, that $G=H\cdot L$ as well, and that $L$ acts transitively on $G/H$. One obtains a natural diffeomorphism $G/H\cong L/L\cap H$, and so if $H\cap L$ is compact, $L$ acts properly and co-compactly on $G/H$. It is important to note that these are the only known examples of compact standard Clifford-Klein forms (with $\mathfrak{g}$ simple and $\mathfrak{h}$ proper and non-compact).
In this paper we prove that this Table contains all non-trivial triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ with semisimple $\mathfrak{h}$ and $\mathfrak{l}$ generating standard compact Clifford-Klein forms.
Theorem 2.
Assume that $\mathfrak{g}$ is absolutely simple and $\mathfrak{h},$ $\mathfrak{l}$ are proper semisimple subalgebras of non-compact type. The only Lie algebra triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ which yield compact standard Clifford-Klein forms are contained in Table 1.
The discussion in Section 2 shows that Theorem 2 yields a complete classification of all standard compact Clifford-Klein forms determined by triples $(G,H,L)$ with $G$ simple and $H,L$ reductive. Note that if $\mathfrak{g}$ is simple exceptional then, by [4], for any proper reductive subalgebra $\mathfrak{h}\subset\mathfrak{g}$ of non-compact type the corresponding homogeneous space does not admit compact standard Clifford-Klein forms. Therefore, in the proof of Theorem 2 we may assume that $\mathfrak{g}$ is classical simple.
It is also important to stress that the proof of Theorem 2 yields an interesting characterization of proper co-compact actions of reductive Lie groups on homogeneous spaces $G/H$ determined by simple absolutely simple Lie groups $G$ and (non-compact) reductive subgroups.
Theorem 3.
Let $G/H$ be a homogeneous space determined by a simple absolutely simple Lie group $G$ and a reductive (non-compact) closed subgroup $H$. A reductive Lie subgroup $L\subset G$ acts on $G/H$ properlyand co-compactly, if and only if $G=H\cdot L$ and $H\cap L$ is compact.
There are various reasons for studying standard compact Clifford-Klein forms. One of them comes from spectral analysis on pseudo-Riemannian manifolds [19], [20]. In this area, one is interested in constructing quotients $M_{\Gamma}=\Gamma\backslash M$ of discrete groups of isometries of a pseudo-Riemannian homogeneous manifold $M=G/H$. However, it may happen that $M$ does not admit infinite discontinuous groups of isometries (the Calabi-Markus phenomenon). On the other hand, there exist examples of manifolds $G/H$ admitting large discontinuous groups such that $M_{\Gamma}$ is compact, or of finite volume. In this case, we will call $M_{\Gamma}$ a compact Clifford-Klein form, or a Clifford-Klein form of finite volume. Thus, it is a challenging problem to understand when do such compact quotients exist. Suppose a Lie subgroup $L\subset G$ acts properly on $G/H$. Then the action of any discrete subgroup $\Gamma$ of $L$ is automatically properly discontinuous, and this action is free whenever $\Gamma$ is torsion-free. Moreover, if $\Gamma$ is co-compact in $L$, and $L$ acts co-compactly on $G/H$, then the quotient $\Gamma\backslash G/H$ is compact. A quotient $\Gamma\backslash G/H$ is called standard if $\Gamma$ is contained in a reductive subgroup $L$ acting properly on $G/H$. Thus, the first question which arises is to find standard quotients, which is equivalent to finding triples $(G,H,L)$ as in Definition 1. Partial results in this direction were obtained in [4] and [32].
One of the most challenging problems in this area of research is a problem of T. Kobayashi: does any homogeneous space $G/H$ admitting compact Clifford-Klein form admits also a standard one? Clearly, one of the ways to attack it is to understand all standard Clifford-Klein forms.
In a more geometric vein, if $\Gamma\subset G$ acts freely and co-compactly on a homogeneous space $G/H$ with non-compact $G$ and $H$, any $G$-invariant geometric structure descends onto the quotient, which yields compact manifolds with such structures. The topology of such manifolds can be understood, and therefore, these manifolds can serve as testing examples. For instance, we mention [3].
In the context of this work, one should mention that there are many non-existence results for compact Clifford-Klein forms (not necessarily standard). For example, there are no standard Clifford-Klein forms of exceptional Lie groups [4]. Topological obstructions to compact Clifford-Klein forms were found in [27],[28], [33] together with examples. A procedure of constructing homogeneous spaces with no compact Clifford-Klein forms was obtained in [5]. On the positive side, all compact standard Clifford-Klein of symmetric spaces were classified in [32]. The existence of non-standard compact Clifford-Klein forms is a deep mathematical problem of a different nature. We refer to a very recent work [26]. Acknowledgement. We thank T. N. Venkataramana for valuable discussions during the meeting "Zariski dense subgroups, number theory and geometric applications" held at ICTS, Bangalore in January 2024, and for pointing out an interpretation of our results described by Theorem 3.
2 Notation and summary of [6]
2.1 Properties of Lie algebra embeddings yielding compact Clifford-Klein forms
Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple, which means that $\mathfrak{g}$ is absolutely simple, $\mathfrak{h},\mathfrak{l}$ are reductive subalgebras of $\mathfrak{g}$ and $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ induces a compact standard Clifford-Klein form. Throughout we work with root systems of complex semisimple Lie algebras and with real root systems of their real forms. Also, we use some relations between them, which we reproduce here in order to fix notation (see [30], Section 4.1). We may assume that there exists a Cartan involution $\theta$ of $\mathfrak{g}$ such that $\theta(\mathfrak{h})=\mathfrak{h},$ $\theta(\mathfrak{l})=\mathfrak{l}.$ Hence the restriction of $\theta$ is a Cartan involution of $\mathfrak{h}$ and $\mathfrak{l}.$ We obtain the following Cartan decompositions
$\mathfrak{g}=\mathfrak{k}+\mathfrak{p},\ \mathfrak{h}=\mathfrak{k}_{h}+\mathfrak{p}_{h},\ \mathfrak{l}=\mathfrak{k}_{l}+\mathfrak{p}_{l},\ \ \mathfrak{k}_{h},\mathfrak{k}_{l}\subset\mathfrak{k},\ \mathfrak{p}_{h},\mathfrak{p}_{l}\subset\mathfrak{p}.$
Denote by $\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c}$ the complexifications of $\mathfrak{g},\mathfrak{h},\mathfrak{l},$ respectively. Let $\mathfrak{a}\subset\mathfrak{p}$ be a maximal abelian subspace of $\mathfrak{p}$ and denote by $\mathfrak{m}_{0}$ the centralizer of $\mathfrak{a}$ in $\mathfrak{k}.$ Choose a Cartan subalgebra $\mathfrak{t}$ of $\mathfrak{m}_{0}.$ Then $\mathfrak{j}:=\mathfrak{t}+\mathfrak{a}$ is a Cartan subalgebra of $\mathfrak{g}$ and the complexification $\mathfrak{j}^{c}$ of $\mathfrak{j}$ is a Cartan subalgebra of $\mathfrak{g}^{c}.$ Let $\Delta$ be a root system of $\mathfrak{g}^{c}$ with respect to $\mathfrak{j}^{c}.$ Consider the root decomposition
$\mathfrak{g}^{c}=\mathfrak{j}^{c}+\sum_{\alpha_{c}\in\Delta}\mathfrak{g}_{\alpha_{c}}.$
The Lie algebra $\mathfrak{j}:=i\mathfrak{t}+\mathfrak{a}$ is a real form of $\mathfrak{j}^{c}$ with
$\mathfrak{j}=\{A\in\mathfrak{j}^{c}\ |\ \alpha_{c}(A)\in\mathbb{R}\ \textrm{for any}\ \alpha_{c}\in\Delta\}.$
The weights of the adjoint representation of $\mathfrak{a}$ constitute a root system $\Sigma\subset\mathfrak{a}^{*}$ (not necessarily reduced). There is a natural map $pr:\Delta\rightarrow\Sigma$ given by $\alpha_{c}\rightarrow\alpha_{c}|_{\mathfrak{a}}$. It is easy to see that $pr(\Delta)=\Sigma\cup\{0\}.$ Therefore, we will call $pr(\alpha_{c})=\alpha_{c}|_{\mathfrak{a}}\in\mathfrak{a}^{*}$ (if non-zero) the restricted roots. Thus, $\Sigma=\{\alpha:=\alpha_{c}|_{\mathfrak{a}}\ |\ \alpha_{c}\in\Delta\}\setminus\{0\},$
$\mathfrak{g}_{\alpha}:=\sum_{\begin{subarray}{c}\alpha_{c}\in\Delta\\ \alpha_{c}|_{\mathfrak{a}}=\alpha\end{subarray}}\mathfrak{g}_{\alpha_{c}}\cap\ \mathfrak{g}.$
The reader should keep in mind, that we use the following convention. Roots of complex Lie algebra $\mathfrak{g}^{c}$ are denoted by $\alpha_{c},\beta_{c},,...$, the corresponding restrictions are $\alpha,\beta,...$ and so on. Vectors from the root spaces $\mathfrak{g}_{\alpha_{c}}$ of $\mathfrak{g}^{c}$ and $\mathfrak{g}_{\alpha}$ from $\mathfrak{g}$ are always denoted by $x_{\alpha},x_{\beta},...$. The meaning of these will be always clear from the context. However, to avoid clumsy notation, in several cases we write $\alpha_{i}$ for complex roots. In each case we state explicitly, which root (complex or its restriction) is used.
We get the following decomposition of $\mathfrak{g}$
$\mathfrak{g}=\mathfrak{m}_{0}+\mathfrak{a}+\sum_{\alpha\in\Sigma}\mathfrak{g}_{\alpha}.$
Let $B$ be the Killing form of $\mathfrak{g}.$ It extends onto $\mathfrak{g}^{c}$ as a Killing form of $\mathfrak{g}^{c}$ and we denote it by the same letter. Consider the compact real form $\mathfrak{g}_{u}:=\mathfrak{k}+i\mathfrak{p},$ of $\mathfrak{g}^{c},$ and denote by $\tau$ the complex conjugation in $\mathfrak{g}^{c}$ with respect to $\mathfrak{g}_{u}.$ Also we denote by $\sigma$ the complex conjugation of $\mathfrak{g}^{c}$ with respect to $\mathfrak{g}$. It is well known ([31], p. 228) that there is a Hermitian non-degenerate form on $\mathfrak{g}^{c}$ given by the formula $B_{\tau}(X,Y)=-B(X,\tau(Y)).$ Choose a set of positive restricted roots $\Sigma^{+}\subset\Sigma.$ There is a system of positive roots $\Delta^{+}\subset\Delta$ such that
$\Sigma^{+}=\{\alpha_{c}|_{\mathfrak{a}}\ |\ \alpha_{c}\in\Delta^{+}\}-\{0\}.$
The subalgebra $\mathfrak{n}:=\sum_{\alpha\in\Sigma^{+}}\mathfrak{g}_{\alpha}$ is a nilpotent subalgebra of $\mathfrak{g}.$ The decomposition
$\mathfrak{g}=\mathfrak{k}+\mathfrak{a}+\mathfrak{n}$
is the Iwasawa decomposition of $\mathfrak{g}.$ Notice that $\mathfrak{a}+\mathfrak{n}$ is a solvable subalgebra and $\mathfrak{a}$ normalizes $\mathfrak{n}.$ On the Lie group level we have $G=KAN,$ where $K$ is a maximal compact subgroup of $G$ and $N$ is a maximal unipotent subgroup of $G.$ Analogously we obtain the Iwasawa decompositions for $\mathfrak{h}$ and $\mathfrak{l}$
$\mathfrak{h}=\mathfrak{k}_{h}+\mathfrak{a}_{h}+\mathfrak{n}_{h},\ \ H=K_{h}A_{h}N_{h},$
$\mathfrak{l}=\mathfrak{k}_{l}+\mathfrak{a}_{l}+\mathfrak{n}_{l},\ \ L=K_{l}A_{l}N_{l}.$
In the same way we get a nilpotent Lie subalgebra of $\mathfrak{g}$ given by the negative restricted roots
$\mathfrak{n}^{-}=\sum_{\alpha\in\Sigma^{-}}\mathfrak{g}_{\alpha},$
as well as $\mathfrak{n}_{h}^{-},\mathfrak{n}_{l}^{-}$ for $\mathfrak{h}$ and $\mathfrak{l}.$ Since $\theta(\mathfrak{g}_{\alpha})=\mathfrak{g}_{-\alpha}$ for all $\alpha\in\Sigma$ ([17], Proposition 5.9) we get $\theta(\mathfrak{n})=\mathfrak{n}^{-}$.
Lemma 1 ([6]).
Under the assumptions of Theorem 4, we may assume that
$\mathfrak{n}=\mathfrak{n}_{h}\oplus\mathfrak{n}_{l},\ \ \mathfrak{n}^{-}=\mathfrak{n}_{h}^{-}\oplus\mathfrak{n}_{l}^{-},\ \ \mathfrak{a}=\mathfrak{a}_{h}\oplus\mathfrak{a}_{l},$
$\theta(\mathfrak{h})=\mathfrak{h},\ \ \theta(\mathfrak{l})=\mathfrak{l}.$
We will denote by $\Sigma_{h}$, $\Sigma_{l}$ the real root systems of $\mathfrak{h}$ and $\mathfrak{l}$, respectively, corresponding to the above decompositions.
Definition 2.
Embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g},\mathfrak{l}\hookrightarrow\mathfrak{g}$ fulfilling the conditions of Lemma 1 will be called good, and the corresponding subalgebras will be called well embedded.
In this terminology we may formulate the problem of classifying standard compact Clifford-Klein forms as a problem of describing Lie algebra triples given by good embeddings. We will say that triples of Lie algebras $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ and $(\mathfrak{g},\mathfrak{h}^{\prime},\mathfrak{l}^{\prime})$ are isomorphic, if there exist inner automorphisms $\varphi_{1},\varphi_{2}\in\operatorname{Int}(\mathfrak{g})$ such that $\mathfrak{h}^{\prime}=\varphi_{1}(\mathfrak{h})$ and $\mathfrak{l}^{\prime}=\varphi_{2}(\mathfrak{l})$.
Proposition 1 ([6]).
Let $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ be a triple of Lie algebras which determines a standard compact semisimple Clifford-Klein form. Then there is an isomorphic Lie algebra triple $(\mathfrak{g},\mathfrak{h}^{\prime},\mathfrak{l}^{\prime})$ such that $\mathfrak{h}^{\prime}$ and $\mathfrak{l}^{\prime}$ are well embedded.
Assume that $0\not=X\in\mathfrak{t}$. Define
$\Delta_{m}=\{\alpha_{c}\in\Delta\,|\,\alpha_{c}|_{\mathfrak{a}}=0\},$
$\Delta_{m}^{+}=\Delta^{+}\setminus\Delta_{m},\Delta_{m}^{-}=\Delta^{-}\setminus\Delta_{m},$
$\Delta_{p}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)>0\},\,\Delta_{n}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)<0\},$
$\Delta_{p}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)>0\},\,\Delta_{n}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)<0\},$
$\Delta_{0}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)=0\},\ \Delta_{0}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)=0\},$
$\Delta_{0}=\Delta_{0}^{+}\cup\Delta_{0}^{-}.$
Theorem 4 ([6]).
Let
$Z=\sum_{\alpha_{c}\in\Delta_{p}^{+}\cup\Delta_{n}^{+}}\mathfrak{g}_{\alpha_{c}}\subset\mathfrak{n}^{c}.$
Let $\pi:\mathfrak{n}^{c}=\sum_{\alpha_{c}\in\Delta_{m}^{+}}\mathfrak{g}_{\alpha_{c}}\rightarrow Z$ be the natural projection, $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$. Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple and that there exists a non-zero $X\in\mathfrak{t}$ such that $X\not\in(\mathfrak{h}+\mathfrak{l})$. We may assume that there exists a non-zero $X\in\mathfrak{t}$ such that $X$ is $B$-orthogonal to $\mathfrak{h}+\mathfrak{l}.$ Then:
  1. 1.
    There exists a basis of $Z_{h}$ of the form
    $S_{h}^{i}=x_{\alpha_{i}}+\sum_{l=1}^{k}a^{i}_{k+l}x_{\alpha_{k+l}},a^{i}_{k+l}\in\mathbb{C},\alpha_{i}\in\Delta_{p}^{+},\alpha_{k+l}\in\Delta_{n}^{+}.$
  2. 2.
    For any $S_{h}^{i}\in Z_{h}$ $\alpha_{i}|_{\mathfrak{a}_{h}}\in\Sigma_{h}$.
  3. 3.
    Each complexified real root space $\mathfrak{h}_{\gamma}^{c}$ is spanned by vectors of the form
    $S_{h}^{i_{1}}+Q_{1},...,S_{h}^{i_{s}}+Q_{s},Q_{s+1},...,Q_{s+w},s+w=\dim\,{h}_{\gamma}^{c},$
    where $\alpha_{i_{1}},...,\alpha_{i_{s}}$ are all roots from $\Delta_{p}^{+}$ whose restrictions onto $\mathfrak{a}_{h}$ coincide with $\gamma$, while all $Q_{j}$ satisfy the conditions
    $Q_{j}\in\sum_{\alpha_{c}\in\Delta_{0},\ \alpha_{c}|_{\mathfrak{a}_{h}}=\gamma}\mathfrak{g}_{\alpha_{c}}.$
Analogous conditions hold for $\mathfrak{l}_{\gamma}^{c}$.
Corollary 1.
If $(\mathfrak{g},\mathfrak{h},\mathfrak{l}),$ with $\mathfrak{g}$ absolutely simple and split, yields a standard compact Clifford-Klein form then $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}.$ Thus $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is contained in Table 1.
We consider triples $(G,H,L)$ which yield Clifford-Klein forms up to an equivalence which is given by the following well-known fact.
Proposition 2.
If a triple $(G,H,L)$ yields a standard compact Clifford-Klein form then so does $(G,g_{1}Hg_{1}^{-1},g_{2}Lg_{2}^{-1})$ for any $g_{1},g_{2}\in G.$
This enables us to choose embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$, $\mathfrak{l}\hookrightarrow\mathfrak{g}$ up to conjugacy in $G$, and we will do this throughout without further notice. It is sufficient to consider the condition $G=HL$ on the Lie algebra level, because of the classical result of Onishchik.
Theorem 5 ([29], Theorem 3.1).
Let $\mathfrak{g}$ be a reductive real Lie algebra and let $\mathfrak{g}^{\prime},$ $\mathfrak{g}^{\prime\prime}$ be two reductive subalgebras. Let $G$ be a connected Lie subgroup with Lie algebra $\mathfrak{g}$ and let $G^{\prime}$ and $G^{\prime\prime}$ be its connected Lie subgroups corresponding to Lie algebras $\mathfrak{g}^{\prime}$ and $\mathfrak{g}^{\prime\prime}$. Then $G=G^{\prime}G^{\prime\prime}$ if and only if $\mathfrak{g}=\mathfrak{g}^{\prime}+\mathfrak{g}^{\prime\prime}.$
Theorem 6 ([2], Corollary 3, see also [21]).
If $G/H$ admits a compact Clifford-Klein form then the center of $H$ is compact.
Let
$d(G):=\dim\,G/K,\,d(H):=\dim\,H/K_{h},\,d(L):=\dim\,L/K_{l}.$
Clearly, we can write $d(\mathfrak{g}),d(\mathfrak{h}),d(\mathfrak{l}).$
Theorem 7 ([24], Theorem 4.7).
Assume that $L$ acts properly on $G/H.$ The triple $(G,H,L)$ induces a standard compact Clifford-Klein form if and only if
$d(G)=d(H)+d(L),\,\text{or, equivalently}\,\,d(\mathfrak{g})=d(\mathfrak{h})+d(\mathfrak{l}).$ (1)
By [2] we may assume that the centers of $H$ and $L$ are compact. Let $\mathfrak{h}^{\prime}\subset\mathfrak{h}$ be a semisimple Lie algebra given by the sum of all simple ideals of $\mathfrak{h}$ which are of non-compact type. Let $H^{\prime}\subset H$ be the closed connected subgroup corresponding to $\mathfrak{h}^{\prime}.$ We see that $L$ acts properly on $G/H^{\prime}$ and that $d(G)=d(H^{\prime})+d(L).$ This way the triple $(G,H^{\prime},L)$ induces a standard compact Clifford-Klein form.
2.2 Chevalley bases compatible with non-compact real forms
Given a root $\alpha_{c}$ and a root space $\mathfrak{g}_{\alpha_{c}}$ we denote by $t_{\alpha}\in\mathfrak{j}^{c}$ the vector determined by the equality $B(t_{\alpha},h)=\alpha_{c}(h)$ for all $h\in\mathfrak{j}^{c}$. Set $h_{\alpha}=2t_{\alpha}/B(t_{\alpha},t_{\alpha}).$
Definition 3.
A Chevalley basis of $\mathfrak{g}^{c}$ with respect to $\mathfrak{j}^{c}$ is a basis of $\mathfrak{g}^{c}$ consisting of $x_{\alpha}\in\mathfrak{g}_{\alpha_{c}}$ and $h_{\alpha}$ with the following properties:
  1. 1.
    $[x_{\alpha},x_{-\alpha}]=-h_{\alpha},\,\forall\alpha_{c}\in\Delta$,
  2. 2.
    for each pair $\alpha_{c},\beta_{c}\in\Delta$ such that $\alpha_{c}+\beta_{c}\in\Delta$ the constants $c_{\alpha,\beta}\in\mathbb{C}$ determined by $[x_{\alpha},x_{\beta}]=c_{\alpha,\beta}x_{\alpha+\beta}$ satisfy $c_{\alpha,\beta}=c_{-\alpha,-\beta}$.
Throughout, we work with a special choice of the Chevalley basis compatible with a non-compact real form $\mathfrak{g}$ of $\mathfrak{g}^{c}$.
Lemma 2.
For the chosen $\mathfrak{j}^{c}$ in $\mathfrak{g}^{c}$ there is a base of $\sum_{\alpha_{c}\in\Delta}\mathfrak{g}_{\alpha_{c}}$ of the form
$\{\tilde{x}_{\alpha}\ |\ \tilde{x}_{\alpha}\in\mathfrak{g}_{\alpha_{c}},\ \alpha_{c}\in\Delta\},$
such that for every $\alpha_{c}\in\Delta$
$\tau(\tilde{x}_{\alpha})=\tilde{x}_{-\alpha},\ \ B(\tilde{x}_{\alpha},\tilde{x}_{-\alpha})=-1.$
Proof.
Note that $\tau(\mathfrak{j}^{c})=\mathfrak{j}^{c}$ and therefore $\tau$ permutes the root spaces. Since $x_{\alpha}+\tau(x_{\alpha})\in\mathfrak{g}_{u}$ we get $B(x_{\alpha}+\tau(x_{\alpha}),x_{\alpha}+\tau(x_{\alpha}))<0$. Note that $B(x_{\alpha},x_{\beta})\not=0$ if and only if $\alpha_{c}=-\beta_{c}$. Hence $B(x_{\alpha},\tau(x_{\alpha}))<0$ and therefore $\tau(x_{\alpha})=ax_{-\alpha}$ for some $a>0$. Since $\tau$ is an involution, $\tau(x_{-\alpha})=\frac{1}{a}x_{\alpha}$. Therefore, one can take $\tilde{x}_{\alpha}=\frac{1}{\sqrt{ab}}x_{\alpha}$ and $\tilde{x}_{-\alpha}=\frac{\sqrt{a}}{\sqrt{b}}x_{-\alpha}$. ∎
Remark 1.
Note that in this subsection we consider only root systems of semisimple complex Lie algebras, and, therefore, follow the standard notation. However, in the subsequent sections we distinguish between real and complex roots.
3 Some properties of root systems
In this subsection we establish some properties of abstract root systems. Our aim is Lemma 5 which describes subsets of the root system consisting of roots simultaneously vanishing on a pair of vectors. This may be known, but we did not find any appropriate reference. Let $\Delta\subset\mathbb{R}^{n}$ be an indecomposable root system in the Euclidean space $(\mathbb{R}^{n},(,))$ ([30], Chapter 3). The system $\Delta$ is not assumed to be reduced. Let $\Delta^{+}\subset\Delta$ be a subset of positive roots and $\Pi\subset\Delta^{+}$ the subset of simple roots. There is a unique maximal root $\beta\in\Delta^{+}$ such that for every $\alpha\in\Delta^{+}$ vector $\beta-\alpha$ is a combination of simple roots with non-negative coefficients.
Lemma 3.
Let $A$ be a vector in the interior of the positive Weyl chamber. Then $\beta(A)>\alpha(A)$ for any $\alpha\in\Delta^{+}-\{\beta\}.$
Proof.
Since $A$ is in the interior of the positive Weyl chamber, $\alpha_{i}(A)>0$ for any simple root $\alpha_{i}\in\Pi.$ Thus for $\alpha\in\Delta^{+},$ $\alpha\neq\beta$ we obtain $(\beta-\alpha)(A)>0.$
Proposition 3 ([30], Chapter 3, Proposition 1.2).
Let $\gamma_{1},\gamma_{2}\in\Delta.$ If $(\gamma_{1},\gamma_{2})<0$ then $\gamma_{1}+\gamma_{2}\in\Delta.$
Lemma 4.
If $\beta\in\Delta^{+}$ is such that for any $\alpha\in\Delta^{+}$ the vector $\beta+\alpha$ is not a root then $\beta$ is the highest root.
Proof.
Assume that $\beta,\gamma\in\Delta^{+}$ are distinct and have the property that $\beta+\alpha$ and $\gamma+\alpha$ are not roots for any $\alpha\in\Delta^{+}.$ By Proposition 3 for any simple root $\alpha_{i}$ we get $(\beta,\alpha_{i}),(\gamma,\alpha_{i})\geq 0.$ As $\beta\neq\gamma,$ for some simple root $\alpha_{i}$ we get $(\beta,\alpha_{i})\neq(\gamma,\alpha_{i})$. So $(\beta+\gamma,\alpha_{i})>0.$ By Proposition 3 (setting $\gamma_{1}=\alpha+\beta,$ $\gamma_{2}=-\alpha_{i}$) we see that $\theta:=\beta+\gamma-\alpha_{i}$ is a root. Since $\beta$ is the highest root, $\gamma$ is a positive root and $\alpha_{i}$ is simple (a lowest root) therefore $\gamma-\alpha_{i}=0$ (otherwise $\beta-\theta$ written as a combination of positive roots would have some negative coefficients). But $\gamma$ is a simple root and so does not have the assumed property that $\gamma+\alpha$ is not a root for every positive $\alpha$ (unless $\Delta=\{\alpha,-\alpha\}$ in which case $\gamma=\beta$). ∎
Lemma 5.
Let $X,H\in\mathbb{R}^{n}$ be non-zero. Define
$C_{X}:=\{\alpha\in\Delta\ |\ \alpha(X)=0\},\ \ C_{H}:=\{\alpha\in\Delta\ |\ \alpha(H)=0\}.$
Then $\Delta\neq C_{X}\cup C_{H}.$
Proof.
Assume that $\Delta=C_{X}\cup C_{H}.$ Consider sets $C_{H}\setminus C_{X}$ and $C_{X}\setminus C_{H}.$ If one of them is empty then all the roots in $\Delta$ are orthogonal to $X$ or to $H$, but this is impossible. Therefore there exist simple roots $\gamma_{1},\gamma_{2}$ (for any choice of a subset of simple roots) such that
$\gamma_{1}\in C_{H}\setminus C_{X},\ \ \gamma_{2}\in C_{X}\setminus C_{H},$
because any root can be represented as a linear combination of simple roots. Since $C_{H}\cap C_{X}$ is a root system, we can take the set of positive roots $C_{HX}^{+}\subset C_{H}\cap C_{X}.$ Let
$C_{X}^{+}:=\{\alpha\in C_{H}\setminus C_{X}\ |\ \alpha(X)>0\},$
$C_{H}^{+}:=\{\alpha\in C_{X}\setminus C_{H}\ |\ \alpha(H)>0\},$
One can see that for $\alpha\in C_{X}^{+}$ and $\beta\in C_{H}^{+}$, vector $\alpha+\beta$ is not a root, because otherwise one would obtain a root which is non-zero on $H$ and on $X.$ Consider the set
$\tilde{\Delta}^{+}:=C_{HX}^{+}\cup C_{X}^{+}\cup C_{H}^{+}.$
Note that for any $\alpha\in\Delta$ exactly one of the roots $\alpha,$ $-\alpha$ is contained in $\tilde{\Delta}^{+}.$ Also if $\alpha,\gamma\in\tilde{\Delta}^{+}$ and $\alpha+\gamma$ is a root, then $\alpha+\gamma\in\tilde{\Delta}^{+}.$ Therefore $\tilde{\Delta}^{+}$ is the set of positive roots from $\Delta$ with the property
$\forall_{\alpha\in\tilde{\Delta}^{+}}\alpha(X)\geq 0,\ \alpha(H)\geq 0.$
Consider the highest root $\tilde{\beta}\in\tilde{\Delta}^{+}.$ Since $\Delta$ is indecomposable $\tilde{\beta}$ is a linear combination of all simple roots of $\tilde{\Delta}^{+}$ with positive coefficients. But we have seen before that there are simple roots $\tilde{\gamma}_{1},$ $\tilde{\gamma}_{2}\in\tilde{\Delta}^{+}$ such that
$\tilde{\gamma}_{1}(X)>0,\ \ \tilde{\gamma}_{2}(H)>0.$
Thus $\tilde{\beta}(X)>0$ and $\tilde{\beta}(H)>0.$ A contradiction. ∎
Proposition 4 ([6]).
Assume that $X\in\mathfrak{t}$ satisfies the assumptions of Theorem 4 and put $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$. Then
$Z=Z_{h}\oplus[X,Z_{h}]=Z_{l}\oplus[X,Z_{l}]=Z_{h}\oplus Z_{l}=Z_{h}\oplus[X,Z_{l}].$
Lemma 6 ([6]).
Under the assumptions of Theorem 4, The involution $\sigma$ permutes the root spaces given by $\Delta^{+}_{m}$ so that if $\alpha_{c}\in\Delta_{p}^{+}$ then $\sigma(\mathfrak{g}_{\alpha_{c}})\subset\mathfrak{g}_{\alpha_{c}^{{}^{\prime}}}$ for some $\alpha_{c}^{{}^{\prime}}\in\Delta_{n}^{+}$ (by abuse of notation we write $\sigma(\Delta_{p})=\Delta_{n}$). Thus Analogous properties hold for $\Delta_{p}^{-},\Delta_{n}^{-}.$
Lemma 7 ([6]).
None of the non-zero root vectors $x_{\beta}$ for $\beta_{c}\in\Delta_{0}-\Delta_{m}$ centralizes $Z+\tau(Z).$
Proof.
Assume there exists $x_{\beta},\beta_{c}\in\Delta_{0}$ such that $[x_{\beta},Z+\tau(Z)]=0$. Take $H_{\beta}\in i\mathfrak{t}+\mathfrak{a}$, defined, as usual by the condition $B(\tilde{H},H_{\beta})=\beta(\tilde{H})$ for any $\tilde{H}\in i\mathfrak{t}+\mathfrak{a}$. Denote by $\mathfrak{g}^{x_{\beta}}$ the centralizer of $x_{\beta}$ in $\mathfrak{g}^{c}$. By [11], Lemma 3.4.3
$\mathfrak{g}^{x_{\beta}}=\oplus_{j\geq 0}(\mathfrak{g}^{c}\cap\mathfrak{g}_{j}),$
where $\mathfrak{g}_{j}$ denote the eigenspaces of $\operatorname{ad}\,H_{\beta}$ (which are integers $j\geq 0$). Write $H_{\beta}=iT+H$, where $T\in\mathfrak{t},H\in\mathfrak{a}$. Note that $H\not=0$, since $\beta\in\Delta_{0}$. We have
$H=\frac{1}{2}(H_{\beta}+\sigma(H_{\beta})).$
Define
$C_{X}=\{\alpha_{c}\in\Delta\,|\alpha_{c}(X)=0\},\,C_{H}=\{\alpha_{c}\in\Delta\,\,|\,\alpha_{c}(H)=0\}.$
Consider the decomposition
$\Delta=\Delta_{m}\cup\Delta_{0}\cup\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$
Note that for any $\mu_{c}\in\Delta_{m}$, $\mu_{c}(H)=0$. Assume that
$\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}\subset C_{H}.$ (2)
Since $\Delta_{0}\subset C_{X}$, and all roots from $\Delta_{m}$ vanish on $H,$ equality (2) would imply
$\Delta=C_{X}\cup C_{H},$
but this contradicts Lemma 5. Thus there exists $\alpha_{c},-\alpha_{c}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha_{c}(H)\not=0$. We have one of the following possibilities
  1. 1.
    $\alpha_{c}(iT)\neq-\alpha_{c}(H).$ In this case (since one can choose between $\alpha_{c}$ and $-\alpha_{c}$) we may assume that $\alpha_{c}(H_{\beta})<0.$
  2. 2.
    $\alpha_{c}(iT)=-\alpha_{c}(H).$ In this case we see that $\sigma(\alpha_{c})(H_{\beta})=\alpha_{c}(-iT+H)=2\alpha_{c}(H)\neq 0$ and by Lemma 6, $\sigma(\alpha_{c})\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ In this case (since one can choose between $\sigma(\alpha_{c})$ and $-\sigma(\alpha_{c})$) we may assume that $\sigma(\alpha_{c})(H_{\beta})<0.$
Therefore in any case there exists $\alpha_{c}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha_{c}(H_{\beta})<0.$ But for any root vector $x_{\alpha}\in Z+\tau(Z)$ we have $\operatorname{ad}_{H_{\beta}}(x_{\alpha})=\alpha_{c}(H_{\beta})x_{\alpha}$ and so for $x_{\alpha}\in\mathfrak{g}^{x_{\beta}},$ $j=\alpha_{c}(H_{\beta})$. This is a contradiction with the inequality $j\geq 0$. The proof of the Lemma is complete. ∎
4 Proof of Theorem 2: eliminating the regular case
We have the following setting
  • $\mathfrak{g}$ is an absolutely simple Lie algebra
  • $\mathfrak{h},\mathfrak{l}$ are semisimple (and consist of simple ideals of non-compact type) subalgebras of $\mathfrak{g}$ s.t. the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ gives a standard compact Clifford-Klein form
  • The triple is not contained in Table 1 and thus there exists $0\neq X\in\mathfrak{m}_{0}$ orthogonal to $\mathfrak{h}\oplus\mathfrak{l}$.
  • $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c},$ that is, some Cartan subalgebra of $\mathfrak{g}^{c}$ normalizes $\mathfrak{h}^{c}.$
Denote by $\mathfrak{d}^{c}$ the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}^{c}.$ We have
$\mathfrak{d}^{c}=\mathfrak{m}_{0}^{c}\oplus(\mathfrak{a}_{h}^{\perp})^{c}\oplus(\Sigma_{\alpha\in\Delta_{0}^{\pm},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha})\oplus\mathfrak{a}_{h}^{c},$
where
$\mathfrak{m}_{0}^{c}=\mathfrak{t}^{c}\oplus\Sigma_{\alpha\in\Delta_{m}}\mathfrak{g}_{\alpha}.$
Notice that $\mathfrak{d}^{c}$ is a reductive subalgebra. Also if $\alpha\in\Delta_{0}^{\pm}$ and $\alpha|_{\mathfrak{a}_{h}}=0$ then (by Theorem 4) necessarily $\alpha(iX)=0.$ Consider the semisimple subalgebra
$\mathfrak{d}_{s}^{c}:=[\mathfrak{d}^{c},\mathfrak{d}^{c}].$
Lemma 8.
The subalgebra $\mathfrak{d}^{c}_{s}$ is not simple or $\Phi:=\{\alpha\in\Delta_{0}^{\pm}\ |\ \alpha|_{\mathfrak{a}_{h}}=0\}=\emptyset.$
Proof.
Assume that $\Phi\neq\emptyset.$ If $\alpha\in\Delta_{m}$ is s.t. $\alpha(iX)\neq 0$ and $\beta\in\Phi$ then $\alpha\pm\beta$ is not a root since otherwise $\alpha\pm\beta\in\Phi$ and $(\alpha\pm\beta)(iX)\neq 0,$ which is impossible. Take any $\alpha_{1}\in\Phi.$ It follows that for any $\alpha\in\Delta_{m},$ $\alpha(iX)\neq 0$ we have $\alpha\perp\alpha_{1}.$ But now every root of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ is orthogonal to $iX$ or to $\alpha_{1}$ and so all roots are contained in two hyperplanes. Lemma 5 implies that $\mathfrak{d}_{s}^{c}$ is not simple. ∎
Corollary 2.
We can decompose $\mathfrak{d}^{c}$ into a sum of three ideals $\mathfrak{d}^{c}=\mathfrak{d}_{1}\oplus\mathfrak{d}_{2}\oplus\mathfrak{a}_{h}^{c}$ such that $iX\in\mathfrak{d}_{1}$ and $\mathfrak{a}_{h}^{\perp},\ (\Sigma_{\alpha\in\Delta_{0}^{\pm},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha})\subset\mathfrak{d}_{2}.$
Proof.
First notice that $\mathfrak{a}_{h}^{c}$ is in the center of $\mathfrak{d}^{c}.$ If $\Phi=\emptyset$ then take $\mathfrak{d}_{1}:=\mathfrak{m}_{0}^{c},$ $\mathfrak{d}_{2}:=(\mathfrak{a}_{h}^{\perp})^{c}.$
Assume that $\alpha_{1}\in\Phi\neq\emptyset.$ All roots of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ lie on two hypersurfaces in $(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c},$ $s_{iX}$ (orthogonal to $iX$) and $s_{\alpha_{1}}$ (orthogonal to $\alpha_{1}$). We see that if $\tilde{\mathfrak{d}}$ is a simple ideal of $\mathfrak{d}_{s}^{c}$ then (by Lemma 5) all roots of $(\tilde{\mathfrak{d}},\tilde{\mathfrak{d}}\cap(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c}))$ lie either in $s_{iX},$ or in $s_{\alpha_{1}}.$ Therefore we may take $\mathfrak{d}_{2}$ the be the sum of all simple ideals which roots lie in $s_{iX}$ plus $(\mathfrak{a}_{h}^{\perp})^{c}.$ Then $\mathfrak{d}_{1}$ is the ideal of $\mathfrak{m}_{0}^{c}$ complementary to $\mathfrak{d}_{2}\cap\mathfrak{m}_{0}^{c}.$
As $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c}$ then some Cartan subalgebra $\mathfrak{c}$ of $\mathfrak{g}^{c}$ normalizes $\mathfrak{h}^{c}.$ We may assume that $\mathfrak{a}_{h}^{c}\subset\mathfrak{c}.$ Indeed, notice that $\mathfrak{h}^{c}\cap\mathfrak{c}$ is a Cartan subalgebra of $\mathfrak{h}^{c}.$ Since any two Cartan subalgebras of $\mathfrak{h}^{c}$ are conjugate by an element of $H^{c}$ thus any Cartan subalgebra of $\mathfrak{h}^{c}$ is of the form $Ad_{h}(\mathfrak{c})\cap\mathfrak{h}^{c}$ for some $h\in H^{c}$ (where $H^{c}$ denotes a connected subgroup of $G^{c}$ corresponding to $\mathfrak{h}^{c}$). Obviously $Ad_{h}(\mathfrak{c})$ normalizes $\mathfrak{h}^{c}.$ As $\mathfrak{a}_{h}^{c}$ is in some Cartan subalgebra of $\mathfrak{h}^{c}$ therefore we may assume that
$\mathfrak{a}_{h}^{c}\subset\mathfrak{c}\subset\mathfrak{d}^{c}.$
Lemma 9.
We may assume that
$\sigma(\mathfrak{c})=\mathfrak{c},\ \ \theta(\mathfrak{c})=\mathfrak{c},\ \ \sigma(\mathfrak{d}_{2})=\mathfrak{d}_{2},\ \ \theta(\mathfrak{d}_{2})=\mathfrak{d}_{2}.$
Proof.
First notice that $\sigma(\mathfrak{a}_{h})=\mathfrak{a}_{h}$ and so $\sigma(\mathfrak{d}^{c})=\mathfrak{d}^{c}.$ Denote by $\mathfrak{d}^{\sigma}$ the real form of $\mathfrak{d}^{c}$ given by $\sigma.$ We see that $\mathfrak{d}^{\sigma}$ is the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}.$ Put $\mathfrak{c}_{h}:=\mathfrak{c}\cap\mathfrak{h}^{c}.$ After conjugating $\mathfrak{c}$ by an element of the centralizer of $\mathfrak{a}_{h}$ in $H^{c}$ we have $\sigma(\mathfrak{c}_{h})=\mathfrak{c}_{h}.$ Therefore $\mathfrak{c}_{h}^{\sigma}$ (the real form of $\mathfrak{c}_{h}$ w.r.t. $\sigma$) is in $\mathfrak{d}^{\sigma}.$ We take a $\sigma$-stable Cartan subalgebra $\mathfrak{c}^{\prime}$ of the normalizer of $\mathfrak{h}^{c}$ in $\mathfrak{d}^{c}$ containing $\mathfrak{c}_{h}^{\sigma}.$ We have $[\mathfrak{c}^{\prime},\mathfrak{h}^{c}]\subset\mathfrak{h}^{c}$ and $\sigma(\mathfrak{c}^{\prime})\subset\mathfrak{c}^{\prime},$ as desired. By the similar argument (by conjugating $(\mathfrak{c}^{\prime})^{\sigma}$ inside $G$) we have $\theta(\mathfrak{c}^{\prime})=\mathfrak{c}^{\prime}.$
Notice that as $\sigma(X)=\theta(X)=X$ we have $\sigma(\mathfrak{d}_{2})=\theta(\mathfrak{d}_{2})=\mathfrak{d}_{2}.$
Lemma 10.
We have $\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}=\{0\}$ and so $\Phi\neq\emptyset.$
Proof.
Let $Y\in\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}.$ For every $1\leq i\leq k$ we have
$[Y,S_{h}^{i}+Q_{i}+\sigma(S_{h}^{i}+Q_{i})]\subset\mathfrak{n}_{h}\cap\mathfrak{n}_{h}^{\perp}=\{0\},$
$[Y,i(S_{h}^{i}+Q_{i}-\sigma(S_{h}^{i}+Q_{i}))]\subset\mathfrak{n}_{h}\cap\mathfrak{n}_{h}^{\perp}=\{0\},$
where $\mathfrak{n}_{h}^{\perp}$ is the orthogonal complement of $\mathfrak{n}_{h}$ in $\mathfrak{n}$ w.r.t. $B_{\theta}(\bullet,\bullet)=B(\bullet,\theta(\bullet)).$ It follows that $Y$ centralizes all roots of $\Delta_{p}^{\pm}$ and $\Delta_{n}^{\pm}.$ By Lemma 7 every root of $\Delta_{0}^{\pm}$ is of the form $\alpha_{1}+\alpha_{2}$ for some $\alpha_{1},\alpha_{2}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ But now $Y$ centralizes $\mathfrak{n}$ and so $Y=0.$
Notice that
$\mathfrak{d}^{\sigma}\subset\mathfrak{k}_{d}\oplus\mathfrak{m}_{0}\oplus\mathfrak{p}_{d}\oplus\mathfrak{a}_{h},$
$\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma},$
$\mathfrak{p}_{d}\subset\mathfrak{d}_{2}^{\sigma},$
where
$\mathfrak{p}_{d}=\mathfrak{a}_{h}^{\perp}\oplus\textrm{Lin}_{\mathbb{R}}\{x_{\alpha}+\sigma(x_{\alpha})-\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))-\theta(i(x_{\alpha}+\sigma(x_{\alpha})))\ |\ \alpha\in\Phi\},$
$\mathfrak{k}_{d}=\textrm{Lin}_{\mathbb{R}}\{x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))+\theta(i(x_{\alpha}+\sigma(x_{\alpha})))\ |\ \alpha\in\Phi\}.$
Previous Lemmas and the fact that $\textrm{rank}(\mathfrak{g})>\textrm{rank}(\mathfrak{m}_{0}\oplus\mathfrak{a}_{h})$ imply the following corollary.
Corollary 3.
We have
$\mathfrak{c}^{\sigma}\cap(\mathfrak{p}_{d}\oplus\mathfrak{a}_{h})=\mathfrak{a}_{h}.$
Proof.
Obviously $\mathfrak{a}_{h}\subset\mathfrak{c}^{\sigma}.$ Take any $Y\in\mathfrak{p}_{d}\cap\mathfrak{c}_{\sigma}.$ Then $Y$ is conjugate by an element $k^{-1}$ of $K_{d}$ (connected closed subgroup of $K$ corresponding to reductive subalgebra $\mathfrak{k}_{d}$) to $\mathfrak{a}_{h}^{\perp}.$ As $\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma}$ thus it centralizes $iX$ and $\mathfrak{a}_{h}$ and so for any $g\in K_{d}$ $Ad_{g}\mathfrak{h}$ is well embedded (as $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$). Also $\mathfrak{a}_{h}^{\perp}$ is orthogonal to $\textrm{Ad}_{g}\mathfrak{h}.$ We can repeat the argument of the proof of Lemma 10 to $\textrm{Ad}_{k}Y$ and $\textrm{Ad}_{k}\mathfrak{h}.$
Because $\theta(\mathfrak{d}_{2}^{\sigma})=\mathfrak{d}_{2}^{\sigma}$ therefore $\mathfrak{d}_{2}^{\sigma}\subset\mathfrak{k}.$ As every element of a Lie algebra of compact type is in some Cartan subalgebra we may conjugate $\mathfrak{c}^{\sigma}$ by an element of $\mathfrak{d}_{2}^{\sigma}$ so that
$0\neq\tilde{Y}:=x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha}))\in\mathfrak{c}^{\sigma},$
where $\alpha\in\Phi.$ As $\mathfrak{d}_{2}^{\sigma}$ centralizes $iX$ and $\mathfrak{a}_{h}$ thus for any $g\in D_{2}^{\sigma}$ (connected closed subgroup of $G$ corresponding to reductive subalgebra $\mathfrak{d}_{2}^{\sigma}$) $Ad_{g}\mathfrak{h}$ is well embedded (as $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$). Thus Theorem 4 holds for $\textrm{Ad}_{g}\mathfrak{h}.$
Lemma 11.
We have $\tilde{Y}\notin\mathfrak{h}.$
Proof.
Since $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}$ therefore there exist $H\in\mathfrak{n}_{h}$ and $L\in\mathfrak{n}_{l}$ s.t. $x_{\alpha}+\sigma(x_{\alpha})=H+L$ and so $\tilde{Y}=H+\theta(H)+L+\theta(L).$ If $\tilde{Y}\in\mathfrak{h}$ then $L+\theta(L)\in\mathfrak{h}.$ But
$\{W+\theta(W)\ |\ W\in\mathfrak{n}\}=\{W+\theta(W)\ |\ W\in\mathfrak{n}_{h}\}\oplus\{W+\theta(W)\ |\ W\in\mathfrak{n}_{l}\}$
and thus $L+\theta(L)=H_{1}+\theta(H_{1})$ for some $H_{1}\in\mathfrak{n}_{h}.$ As $\mathfrak{n}\cap\theta(\mathfrak{n})=\{0\}$ therefore $L\in\mathfrak{n}_{h}$ and so $L=0.$ Thus $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}_{h}$ and $\alpha|_{\mathfrak{a}_{h}}=0,$ a contradiction. ∎
It follows that there exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$ which is orthogonal (with respect to $B$) to $\mathfrak{h}.$ Since $\mathfrak{h}^{c}$ is regular we obtain the following
Corollary 4.
There exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$ which centralizes $\mathfrak{h}.$
Theorem 8.
Assume that neither $G/H$ nor $H$ is compact. In the setting of this section (there exists $iX\neq 0$ and $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c}$) the pair $(\mathfrak{g},\mathfrak{h})$ cannot induce standard Clifford-Klein form.
Proof.
Put
$Z_{p}:=\sum_{\alpha\in\Delta_{p}^{\pm}}\mathfrak{g}_{\alpha},\ \ Z_{n}:=\sum_{\alpha\in\Delta_{n}^{\pm}}\mathfrak{g}_{\alpha},\ \ O:=\sum_{\alpha\in\Delta_{0}^{\pm}}\mathfrak{g}_{\alpha}.$
As $[Y,iX]=[Y,\mathfrak{a}_{h}]=0$ therefore
$[Y,Z_{n}]\subset Z_{n},\ \ [Y,Z_{p}]\subset Z_{p},\ \ [Y,O]\subset O.$
But $Y$ centralizes $\mathfrak{h}^{c}$ and so for $1\leq i\leq k$ we have $[Y,S_{h}^{i}+O_{i}]=0.$ By Theorem 4 (1.) we have
$[Y,x_{\alpha_{i}}]=0,\ \ \alpha_{i}\in\Delta_{p}^{\pm}$
and thus (as $[Y,\sigma(x_{\alpha_{i}})]=0$) $[Y,x_{\alpha_{j}}]=0$ for any $\alpha_{j}\in\Delta_{n}^{\pm}.$ By Lemma 7 this implies $[Y,x_{\beta_{i}}]=0$ for every $\beta_{i}\in\Delta_{0}^{\pm}.$ Thus
$[Y,\mathfrak{n}]=[Y,\theta(\mathfrak{n})]=0$
and so $[Y,\mathfrak{a}]=0$ as $\mathfrak{a}\subset[\mathfrak{n},\theta(\mathfrak{n})].$ But now $[Y,\mathfrak{p}]=0$ and so $[Y,\mathfrak{k}]=0$ since $\mathfrak{k}=[\mathfrak{p},\mathfrak{p}].$ Therefore $[Y,\mathfrak{g}]=0$ and $Y\neq 0.$ But now $\mathfrak{g}$ cannot be simple, a contradiction. ∎
5 Completion of proof of Theorem 2: non-regular case
In view of Theorem 8 we assume that the embeddings $\mathfrak{h}^{c},\mathfrak{l}^{c}\hookrightarrow\mathfrak{g}^{c}$ are not regular. The proof is divided into two parts:
  1. 1.
    $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple,
  2. 2.
    at least one of $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ is semisimple but not simple.
The proof of the main theorem will follow from Propositions 6, 8, 9 proved in this section. We begin with case 1) and consider embeddings of classical Lie algebras $\mathfrak{h}^{c}$ as linear representations, and, following Dynkin [14], divide these into two classes: irreducible and reducible. In what follows we use the notation from the paper [14]. We use tables of inclusions of simple classical Lie algebras from [14], however, we also reproduce parts of them as Tables 3,4,5 in Section 6.
5.1 Proof in the case when $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple
5.1.1 Irreducible representations and simple classical subalgebras
We begin with the classification of maximal simple classical subalgebras in simple Lie algebras (see, [30], Chapter 6, Theorem 3.1). Following Dynkin [14] we introduce the following terminology. Let $\mathfrak{g}^{c}$ be a classical Lie algebra, that is $\mathfrak{g}^{c}=\mathfrak{sl}(V),\mathfrak{so}(V),\mathfrak{sp}(V)$ where $V$ is a finite-dimensional complex linear space. Note that by assumption, $\mathfrak{h}^{c}$ is also classical, so, $\mathfrak{h}^{c}=\mathfrak{sl}(U),\mathfrak{so}(U),\mathfrak{sp}(U)$ for some finite-dimensional complex linear space.
Definition 4.
We say that $\mathfrak{h}^{c}$ is an irreducible/reducible subalgebra in $\mathfrak{g}^{c}$ if the linear representation $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is irreducible/reducible.
Theorem 9 (Dynkin, [30]).
The classification of maximal semisimple Lie subalgebras $\mathfrak{h}^{c}$ in simple complex Lie algebras $\mathfrak{g}^{c}$ of classical type falls into the classes:
  1. 1.
    $\mathfrak{h}^{c}$ is reducible, and then a) $\mathfrak{h}^{c}$ is maximal parabolic, b) if $\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C})$; if $\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})$, $0<k<n$;
  2. 2.
    $\mathfrak{h}$ is irreducible and non-simple, and then the following possibilities may occur: $\mathfrak{g}^{c}=\mathfrak{sl}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C}),$ $n=st,s,t\geq 2$, if $\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate either to $\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})$, where $n=st,s,t\geq 3$, $s,t\not=4$, or to $\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})$, where $n=st,s,t\geq 2$, if $\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{sp}(t,\mathbb{C})$, where $n=st,s\geq 2,t\geq 3$, or $s=2,t=4$.
  3. 3.
    $\mathfrak{h}^{c}$ is irreducible and simple.
We begin with Dynkin’s classification of simple complex $\mathfrak{h}^{c}$ such that the embedding $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is an irreducible representation of $\mathfrak{h}^{c}$ in one of the classical Lie algebras $\mathfrak{g}^{c}=\mathfrak{sl}(V)$,$\mathfrak{so}(V),\mathfrak{sp}(V)$ for some complex vector space $V$. It says that essentially, maximal and irreducible classes coincide with several exceptions.
Theorem 10 (Dynkin, [30]).
Let $R:\mathfrak{h}\rightarrow\mathfrak{gl}(V)$ be a nontrivial irreducible representation of a simple complex Lie algebra $\mathfrak{h}$. If there are no $R(\mathfrak{h})$-invariant bilinear form on $V$, then $R(\mathfrak{h})$ is maximal in $\mathfrak{sl}(V)$. If $R(\mathfrak{h})$ preserves some orthogonal or symplectic form, then $R(\mathfrak{h})$ is maximal in $\mathfrak{so}(V)$ or $\mathfrak{sp}(V)$. The only exceptions are given in Table 1 of [14] (or Table 7 of [30]).
Remark 2.
Note that exception means "non-maximal" in this theorem.
Theorem 11 (Dynkin, [14], Theorem 2.3).
The list of all triples of irreducible inclusions $\mathfrak{h}^{c}\subset\tilde{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ of simple Lie algebras $\mathfrak{h}^{c},\tilde{\mathfrak{h}}^{c},\mathfrak{g}^{c}$ is given by Table 5 in [14].
Comparing Theorems 9 and 10 we get the following corollary.
Corollary 5.
Any irreducible simple classical subalgebra $\mathfrak{h}^{c}$ of classical Lie algebra $\mathfrak{g}^{c}$ is either maximal, or is contained in Table 1 of [14] (or, equivalently, in Table 7 of [30]), or in Table 5 in [14]. If the representation of $\mathfrak{h}^{c}$ in $\mathfrak{g}^{c}$ is fundamental, then it is contained in Tables 1 and 2 of [7].
Remark 3.
Reading the formulations in [14] one may get an impression that Table 1 contains all the possibilities for irreducible simple classical subalgebras. However Table 5 [14] is necessary, since Table 1 of [14] and tables of fundamental representations are organized in the way which does not cover several possible triples. This is because, for example, $\mathfrak{sp}(n,\mathbb{C})$ is an irreducible representation of minimal dimension $2n$, which belongs to the table of symplectic irreducible representations, but also $\mathfrak{sp}(n,\mathbb{C})$ is contained in $\mathfrak{sl}(2n,\mathbb{C})$. Such cases are covered by Theorem 11.
5.1.2 Using Dynkin’s classification
We summarize the procedure for classification of standard Clifford-Klein forms based on the previous results as follows.
Proposition 5.
The classification of possible standard Clifford-Klein forms of classical simple Lie algebras amounts to finding these as real forms of the following classes of triples $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c}$, where $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple subalgebras (not necessarily proper) of the following lists of subalgebras:
  1. 1.
    the class of Levi subalgebras of parabolic subalgebras,
  2. 2.
    the class of all maximal irreducible simple subalgebras in classical $\mathfrak{g}^{c}$, or belong to Tables 1 and 5 of [14],
  3. 3.
    $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ is simple, reducible and is contained in one of the following inclusions:
    $\mathfrak{h}^{c}\subset\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k)\subset\mathfrak{so}(n,\mathbb{C}),$
    $\mathfrak{h}^{c}\subset\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})\subset\mathfrak{sp}(n,\mathbb{C}),0<k<n,$
    $\mathfrak{h}^{c}\subset\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{sl}(st,\mathbb{C}),\ s,t\geq 2,$
    $\mathfrak{h}^{c}\subset\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{so}(st,\mathbb{C}),\ n=st,s,t\geq 3,s,t\neq 4,$
    $\mathfrak{h}^{c}\subset\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{so}(st,\mathbb{C}),\ s,t\geq 2,$
    $\mathfrak{h}^{c}\subset\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{sp}(t,\mathbb{C})\subset\mathfrak{sp}(st,\mathbb{C}),\ s\geq 2,t\geq 3,\ or\ s=2,t=4,$
    (and the same inclusions may hold for $\mathfrak{l}^{c}$),
  4. 4.
    a class of simple reducible subalgebras such that they are contained in some simple irreducible subalgebra $\tilde{\mathfrak{h}}^{c}$ of $\mathfrak{g}^{c}$ (with $\tilde{\mathfrak{h}}^{c}$ necessarily satisfying assumptions of 1)).
  5. 5.
    exceptional simple Lie subalgebras with the following simplest representations $\mathfrak{g}_{2}\subset\mathfrak{so}(7,\mathbb{C}),$ $\mathfrak{f}_{4}\subset\mathfrak{so}(26,\mathbb{C}),$ $\mathfrak{e}_{6}\subset\mathfrak{sl}(27,\mathbb{C}),$ $\mathfrak{e}_{7}\subset\mathfrak{sp}(56,\mathbb{C}),$ $\mathfrak{e}_{8}\subset\mathfrak{so}(248,\mathbb{C}).$
5.1.3 Non-generic case $\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1}$
This case is in a sense "non-generic" and we consider it first. Note that it will be used in other considerations as well. We will essentially use the following.
Theorem 12 ([21], Theorem 1.5).
Let $G/H$ be a homogeneous space of reductive type. If $G/H$ has a compact Clifford-Klein form, then there does not exist a closed subgroup $G^{\prime}$ reductive in $G$ satisfying the conditions
  • $\mathfrak{a}_{g^{\prime}}\subseteq\mathfrak{a}_{h}$,
  • $d(G^{\prime})>d(H)$.
Proposition 6.
Let
$\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1}$
where $\mathfrak{h}^{c}$ is a regular simple subalgebra of $B_{n}$, and $B_{n}$ is a subalgebra of all fixed points of an outer automorphism of $D_{n+1}$. The pair $(\mathfrak{g},\mathfrak{h})$ cannot generate a standard compact Clifford-Klein form such that $\mathfrak{g}\neq\mathfrak{h}+\mathfrak{l}$ (for some $\mathfrak{l}\subset\mathfrak{g}$).
Proof.
Choose a Cartan subalgebra $\mathfrak{c}\subset D_{n+1}$ such that $\mathfrak{c}\cap\mathfrak{h}^{c}$ is a Cartan subalgebra of $\mathfrak{h}^{c}$. The relation between roots and root vectors for the embedding $B_{n}\subset D_{n+1}$ is well known. Following [16], p. 507, we write $H_{1},...,H_{n+1}$ for the simple co-roots and the root vectors corresponding to them as $X_{1},...,X_{n+1},$. The opposite root vectors are denoted by $Y_{1},...,Y_{n+1}$. The following vectors determine the embedding of $B_{n}$ into $D_{n+1}$:
$H_{1},...,H_{n-1},H_{n}+H_{n+1},$
$X_{1},...,X_{n-1},X_{n}+X_{n+1},Y_{1},...,Y_{n-1},Y_{n}+Y_{n+1}.$
Since any subalgebra $\mathfrak{h}^{c}$ which is regular in $B_{n}$ is obtained from the extended Dynkin diagram of $B_{n}$ by deleting some vertices, we can write all the possibilities for the simple roots and root vectors of $\mathfrak{h}^{c}$ in terms of the simple roots, or corresponding co-roots and root vectors expressed via the data for the type $D_{n+1}$. The extended Dynkin diagram for $B_{n}$ is given by vertices
$\alpha_{0},\alpha_{1},...,\alpha_{n-1},\bar{\alpha}_{n}:=\frac{1}{2}(\alpha_{n}+\alpha_{n+1})$
where $\alpha_{0}$ is the minimal root of $B_{n}$. Looking at the explicit expressions for maximal roots (for example, in [9]) one obtains the expression for each root of this diagram expressed via roots in of $D_{n+1}$:
$\alpha_{0}=-(\alpha_{1}+2\alpha_{2}+\cdots+2\alpha_{n-1}+(\alpha_{n}+\alpha_{n+1})).$ (3)
Notice that $\alpha_{0}$ is a root of $D_{n+1}.$ Moreover, it is the minimal root of $D_{n+1}$ (see [8], Table IV). We have shown the following: the system of simple roots for any $\mathfrak{h}^{c}$ is obtained by deleting some of the roots from (4) in a way to get a connected diagram. In particular, if $\mathfrak{h}^{c}$ is simple and not regular in $\mathfrak{g}^{c},$ then $\mathfrak{h}^{c}$ is obtained by deleting $\alpha_{0},\alpha_{1},...,\alpha_{t},$ $1\leq t<n,$ because otherwise $\mathfrak{h}^{c}$ would be regular in $D_{n+1}$. Therefore, $\mathfrak{h}^{c}$ is necessarily of the form $B_{p}$. Looking at [30], Table 4, one can see that there is only one possible real form (up to conjugation):
$\mathfrak{h}=\mathfrak{so}(q,2p+1-q).$
Let $\theta$ be a Cartan involution of $\mathfrak{g}$ such that $\theta(\mathfrak{h})=\mathfrak{h}.$ We may assume that $\theta(\mathfrak{c})=\mathfrak{c}.$ It follows from the classification of involutive automorphisms (see Section 1.4, Chapter 4 in [30]) and the labels of the extended Dynkin diagram of $D_{n+1}$ that $\theta=\operatorname{Ad}I_{m,2n+2-m}$ for some $p\in\{2,...,n+1\}$ ([30], Theorem 1.4 (2), Chapter 4). Indeed, the Kac diagram of $\theta$ cannot be given by two $1^{\prime}s$ on the opposite sides of the extended Dynkin diagram of $D_{n+1}$ (because $\theta(\mathfrak{h}^{c})=\mathfrak{h}^{c}$). Therefore,
$\mathfrak{h}=\mathfrak{so}(q,2p+1-q)\subset\mathfrak{g}=\mathfrak{so}(m,2n+2-m),$
where $q<m\leq n+1.$ By Theorem 12 we have $2p+1-q=2n+2-m.$ Indeed, if $2p+1-q<2n+2-m$, one can construct $\mathfrak{h}\subset\mathfrak{g}^{\prime}\subset\mathfrak{g}$ with $\mathfrak{a}_{g^{\prime}}=\mathfrak{a}_{h}$ and $d(\mathfrak{g}^{\prime})>d(\mathfrak{h})$ by increasing $2p+1-q$ but still keeping it less than $2n+2-m$. One can do this analyzing separately all the possible choices of $q$ and $m$ even or odd, and using the formulas for the non-compact dimensions from Table 2. The inequality $2p+1-q>2n+2-m$ is not possible for dimensional reasons. This yields an embedding
$\mathfrak{so}(q,2n+2-m)\subset\mathfrak{so}(m,2n+2-m).$
Since $\theta(\mathfrak{h})\subset\mathfrak{h}$, we get the inclusion of maximal compact subalgebras of $\mathfrak{h}$ and $\mathfrak{g}$:
$\mathfrak{so}(q)\oplus\mathfrak{so}(2n+2-m)\subset\mathfrak{so}(m)\oplus\mathfrak{so}(2n+2-m).$
Therefore there is an obvious inclusion
$\mathfrak{so}(m-q)\oplus\mathfrak{so}(q,2n+2-m)\subset\mathfrak{so}(m,2n+2-m).$
We see that the pair $(\mathfrak{so}(m,2n+2-m),\mathfrak{so}(m-q)\oplus\mathfrak{so}(q,2n+2-m))$ is symmetric and $(\mathfrak{so}(m,2n+2-m),\mathfrak{so}(m-q)\otimes\mathfrak{so}(q,2n+2-m))$ induces a compact Clifford-Klein form if and only if $(\mathfrak{so}(m,2n+2-m),\mathfrak{so}(q,2n+2-m))$ induces a compact Clifford-Klein form. Therefore, [32] yields the desired result.
5.1.4 Completion of proof for $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ simple
Proposition 7.
If $\mathfrak{h}^{c}$ is a simple non-regular subalgebra in a simple Lie algebra $\mathfrak{g}^{c}$ of classical type, then
$\operatorname{rank}\mathfrak{h}^{c}\leq\frac{1}{2}(\operatorname{rank}\mathfrak{g}^{c}+1)$
with the following exceptions:
  1. 1.
    $\mathfrak{h}^{c}:=B_{p-1}\subset D_{p}$
  2. 2.
[…]
  3. 3.
[…]
Proof.
Recall that by Proposition 5 the inclusions of a simple Lie algebra $\mathfrak{h}^{c}$ into classical $\mathfrak{g}^{c}$ fall into classes:
  1. 1.
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$, where $\hat{\mathfrak{h}}^{c}$ is maximal parabolic,
  2. 2.
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$, where
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}=\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C}),$
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}=\mathfrak{sp}(k,\mathbb{C}\oplus\mathfrak{sp}(n-k,\mathbb{C}).$
  3. 3.
    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}$, where $\hat{\mathfrak{h}}^{c}$ is of tensor type,
  4. 4.
    $\mathfrak{h}^{c}$ is irreducible and simple.
We begin the proof with case (4) which requires the longest calculations. Since $\hat{\mathfrak{h}}^{c}$ is simple and irreducible, we consider first the case when the inclusion $\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ is a fundamental representation. Looking at Tables 4,5,6 we calculate the possible ranks of the inclusions, using formulas from Table 4. In any case we obtain the required inequality. If the representation $\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ is not fundamental, we use the Weyl dimension formula [18], Theorem 5.84. The proof of case $(2)$ can be obtained from the following observation. If $\mathfrak{h}\hookrightarrow\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$, then necessarily there is an embedding of $\mathfrak{h}$ to at least one of $\mathfrak{g}_{i}$ ($i=1,2$). Otherwise, the kernels of the projections onto $\mathfrak{g}_{i}$ composed with the embedding would yield non-trivial ideals in $\mathfrak{h}$. Thus, $\mathfrak{h}^{c}\subset\mathfrak{g}_{1}$ or $\mathfrak{h}^{c}\subset\mathfrak{g}_{2}$. In any case, one obtains an embedding of simple Lie algebras with $\mathfrak{g}_{i}$ of lower ranks. One proceeds by induction using $(4)$. This argument also settles case $(1)$. Since $\hat{\mathfrak{h}}^{c}$ is maximal parabolic, its reductive Levi subalgebra has the form
$\mathfrak{f}(\mathfrak{c},[M]),\,M\subset\Pi,$
where $M$ is a subset of $\Pi$, and $[M]$ is a subsystem of the root system $\Delta$. Thus, $\mathfrak{h}^{c}$ is a simple subalgebra contained in a semisimple subalgebra of a lower rank. In the same manner one settlles (3), since $\hat{\mathfrak{h}}^{c}$ is a direct sum of two simple Lie subalgebras of $\mathfrak{sl}(V_{1})\oplus\mathfrak{sl}(V_{2})$ embedded into $\mathfrak{sl}(V_{1})\otimes\mathfrak{sl}(V_{2})$. Thus, the same argument works.
Corollary 6.
If $\mathfrak{h}^{c}$ belongs to any of the inclusions in Proposition 7 then no real form of the pair $(\mathfrak{g}^{c},\mathfrak{h}^{c})$ cannot generate compact standard Clifford-Klein form.
Proof.
The proof follows from the proof of Proposition 6 with small modifications. One can notice that if $\mathfrak{g}^{c}=D_{n}$, or $\mathfrak{g}^{c}=B_{n}$ then the same argument works, this time applied to $\mathfrak{h}^{c}$ and the ambient Lie algebra $\mathfrak{g}^{c}$. Still, one obtains the embeddings
$\mathfrak{so}(q,2p-1-q)\subset\mathfrak{so}(m,2n+1-m),\,\,\text{or}\,\,\mathfrak{so}(q,2p-1-q)\subset\mathfrak{so}(m,2n-m)$
and the same argument using Theorem 12 completes the proof. The case $\mathfrak{g}^{c}=C_{n}$ need not be considered for the dimensional reasons: the smallest possible dimension of a symplectic representation of $D_{p}$ is $\frac{n}{2}$. ∎
Let us now list all simple real forms of non-compact type. By a slight abuse of notation we will treat simple but not absolutely simple real Lie algebras of rank 2n as Lie algebras of type $n.$ For instance $\mathfrak{g}=\mathfrak{sl}(n+1,\mathbb{C})$ is classified as type $A_{n}.$ Notice that in this case $\mathfrak{g}^{c}=\mathfrak{sl}(n+1,\mathbb{C})\oplus\mathfrak{sl}(n+1,\mathbb{C}).$
Simple real forms of non-compact type
$A_{t}$:
  • $\mathfrak{su}(a,t+1-a),$ $a\leq\frac{t+1}{2},$ $d:=d(\mathfrak{g})=2at+2a-2a^{2},$ $r:=\textrm{rank}_{\mathbb{R}}(\mathfrak{g})=a,$
  • $\mathfrak{su}^{\ast}(t+1),$ $\frac{t-1}{2}\in\mathbb{N},$ $d=\frac{1}{2}(t^{2}+t-2),$ $r=\frac{t-1}{2},$
  • $\mathfrak{sl}(t+1,\mathbb{R}),$ $d=\frac{1}{2}t(t+3),$ $r=t,$
  • $\mathfrak{sl}(t+1,\mathbb{C}),$ $d=t^{2}+2t,$ $r=t.$
$B_{t}$:
  • $\mathfrak{so}(a,2t+1-a),$ $a<t+1,$ $d=2at+a-a^{2},$ $r=a,$
  • $\mathfrak{so}(2t+1,\mathbb{C}),$ $d=t(2t+1),$ $r=t.$
$C_{t}$:
  • $\mathfrak{sp}(a,t-a),$ $a\leq\frac{t}{2},$ $d=4at-4a^{2},$ $r=a,$
  • $\mathfrak{sp}(t,\mathbb{R}),$ $d=t^{2}+t,$ $r=t,$
  • $\mathfrak{sp}(t,\mathbb{C}),$ $d=2t^{2}+t,$ $r=t,$
$D_{t}$:
  • $\mathfrak{so}(a,2t-a),$ $a<t+1,$ $d=2at-a^{2},$ $r=a,$
  • $\mathfrak{so}^{\ast}(t),$ $d=t^{2}-t,$ $r=[\frac{t}{2}],$
  • $\mathfrak{so}(2t,\mathbb{C}),$ $d=2t^{2}-t,$ $r=t.$
$G_{2}$ (simplest representation $G_{2}\subset B_{3}$):
  • $\mathfrak{g}_{2(2)},$ $d=8,$ $r=2,$
  • $\mathfrak{g}_{2}(\mathbb{C}),$ $d=14,$ $r=2,$
$F_{4}$ (simplest representation $F_{4}\subset D_{13}$):
  • $\mathfrak{f}_{4(4)},$ $d=28,$ $r=4,$
  • $\mathfrak{f}_{4(-2)},$ $d=16,$ $r=1,$
  • $\mathfrak{f}_{4}(\mathbb{C}),$ $d=52,$ $r=4,$
$E_{6}$ (simplest representation $E_{6}\subset A_{26}$):
  • $\mathfrak{e}_{6(6)},$ $d=42,$ $r=6,$
  • $\mathfrak{e}_{6(2)},$ $d=40,$ $r=4,$
  • $\mathfrak{e}_{6(-14)},$ $d=32,$ $r=2,$
  • $\mathfrak{e}_{6(-26)},$ $d=26,$ $r=2,$
  • $\mathfrak{e}_{6}(\mathbb{C}),$ $d=78,$ $r=6,$
$E_{7}$ (simplest representation $E_{7}\subset C_{56}$):
  • $\mathfrak{e}_{7(7)},$ $d=70,$ $r=7,$
  • $\mathfrak{e}_{7(-5)},$ $d=64,$ $r=4,$
  • $\mathfrak{e}_{7(-25)},$ $d=54,$ $r=3,$
  • $\mathfrak{e}_{7}(\mathbb{C}),$ $d=133,$ $r=7,$
$E_{8}$ (simplest representation $E_{8}\subset D_{124}$):
  • $\mathfrak{e}_{8(8)},$ $d=128,$ $r=8,$
  • $\mathfrak{e}_{8(-24)},$ $d=112,$ $r=4,$
  • $\mathfrak{e}_{8}(\mathbb{C}),$ $d=248,$ $r=8,$
Remarks:
  1. 1.
    For all classical types with $\textrm{rank}=t$ and $\textrm{rank}_{\mathbb{R}}=a$ one of the following three algebras has the highest $d$
    • I$:=\mathfrak{so}(a,2t+1-a),$ $d=2at+a-a^{2},$
    • II$:\mathfrak{sp}(a,t-a),$ $d=4at-4a^{2},$
    • III$:=\mathfrak{sp}(t,\mathbb{C}),$ $d=2t^{2}+t.$
  2. 2.
    For each real form of type $F_{4},E_{6},E_{7},E_{8}$ one can find a real form of type I or II with the same real rank, with greater or equal $d$ and with rank at most half of the rank of the simplest representation. For instance consider type $F_{4}.$
    • $\mathfrak{f}_{4(4)}$ ($d=28,$ $r=4$) has the simplest representation in $D_{13}$ (after complexification). One can substitute $\mathfrak{f}_{4(4)}$ with $\mathfrak{so}(4,9)$ ($d=36,$ $r=4$) which is of the type $B_{6}.$
    • $\mathfrak{f}_{4(-2)}$ ($d=16,$ $r=1$) has the simplest representation in $D_{13}$ (after complexification). One can substitute $\mathfrak{f}_{4(-2)}$ with $\mathfrak{sp}(1,5)$ ($d=20,$ $r=1$) which is of the type $C_{6}.$
    • $\mathfrak{f}_{4}(\mathbb{C})$ ($d=52,$ $r=4$) has the simplest representation in $D_{13}\oplus D_{13}\subset D_{26}$ (after complexification). One can substitute $\mathfrak{f}_{4}(\mathbb{C})$ with $\mathfrak{so}(4,21)$ ($d=83,$ $r=4$) which is of the type $B_{12}.$
  3. 3.
    For $\mathfrak{g}_{2}(\mathbb{C})$ one can find a real form of type I ($\mathfrak{so}(2,7)$) with the same real rank, with greater or equal $d$ and with rank at most half of the rank of the representation into algebra of rank 8 or greater. As the simplest representation of $\mathfrak{g}_{2}(\mathbb{C})$ is $B_{6}.$ Thus in each case one has to verify separately the embeddings of $\mathfrak{g}_{2}(\mathbb{C})$ into lie algebras of rank $6$ and $7$
  4. 4.
    For $\mathfrak{g}_{2(2)}$ one can find a real form of type I ($\mathfrak{so}(2,5)$) with the same real rank, with greater or equal $d$ and with rank at most half of the rank of the representation into algebra of rank 6 or greater. As the simplest representation of $\mathfrak{g}_{2(2)}$ is $B_{3}$ then one should in each case verify the embeddings of $\mathfrak{g}_{2(2)}$ into lie algebras of rank $4$ and $5$ This is because the real rank of $\mathfrak{g}_{2(2)}$ equals 2 and so if $(\mathfrak{g},\mathfrak{g}_{2(2)},\mathfrak{l})$ gives a standard compact Clifford-Klein form then necessarily the real rank of $\mathfrak{g}$ is at lest 3. If the $\textrm{rank}(\mathfrak{g})=3$ then $\mathfrak{g}$ is split and we can use Theorem “split”.
We only have to check the triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ such that $\mathfrak{h}$ and $\mathfrak{l}$ are one of the following
  • absolutely simple non-regular,
  • simple but not absolutely simple,
and $\textrm{rank}_{\mathbb{R}}(\mathfrak{h})=:a>0,$ $\textrm{rank}_{\mathbb{R}}(\mathfrak{l})=:b>0,$ $\textrm{rank}_{\mathbb{R}}(\mathfrak{g})=a+b.$
Proposition 8.
For any classical absolutely simple $\mathfrak{g}$ in every case described above
$D:=d(\mathfrak{g})-d(\mathfrak{h})-d(\mathfrak{l})>0$
and so one cannot obtain compact standard Clifford-Klein forms.
Proof.
By the remarks in the previous Section we only have to check the inequality for types I,II,III (and $G_{2}$ for low ranks).
By $X_{t}$ we will denote an arbitrary type of rank $t.$ By $(i,j),$ $i,j\in\{I,II,III\}$ we denote a pair of subalgebras of type i and j, respectively. We exclude the cases in which $\mathfrak{g}$ is split, becuse of [6].
Type: $A_{2n}$ Case $\mathfrak{g}=\mathfrak{su}(a+b,2n+1-a-b)$ We start with $\mathfrak{h}$ of type $G_{2}.$ First we have to check $\mathfrak{g}_{2(2)}$ in $A_{4}.$ But in this case real rank of $\mathfrak{g}$ is at most 2. Secondly we have $\mathfrak{g}_{2}(\mathbb{C})$ in $A_{6}.$ So necessarily $\mathfrak{g}=\mathfrak{su}(3,4).$ One easily checks that there is no simple $\mathfrak{l}$ of real rank one with $d(\mathfrak{l})>8.$
Clearly, the only proper regular subalgebras in $A_{2n}$ are of type $A_{p},$ $p<2n.$ Note that any non-regular simple Lie subalgebra of $A_{2n}$ is of the rank at most $n,$ by Proposition 7.
If $\mathfrak{h}$ is simple but not absolutely simple then it is obviously of type $X_{p}$ with $p\leq n.$ Thus we have to check Lie subalgebras of type I,II and III for $t=n.$ We have
  • $d(\mathfrak{g})=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab,$ $r=a+b,$
  • $d(I)=2an+a-a^{2},$ $r=a,$
  • $d(II)=4an-4a^{2},$ $r=a,$
  • $d(III)=2a^{2}+a,$ $r=a.$
We also have $a+b\leq n.$
(I,I) $D=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab-2an-a+a^{2}-2bn-b+b^{2}\geq a^{2}+b^{2}+a+b>0,$
(I,II) $D=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab-2an-a+a^{2}-4bn+4b^{2}\geq(a-b)^{2}+a+2b+b^{2}>0,$
(I,III) $D=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab-2an-a+a^{2}-2b^{2}-b\geq a^{2}+2ab+a+b>0,$
(II,II) $D=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab-4an+4a^{2}-4bn+4b^{2}=2(a-b)^{2}+2a+2b>0,$
(II,III) $D=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab-4an+4a^{2}-2b^{2}-b\geq 2a^{2}+2a+b>0,$
(III,III) $D=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab-2a^{2}-a-2b^{2}-b\geq 4ab+a+b>0.$
Type: $A_{2n+1}$
Again, by Proposition 7 we see that if $X_{p}\subset A_{2n+1}$ is classical simple non-regular subalgebra or simple but not absolutely simple subalgebra then $p\leq n+1.$
We begin with subalgebras of type $G_{2}.$
  • If $\mathfrak{h}:=\mathfrak{g}_{2(2)}\subset A_{5}$ then (since the real rank of $\mathfrak{g}$ has to be greater than 2 and we omit the split case) necessarily $\mathfrak{g}=\mathfrak{su}(3,3).$ But $\mathfrak{g}_{2(2)}$ is not a subalgebra of $\mathfrak{g}.$
  • If $\mathfrak{h}:=\mathfrak{g}_{2}(\mathbb{C})\subset A_{7}$ then (since the real rank of $\mathfrak{g}$ has to be greater than 2 and we omit the split case) necessarily $\mathfrak{g}=\mathfrak{su}(3,5),$ $\mathfrak{su}(4,4),$ $\mathfrak{su}^{\ast}(8).$ But the maximal compact subgroup of $\mathfrak{g}_{2}(\mathbb{C})$ is $\mathfrak{g}_{2}.$ The maximal compact subgroups of possible $\mathfrak{g}$ are $\mathfrak{su}(3)\oplus\mathfrak{u}(5),$ $\mathfrak{su}(4)\oplus\mathfrak{u}(4),$ $\mathfrak{sp}(4)$ and so $\mathfrak{g}_{2(2)}$ is not a subgroup of $\mathfrak{g}.$
Case $\mathfrak{g}=\mathfrak{su}^{\ast}(2n+2)$ In this case we have to check Lie subalgebras of type I,II and III for $t=n+1.$ We have
  • $d(\mathfrak{g})=2a^{2}+2b^{2}+4ab+3a+3b,$ $r=a+b=n$
  • $d(I)=2an+3a-a^{2},$ $r=a,$
  • $d(II)=4an-4a^{2}+a,$ $r=a,$
  • $d(III)=2a^{2}+a,$ $r=a.$
We also have $a+b=n.$
(I,I) $D=2a^{2}+2b^{2}+4ab+3a+3b-2a(a+b)+a^{2}-3a-2b(a+b)+b^{2}-3b=a^{2}+b^{2}>0,$
(I,II) $D=2a^{2}+2b^{2}+4ab+3a+3b-2a(a+b)+a^{2}-3a-4b(a+b)+4b^{2}-b=(a-b)^{2}+b^{2}+2b>0,$
(I,III) $D=2a^{2}+2b^{2}+4ab+3a+3b-2a(a+b)+a^{2}-3a-2b^{2}-b=a^{2}+2ab+2b>0,$
(II,II) $D=2a^{2}+2b^{2}+4ab+3a+3b-4a(a+b)+4a^{2}-a-4b(a+b)+4b^{2}-b=2(a-b)^{2}+2(a+b)>0,$
(II,III) $D=2a^{2}+2b^{2}+4ab+3a+3b-4a(a+b)+4a^{2}-a-2b^{2}-b=2a^{2}+2(a+b)>0,$
(III,III) $D=2a^{2}+2b^{2}+4ab+3a+3b-2a^{2}-a-2b^{2}-b=4ab+2(a+b)>0.$
[…]
  • $d(\mathfrak{g})=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b,$
  • $d(I)=2an+3a-a^{2},$ $r=a,$
  • $d(II)=4an-4a^{2}+a,$ $r=a,$
  • $d(III^{\prime})=a^{2}+2a,$ $r=a.$
We have $a+b\leq n+1,$ $n\geq 2.$
(I,I) $D=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b-2an-3a+a^{2}-2bn-3b+b^{2}\geq a(a-1)+b(b-1)>0,$ unless $a=b=1.$ But if $a=b=1$ then $D=4(n-1)>0,$
(I,II) $D=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b-2an-3a+a^{2}-4bn+4b^{2}-b\geq(a-b)(a-b-1)+b^{2}+2b>0,$
(I,III’) $D=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b-2an-3a+a^{2}-b^{2}-2b\geq(a-1)(a+b)+ab+b(b-1)>0,$
(II,II) $D=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b-4an+4a^{2}-a-4bn+4b^{2}-b=2(a-b)^{2}+3(a+b)>0,$
(II,III’) $D=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b-4an+4a^{2}-a-b^{2}-2b=b(b-2)+2a^{2}+3a>0,$
(III’,III’) $D=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b-a^{2}-2a-b^{2}-2b\geq(a+b)(a+b-2)+2ab>0.$
Type: $C_{n}$
In this type we only have to check $\mathfrak{g}=\mathfrak{sp}(a+b,n-a-b).$
We begin with subalgebras of type $G_{2}.$ We can eliminate (in the case $\mathfrak{h}=\mathfrak{g}_{2(2)}$) $C_{4}$ and $C_{5}$ since the real of $\mathfrak{sp}(2,2)$ and $\mathfrak{sp}(2,3)$ equals $2$ If $\mathfrak{h}:=\mathfrak{g}_{2}(\mathbb{C})\subset A_{6}$ then (since the real rank of $\mathfrak{g}$ has to be greater than 2 and we omit the split case) necessarily $\mathfrak{g}=\mathfrak{sp}(3,3).$ But now $\mathfrak{l}\subset\mathfrak{sp}(3,3)$ has to have real rank equal to 1 and $d=22.$ One easily verifies that this is impossible. We can use the same argument for $C_{7}$ (in this case necessarily $\mathfrak{g}=\mathfrak{sp}(3,4)$).
Regular proper subalgebras of $C_{n}$ are given by $A_{p},C_{p},D_{s},$ $p+1,s\leq n.$ By Proposition 7 to obtain non-regular subalgebra in $C_{n}$ of rank greater than $\frac{n}{2}$ one has to “use” the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $C_{n}.$ By Corollary 6 any non-regular simple subalgebra in $D_{p}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
$A_{\frac{n}{2}},B_{n-1},C_{{\frac{n}{2}}},D_{{\frac{n}{2}}}\subset C_{n},$
and so we have
  • $d(\mathfrak{g})=4an+4bn-4a^{2}-4b^{2}-8ab,$
  • $d(I)=2an-a-a^{2},$ $r=a,$
  • $d(II)=2an-4a^{2},$ $r=a,$
  • $d(III)=2a^{2}+a,$ $r=a,$ $a\leq\frac{n}{2}$
We have $a+b\leq\frac{n}{2},$ $n\geq 2.$ Notice that we only have to check I and III.
(I,I) $D=4an+4bn-4a^{2}-4b^{2}-8ab-2an+a+a^{2}-2bn+b+b^{2}\geq a^{2}+b^{2}+a+b>0,$
(I,III) $D=4an+4bn-4a^{2}-4b^{2}-8ab-2an+a+a^{2}-2b^{2}-b\geq a^{2}+4ab+a+b(2b-1)>0,$
(III,III) $D=4an+4bn-4a^{2}-4b^{2}-8ab-2a^{2}-a-2b^{2}-b\geq 2(a^{2}+b^{2})+a(4b-1)+b(4a-1)>0.$
Type: $B_{n}$
In this type we only have to check $\mathfrak{g}=\mathfrak{so}(a+b,2n+1-a-b).$
We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then $\mathfrak{g}=\mathfrak{so}(3,6),$ $\mathfrak{so}(4,5).$ If $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then $\mathfrak{g}=\mathfrak{so}(s,13-s),\mathfrak{so}(s,15-s),$ $s=3,4,5,6,7.$ One verifies that there does not exist $\mathfrak{l}$ with adequate $r,d$ with one exception:
$(\mathfrak{so}(6,7),\mathfrak{so}(4,7),\mathfrak{g}_{2}(\mathbb{C})),$
but in this case $\mathfrak{g}$ is split.
Regular proper subalgebras of $B_{n}$ are given by $A_{p},B_{p},D_{s},$ $p+1,s\leq n.$ By Corollary 6 to obtain non-regular subalgebra in $B_{n}$ of rank greater than $\frac{n}{2}$ one has to consider the chain
$\mathfrak{h}^{c}\subset B_{p-1}\subset D_{p}\subset B_{n}$
where $D_{p}$ is regular in $B_{n}.$ In such case no real form of the pair $(\mathfrak{g}^{c},\mathfrak{h}^{c})$ can yield a compact standard Clifford-Klein form. Thus we may assume that each subalgebra is of rank at most $\frac{n}{2}$.
In this case we have to check Lie subalgebras of type I,II,III for $t=\frac{n}{2}.$
  • $d(\mathfrak{g})=2an+2bn-a^{2}-b^{2}-2ab+a+b,$
  • $d(I)=an+a-a^{2},$ $r=a,$
  • $d(II)=2an-4a^{2},$ $r=a,$
  • $d(III)=2a^{2}+a,$ $r=a,$ $a\leq\frac{n}{2}.$
We have $a+b\leq n.$
(I,I) $D=2an+2bn-a^{2}-b^{2}-2ab+a+b-an-a+a^{2}-bn-b+b^{2}\geq a^{2}+b^{2}>0.$
(I,II) $D=2an+2bn-a^{2}-b^{2}-2ab+a+b-an-a+a^{2}-2bn+4b^{2}\geq(a-b)^{2}+ab+b+2b^{2}>0,$
(I,III) $D=2an+2bn-a^{2}-b^{2}-2ab+a+b-an-a+a^{2}-2b^{2}-b\geq(a-b)^{2}+ab>0,$
(II,II) $D=2an+2bn-a^{2}-b^{2}-2ab+a+b-2an+4a^{2}-2bn+4b^{2}\geq(a-b)^{2}+2(a^{2}+b^{2})+a+b>0,$
(II,III) $D=2an+2bn-a^{2}-b^{2}-2ab+a+b-2an+4a^{2}-2b^{2}-b\geq(a-b)^{2}2a^{2}+a+2b>0,$
(III,III) $D=2an+2bn-a^{2}-b^{2}-2ab+a+b-2a^{2}-a-2b^{2}-b\geq(a-b)^{2}.$ Thus $D>0$ unless $a=b=\frac{n}{2}.$ But then $\mathfrak{g}=\mathfrak{so}(n,n+1),$ which is split.
Type: $D_{n}$
Notice that if $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then we have to check $D_{6}$ and $D_{7}.$ But if $\mathfrak{g}_{2}(\mathbb{C})$ is contained in Lie algebra of type $D_{p}$ then necessarily $p\geq 8.$
Regular proper subalgebras of $D_{n}$ are given by $A_{p},D_{p},$ $p\leq n.$ Corollary 6 eliminates the possibility $\operatorname{rank}\mathfrak{h}^{c}>\frac{1}{2}\operatorname{rank}\mathfrak{g}^{c}$. The opposite inequality is analyzed below.
Case $\mathfrak{g}=\mathfrak{so}^{\ast}(n)$
We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then we have to check $\mathfrak{g}=\mathfrak{so}^{\ast}(4),\mathfrak{so}^{\ast}(5).$ In both cases real rank equals 2 (which equals the real rank of $\mathfrak{g}_{2(2)}$).
By Corollary 6 any non-regular simple subalgebra in $B_{n}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
$A_{\frac{n}{2}},B_{n-1},C_{{\frac{n}{2}}},D_{{\frac{n}{2}}}\subset D_{n},$
and so we have
  • $d(\mathfrak{g})=n^{2}-n,$
  • $d(I)=2an-a-a^{2},$ $r=a,$
  • $d(II)=2an-4a^{2},$ $r=a,$
  • $d(III)=2a^{2}+a,$ $r=a,$ $a\leq\frac{n}{2}.$
We see that one only needs to check I and III. Also $n\geq 4,$ $\frac{n-1}{2}\leq a+b\leq\frac{n}{2}$ and so $a^{2}+b^{2}\leq\frac{n^{2}}{4}-2ab.$ (I,I) $D=n^{2}-n-2an+a+a^{2}-2bn+b+b^{2}\geq-n+a^{2}+a+b^{2}+b>0,$ unless $a=b=1$ and $n=4.$ But in this case we have embedding $B_{3}\subset D_{4},$ which is symmetric and we can conclude by [32].
(I,III) $D=n^{2}-n-2an+a+a^{2}-2b^{2}-b\geq 2bn-n+a+a^{2}-2b^{2}-b\geq 2b^{2}-3b+4ab+a^{2}-a-1=b(2b-3)+2a+a(a-1)+2ab-1>0,$
(III,III) $D=n^{2}-n-2a^{2}-a-2b^{2}-b\geq\frac{n^{2}}{2}-n+2ab-a-b\geq\frac{n}{2}(n-3)+2ab>0.$
Case $\mathfrak{g}=\mathfrak{so}(a+b,2n-a-b)$
We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then we have to check $\mathfrak{g}=\mathfrak{so}(3,5),$ $\mathfrak{so}(3,7),$ $\mathfrak{so}(4,6).$ One verifies that there does not exist $\mathfrak{l}$ with adequate $r,d.$
We have the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $D_{n}.$ By Corollary 6 we can eliminate this case. Thus we may assume that each subalgebra of $\mathfrak{so}(a+b,2n-a-b)$ is of rank at most $\frac{n}{2}.$
We can substitute III by III’$:=\mathfrak{so}(2a,\mathbb{C}),$ $\mathfrak{so}(2a+1,\mathbb{C}),$ $a\leq\frac{n}{2}$ We have
  • $d(\mathfrak{g})=2an+2bn-a^{2}-b^{2}-2ab,$
  • $d(I)=an-a-a^{2},$ $r=a,$
  • $d(II)=2an-4a^{2},$ $r=a,$
  • $d(III^{\prime})\leq 2a^{2}+a,$ $r=a,$ $a\leq\frac{n}{2}.$
We have $a+b\leq n.$
(I,I) $D=2an+2bn-a^{2}-b^{2}-2ab-an+a+a^{2}-bn+b+b^{2}\geq a^{2}+b^{2}-a-b>0,$ unless $a=b=1.$ In such case $D=2n-4>0,$
(I,II) $D=2an+2bn-a^{2}-b^{2}-2ab-an+a+a^{2}-2bn-4b^{2}\geq(a-b)^{2}=3b^{2}+a(b-1)>0,$
(I,III’) $D=2an+2bn-a^{2}-b^{2}-2ab-an+a+a^{2}-2b^{2}-b\geq(a-b)^{2}+a(\frac{1}{2}b-1)+b(\frac{1}{2}a-1)>0,$ unless $(a,b)=(1,1),$ $(1,2),$ $(2,1),$ $(2,2).$ As $n\geq 4$ thus $D>0$ unless $n=4,$ $a=b=2.$ In such case we have $\mathfrak{g}=\mathfrak{so}(4,4),$ $\mathfrak{h}=\mathfrak{so}(2,3),$ $\mathfrak{l}=\mathfrak{so}(4,\mathbb{C}),$ and so $D=4>0,$
(II,II) $D=2an+2bn-a^{2}-b^{2}-2ab-2an+4a^{2}-2bn+4b^{2}=(a-b)^{2}+2(a^{2}+b^{2})>0,$
(II,III’) $D=2an+2bn-a^{2}-b^{2}-2ab-2an+4a^{2}-2b^{2}-b\geq(a-b)(a-b+1)+a^{2}>0,$
(III’,III’) Notice that for the case III we have one of the following
  1. 1.
    $d=2a^{2}+a,$ with $2a+1\leq n,$
  2. 2.
    $d=2a^{2}-a,$ with $2a\leq n,$
For any combination (that is (III.1,III.1), (III.1,III.2), (III.2,III.2)) we have $D\geq(a-b)^{2}+a+b>0.$
Proposition 9.
No real form of the triple $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c})$ with at least one of the pair $(\mathfrak{h}^{c},\mathfrak{l}^{c})$ satisfies the assumptions 1) or 3) of Proposition 5 can yield a standard compact Clifford-Klein form.
Proof.
Consider first $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ which are simple, reducible and contained in direct sums of simple classical Lie algebras $\mathfrak{g}^{c}=\mathfrak{g}_{1}^{c}\oplus\mathfrak{g}^{c}_{2}$. One can notice that such subalgebra is necessarily isomorphic to a subalgebra in one of the summands, $\mathfrak{g}_{1}^{c}$ or $\mathfrak{g}_{2}^{c}$. But since $\dim\mathfrak{g}_{1}^{c}<\dim\mathfrak{g}^{c}$ and $\dim\mathfrak{g}^{c}_{2}<\dim\mathfrak{g}^{c}$, one can apply the previous considerations to the inclusion of reducible $\tilde{\mathfrak{h}}^{c}$ or $\tilde{\mathfrak{l}}^{c}$ (isomorphic to $\mathfrak{h}^{c}$ or to $\mathfrak{l}^{c}$) into one of the $\mathfrak{g}_{1}^{c}$ or $\mathfrak{g}_{2}^{c}$. Since the non-compact dimension is the invariant of a Lie algebra itself we get the same inequality for non-compact dimensions of $\mathfrak{g},\mathfrak{h},\mathfrak{l}$ ensuring that these triples cannot generate compact standard Clifford-Klein forms.
5.1.5 Proof of Theorem 2 for the semisimple case
Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ corresponds to a standard triple and let $\mathfrak{h}=\mathfrak{h}_{1}\oplus\mathfrak{h}_{2}.$ We may also assume that $\mathfrak{g}$ is of classical type and non-split. Let
$\mathfrak{h}_{i}=\mathfrak{k}_{i}\oplus\mathfrak{a}_{i}\oplus\mathfrak{n}_{i},\ \ i=1,2$
be the compatible Iwasawa decompositions. We see that $\mathfrak{h}_{1}$ is in the centralizer of $\mathfrak{a}_{2}$ in $\mathfrak{g}.$ Similarly, $\mathfrak{h}_{2}$ is in the centralizer of $\mathfrak{a}_{1}.$ Therefore, by Theorem 4 all the roots of $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ are contained on two planes: one orthogonal to $\mathfrak{a}_{1},$ the second orthogonal to $\mathfrak{a}_{2}.$ But now, by Lemma 5 for $\Sigma,$ there exists $\beta\in\Delta_{0}^{\pm}$ such that the restriction of $\beta$ to $\mathfrak{a}_{i},$ $i=1,2,$ is non-zero.
By Lemma 7 every root of $\Delta_{0}^{\pm}$ is a sum of some two roots in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ Thus $\beta=\alpha_{1}+\sigma(\alpha_{2}),$ so that $\alpha_{1}|_{\mathfrak{a}_{2}}=0$ and $\sigma(\alpha_{2})|_{\mathfrak{a}_{1}}=0.$ It follows that $\alpha_{1}$ and $\sigma(\alpha_{2})$ are orthogonal and so the root subsystem generated by those roots is necessarily of type $B_{2}.$ Thus $\mathfrak{g}$ has to by of type $B_{t}$ or $C_{t}.$ We thus have only two possibilities (we may assume that $\alpha_{1}(iX)>0$).
Case 1. $\mathfrak{g}=\mathfrak{so}(n,m),$ $n+m$ is odd. In this case for any root $\alpha$ in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ the root $\sigma(\alpha)$ has the same projection on $\mathfrak{a}.$ But in $\mathfrak{so}(n,m)$ any long root of $\Delta$ can correspond only to a 1-dimensional root space of $\mathfrak{g}.$ It follows that $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ consists only of short roots. But now $\alpha_{1}+\alpha_{2}$ is a root because a sum of any two (non-colinear) short roots in $B_{t}$ is a root.
By Theorem 4 we see that
$S_{h}^{1}=x_{\alpha_{1}}+\sum_{l=1}^{k}a^{i}_{k+l}x_{\alpha_{k+l}}+Q_{1},$
$S_{h}^{2}=x_{\alpha_{2}}+\sum_{l=1}^{k}a^{i}_{k+l}x_{\alpha_{k+l}}+Q_{2},$
belong to $\mathfrak{h}^{c}.$ We also see that $[x_{\alpha_{1}},x_{\alpha_{2}}]$ is non-zero and it is a unique factor in $[S_{h}^{1},S_{h}^{2}]$ with maximal value on $ad_{iX}.$ Thus $0\neq Y:=[S_{h}^{1},S_{h}^{2}]\in\mathfrak{h}^{c}$ and
$ad_{\mathfrak{a}_{i}}(Y)\neq 0,\ \ i=1,2.$
But this shows that $\mathfrak{h}\neq\mathfrak{h}_{1}\oplus\mathfrak{h}_{2},$ a contradiction.
Case 2 One can deal with case 2 only by using general structure theory of Lie algebras. Indeed, let $\mathfrak{g}=\mathfrak{sp}(n,m),$ where $n\leq m.$ Since $\mathfrak{g}^{c}$ is of type $C_{p}$ therefore we may represent $\Delta$ as
$\Delta=\{\pm e_{i}\pm e_{j},\pm 2e_{i}\ |\ i\neq j,1\leq i,j\leq p\},$
where $(e_{i})_{i=1}^{p}$ is a standard basis of $\mathbb{R}^{p}.$ From the general consideration in Section 5.1.5 we know that there must exist roots $\alpha_{1}$ and $\sigma(\alpha_{2})$ in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that
  1. 1.
    $\alpha_{1}$ and $\sigma(\alpha_{2})$ are orthogonal,
  2. 2.
    $\alpha_{1}(iX)=-\sigma(\alpha_{2})(iX)\neq 0,$
  3. 3.
    $\beta=\alpha_{1}+\sigma(\alpha_{2})$ is a (long) root of $\Delta$ contained in $\Delta_{0}^{\pm},$
  4. 4.
    $\alpha_{1}|_{\mathfrak{a}_{2}}=0$ and $\sigma(\alpha_{2})|_{\mathfrak{a}_{1}}=0.$
Therefore necessarily
$\alpha_{1}=e_{i}+e_{j},\ \ \ \sigma(\alpha_{2})=e_{i}-e_{j},$
for some $i\neq j$ and $e_{i}(iX)=0,$ $e_{j}(iX)>0.$ Consider the root
$\tilde{\beta}:=\alpha_{1}-\sigma(\alpha_{2})=2e_{j}.$
We see that
$\tilde{\beta}(iX)\neq 0,\ \ \tilde{\beta}|_{\mathfrak{a}_{2}}=\sigma(\alpha_{2}),\ \ \tilde{\beta}|_{\mathfrak{a}_{1}}=\alpha_{1}.$
Thus $\tilde{\beta}$ is in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and is non-zero on $\mathfrak{a}_{1}$ and on $\mathfrak{a}_{2}.$ But, as is shown in Section 5.1.5, the roots of $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ are contained in two planes: one orthogonal to $\mathfrak{a}_{1},$ the second orthogonal to $\mathfrak{a}_{2}.$ A contradiction. $\square$.
Remark 4.
The case of $\mathfrak{g}=\mathfrak{sp}(n,m),$ $n\leq m.$ can be settled also by case-by-case analysis as in Section 5.1.4. One shows that the only possibility is that $\mathfrak{h},\mathfrak{l}$ are of the form
$\mathfrak{h}=\oplus_{s=1}^{w_{h}}\mathfrak{sp}(n_{s},m_{s}),\ \ \ \mathfrak{l}=\oplus_{s=1}^{w_{l}}\mathfrak{sp}(n_{s},m_{s}).$
By [21] (Theorem 1.5), for any $s$ we have $m_{s}=m$ (otherwise one could construct a reductive subalgebra $\mathfrak{g}^{\prime}$ in $\mathfrak{g}$ with $d(\mathfrak{g}^{\prime})>d(\mathfrak{g})$ and $\mathfrak{a}_{g^{\prime}}=\mathfrak{a}_{h}$, which is impossible. But now we see that $\mathfrak{h}$ and $\mathfrak{l}$ have to be simple. A contradiction.
6 Addendum: tables and explicit calculations
6.1 Tables
Dimensions of fundamental representations
Type dimension $\bar{\omega}_{r}$
$A_{l}$ $\binom{l+1}{r}$ $20,r\not=(l+1)/2,(-1)^{r},r=(l+1)/$ 2
$B_{l}$ $\binom{2l+1}{r}$ $1,r\not=l,(-1)^{l(l+1)/2},r=l$
$B_{l}$ $2^{l}$ $(-1)^{l(l+1)/2}$
$C_{l}$ $\binom{2l}{r}-\binom{2l}{r-2}$ $(-1)^{r}$
$D_{l}$ $\binom{2l}{r}$ $21,1\leq r\leq l-$ 2
$D_{l}$ $2^{l-1}$ $0,r=l-1,\,l\,\text{odd},(-1)^{l/2},l\,\text{even}$
$D_{l}$ $2^{l-1}$ $0,\,r=l,\,l\,\text{odd},(-1)^{l/2},l\,\text{even}$
Table 2:
Non-maximal irreducible subalgebras
Type of inclusion dimension $\omega_{r}$
$A_{n}\subset A_{l}$ $3\binom{n+2}{4}$ 00
$A_{n}\subset A_{l}$ $3\binom{n+3}{4}$ 00
$B_{2n+1}\subset\mathfrak{g}^{c}$ $\prod_{s=1}^{2n}{\binom{k+s-1}{k}/\binom{k+s}{k}}$ $(-1)^{n+1)k}$
Table 3:
Irreducible triples
Inclusion 1 Inclusion 2 $\dim$ of the representation
$C_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n-1+k}{k}$
$B_{n}\subset A_{2n}$ $A_{2n}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n+1}{k}$
$D_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n}{k}$
$A_{n}\subset A_{\frac{(n-1)(n+2)}{2}}$ $A_{\frac{(n-1)(n+2)}{2}}\subset\tilde{\mathfrak{g}}^{c}$ $\frac{1}{8}(n-1)(n+1)(n+2)$
$A_{n}\subset A_{\frac{n(n+3)}{2}}$ $A_{\frac{n(n+3)}{2}}\subset\tilde{\mathfrak{g}}^{c}$ $\frac{1}{8}(n+1)(n+2)(n+3)$
$B_{n}\subset D_{n+1}$ $D_{n+1}\subset\tilde{\mathfrak{g}}^{c}$ $\prod_{s=1}^{n-1}\binom{k+2s-1}{k}/\binom{k+s}{k}$
Table 4:
6.2 Explanations to Tables
In Table 2 $\bar{\omega}_{r}$ is 00, if the representation is linear, but neither symplectic, nor orthogonal, it is $1$ is the representation is orthogonal, and it is $(-1)$ if the representation is symplectic. Table 3 reproduces inclusions $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$ of complex simple Lie algebras when $\mathfrak{h}^{c}$ is not maximal in $\mathfrak{g}^{c}$ from [14], Table 1 (recall that we reproduce only the infinite series). Table 4 reproduces inclusions of triples of irreducible inclusions $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}\subset\tilde{\mathfrak{g}}^{c}$ together with the dimensions of $\tilde{\mathfrak{g}}^{c}$. Here we need to explain some simplification in Tables 4 and 5. In [14] the author is concerned with the classification of subalgebras up to conjugation. Then non-conjugate subalgebras differ by integer labels $k$ in the corresponding Dynkin diagrams. These numbers yield different values of $\dim\tilde{\mathfrak{g}}^{c}$. However, our argument does not use this more rained classification, since we consider all possible real ranks for all possible real forms. Therefore, we do not indicate different (non-conjugate) inclusions of the same subalgebra.
6.3 An estimate
Lemma 12.
The following inequality holds for $l\geq 3$
$\binom{2l}{r}-\binom{2l}{r-2}>l^{2},2\leq r\leq l.$ (4)
Proof.
For $r=2$ one obtains
$\binom{2l}{2}-1=\frac{(2l-1)2l}{2}-1=2l^{2}-l-1>l^{2},\,\text{for}\,l>1.$
For $r=3$ one obtains
$\binom{2l}{3}-\binom{2l}{1}=\frac{1}{3}l(4l^{2}-2l-\frac{4}{3}),\,\text{for}\,l\geq 3.$
For $r=4$ one obtains
$\binom{2l}{4}-\binom{2l}{2}=\frac{1}{12}l(2l-1)(4l^{2}-10l+6),\,\text{for}\,l\geq 4.$
For $r\geq 5$ we use the formula
$\binom{n}{k}=\prod_{i=1}^{k}\frac{n+1-i}{i}.$
In our case
$\binom{2l}{r}-\binom{2l}{r-2}=\prod_{i=1}^{r}\frac{2l+1-i}{i}-\prod_{i=1}^{r-2}\frac{2l+1-i}{i}=\binom{2l}{r-2}\cdot(\frac{(2l+2-r)(2l+1-r)}{(r-1)r}-1)=$
$=\binom{2l}{r-2}\cdot\frac{4l^{2}-4lr+6l-2r+2}{r(r-1)}\geq\binom{2l}{3}\cdot\frac{4l^{2}-4lr+6l-2r+2}{r(r-1)}=(*)$
Notice that (*)(*)( * ) obtains the minimal value for $r=l$ and so
(*$(*)\geq\binom{2l}{3}\cdot\frac{4l+2}{l(l-1)}>2l^{2}-3l+1>l^{2},\,\text{for}\,l>4.$

### v3

@@
 Abstract
-We solve the long standing problem of classification of standard compact Clifford-Klein forms of homogeneous spaces of simple non-compact real Lie groups.
-The result is that standard compact Clifford-Klein forms always arise from triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ of real Lie algebras such that $\mathfrak{h}\subset\mathfrak{g},\mathfrak{l}\subset\mathfrak{g}$, $\mathfrak{g}$ is simple and absolutely simple, $\mathfrak{h},\mathfrak{l}$ are (non-compact) reductive, $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$, and the intersection $\mathfrak{h}\cap\mathfrak{l}$ is compact.
-The consequence of this is the following characterization of proper co-compact actions of reductive Lie subgroups $L\subset G$ on a homogeneous spaces $G/H$ determined by absolutely simple real Lie group $G$ and a closed reductive subgroup $H$: $L$ acts on $G/H$ properly and co-compactly if and only if $G=H\cdot L$ and $H\cap L$ is compact.
-Key words and phrases: proper action, homogeneous space, semisimple Lie algebra. 2020 Mathematics Subject Classification: 17B05, 17B10, 53C30, 57S25, 57S30.
+We solve the problem of classification of standard compact Clifford-Klein forms of homogeneous spaces of simple non-compact real Lie groups under the following assumptions:
+  1. 1.
+    either standard compact Clifford-Klein forms are given by triples$(G,H,L),$ with $G,H,L$ linear connected which arise from triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ of real Lie algebras with $\mathfrak{g}$ absolutely simple, and at least one of the complexified subalgebras $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$ or $,\mathfrak{l}^{c}\subset\mathfrak{g}^{c}$, is regular, or
+  2. 2.
+    $\mathfrak{g},\mathfrak{h},\mathfrak{l}$ are absolutely simple.
+The solution is that such triples yield standard compact Clifford-Klein forms if and only if $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$ and $\mathfrak{h}\cap\mathfrak{l}$ is compact.
+Key words and phrases: proper action, homogeneous space, semisimple Lie algebra, Clifford-Klein form, Lie algebra decomposition.
+2020 Mathematics Subject Classification: 17B05, 17B10, 53C30, 57S25, 57S30.
 1 Introduction
+1.1 The problem and motivation
 Let $G$ be a connected real semisimple linear Lie group and $H\subset G$ a closed connected reductive subgroup.
 Definition 1 (T. Kobayashi, [19],[24]).
-A homogeneous space $G/H$ admits a standard compact Clifford-Klein form, if there exists a closed reductive subgroup $L\subset G$ such that $L$ acts properly and co-compactly on $G/H.$
+A homogeneous space $G/H$ admits a standard compact Clifford-Klein form, if there exists a discrete subgroup $\Gamma\subset G$ and closed reductive subgroup $L\subset G$ containing $\Gamma$ as a co-compact lattice such that $\Gamma$ (and thus $L$) acts properly and co-compactly on $G/H.$
 This notion was first introduced in [24].
@@
 The term "standard" was first used in [19].
-Standard Clifford-Klein forms are determined by triples $(G,H,L)$, where $G$ is a semisimple real linear Lie group, and $H$ and $L$ are reductive subgroups.
+Standard Clifford-Klein forms are determined by triples $(G,H,L)$, where $H$ and $L$ are reductive subgroups.
 The general theory [24] shows that we can restrict ourselves to considering semisimple Lie subgroups $H,L$ (see Section 2.2 for details), and we will do that throughout this article.
+The following is well known.
 Theorem 1.
@@
 Table 1:
-The proof of this result is well known: all triples in Table 1 are Lie algebra decompositions in the sense of Onishchik: $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$ [29].
-It follows from this work, that $G=H\cdot L$ as well, and that $L$ acts transitively on $G/H$.
-One obtains a natural diffeomorphism $G/H\cong L/L\cap H$, and so if $H\cap L$ is compact, $L$ acts properly and co-compactly on $G/H$.
-It is important to note that these are the only known examples of compact standard Clifford-Klein forms (with $\mathfrak{g}$ simple and $\mathfrak{h}$ proper and non-compact).
-In this paper we prove that this Table contains all non-trivial triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ with semisimple $\mathfrak{h}$ and $\mathfrak{l}$ generating standard compact Clifford-Klein forms.
+It is important to note that these are the only known examples of homogeneous spaces admitting compact Clifford-Klein forms (with $\mathfrak{g}$ simple and $\mathfrak{h}$ proper and non-compact).
+All Lie algebra triples in Table 1 come from [32], Table 2.
+In this work, Onishchik classifies all triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ such that $\mathfrak{g}$ is simple, $\mathfrak{h},\mathfrak{l}$ are reductive and $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$.
+Such triples are called Lie algebra decompositions.
+We omit the complete description of the embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$, referring to [32].
+Foe example, in the last four lines one of the subalgebras is embedded via an appropriate spin representation (see [32], the proof of Theorem 4.1).
+Also, the last three lines in the table are already contained in the table as special cases with $n=1$.
+However, we prefer to leave it in the form given by Table 2 in [32], for the convenience of the reader.
+Let us also note that according to the general theory of compact Clifford-Klein forms [24], it is allowed to interchange the roles of $\mathfrak{h}$ and $\mathfrak{l}$, and we will do that throughout this article.
+In this paper we prove that under some mild restrictions Table 1 exhausts all the possibilities for compact standard Clifford-Klein forms. The main result is as follows.
 Theorem 2.
-Assume that $\mathfrak{g}$ is absolutely simple and $\mathfrak{h},$ $\mathfrak{l}$ are proper semisimple subalgebras of non-compact type.
-The only Lie algebra triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ which yield compact standard Clifford-Klein forms are contained in Table 1.
-The discussion in Section 2 shows that Theorem 2 yields a complete classification of all standard compact Clifford-Klein forms determined by triples $(G,H,L)$ with $G$ simple and $H,L$ reductive.
-Note that if $\mathfrak{g}$ is simple exceptional then, by [4], for any proper reductive subalgebra $\mathfrak{h}\subset\mathfrak{g}$ of non-compact type the corresponding homogeneous space does not admit compact standard Clifford-Klein forms.
+Assume that $\mathfrak{g}$ is absolutely simple, and $\mathfrak{h},\mathfrak{l}$ satisfy the following:
+  1. 1.
+    either the complexified embedding $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is regular, or
+  2. 2.
+    both $\mathfrak{h}$ and $\mathfrak{l}$ are proper absolutely simple subalgebras of non-compact type in $\mathfrak{g}$.
+Then only Lie algebra triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ which yield compact standard Clifford-Klein forms are Lie algebra decompositions, that is $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$.
+Note that if $\mathfrak{g}$ is absolutely simple exceptional then, by [6], for any proper reductive subalgebra $\mathfrak{h}\subset\mathfrak{g}$ of non-compact type the corresponding homogeneous space does not admit compact standard Clifford-Klein forms.
 Therefore, in the proof of Theorem 2 we may assume that $\mathfrak{g}$ is classical simple.
-It is also important to stress that the proof of Theorem 2 yields an interesting characterization of proper co-compact actions of reductive Lie groups on homogeneous spaces $G/H$ determined by simple absolutely simple Lie groups $G$ and (non-compact) reductive subgroups.
+It is interesting that the proof of Theorem 2 yields a characterization of proper co-compact actions of reductive Lie groups on homogeneous spaces $G/H$ determined by simple absolutely simple Lie groups $G$ and (non-compact) absolutely simple subgroups (we thank T.
+N.
+Venkataramana for this observation).
 Theorem 3.
-Let $G/H$ be a homogeneous space determined by a simple absolutely simple Lie group $G$ and a reductive (non-compact) closed subgroup $H$.
-A reductive Lie subgroup $L\subset G$ acts on $G/H$ properlyand co-compactly, if and only if $G=H\cdot L$ and $H\cap L$ is compact.
-There are various reasons for studying standard compact Clifford-Klein forms.
-One of them comes from spectral analysis on pseudo-Riemannian manifolds [19], [20].
-In this area, one is interested in constructing quotients $M_{\Gamma}=\Gamma\backslash M$ of discrete groups of isometries of a pseudo-Riemannian homogeneous manifold $M=G/H$.
-However, it may happen that $M$ does not admit infinite discontinuous groups of isometries (the Calabi-Markus phenomenon).
-On the other hand, there exist examples of manifolds $G/H$ admitting large discontinuous groups such that $M_{\Gamma}$ is compact, or of finite volume.
-In this case, we will call $M_{\Gamma}$ a compact Clifford-Klein form, or a Clifford-Klein form of finite volume.
-Thus, it is a challenging problem to understand when do such compact quotients exist.
-Suppose a Lie subgroup $L\subset G$ acts properly on $G/H$.
-Then the action of any discrete subgroup $\Gamma$ of $L$ is automatically properly discontinuous, and this action is free whenever $\Gamma$ is torsion-free.
-Moreover, if $\Gamma$ is co-compact in $L$, and $L$ acts co-compactly on $G/H$, then the quotient $\Gamma\backslash G/H$ is compact.
-A quotient $\Gamma\backslash G/H$ is called standard if $\Gamma$ is contained in a reductive subgroup $L$ acting properly on $G/H$.
-Thus, the first question which arises is to find standard quotients, which is equivalent to finding triples $(G,H,L)$ as in Definition 1.
-Partial results in this direction were obtained in [4] and [32].
-One of the most challenging problems in this area of research is a problem of T.
+Let $G/H$ be a homogeneous space determined by an absolutely simple Lie groups $G,H,$ $H\subset G$.
+An absolutely simple Lie subgroup $L\subset G$ acts on $G/H$ properly and co-compactly, if and only if $G=H\cdot L$ and $H\cap L$ is compact.
+There are various reasons for studying compact Clifford-Klein forms (we mention only some of them):
+  1. 1.
+    spectral analysis on pseudo-Riemannian manifolds [19], [20],[27],
+  2. 2.
+    a problem of T.
 Kobayashi: does any homogeneous space $G/H$ admitting compact Clifford-Klein form admits also a standard one?
-Clearly, one of the ways to attack it is to understand all standard Clifford-Klein forms.
-In a more geometric vein, if $\Gamma\subset G$ acts freely and co-compactly on a homogeneous space $G/H$ with non-compact $G$ and $H$, any $G$-invariant geometric structure descends onto the quotient, which yields compact manifolds with such structures.
-The topology of such manifolds can be understood, and therefore, these manifolds can serve as testing examples.
-For instance, we mention [3].
+Clearly, one of the ways to attack it is to understand all standard Clifford-Klein forms,
+  3. 3.
+    constructing various compact manifolds with particular geometric structures.
+For example, a homogeneous space of reductive type is an important example of a homogeneous pseudo-Riemannian manifold.
+This observation was exploited in [4] to explore Einstein metrics on compact Clifford-Klein forms.
+The existence of some invariant geometric structures is an obstruction to compact Clifford-Klein forms, for example, none of para-Hermitian symmetric spaces admits a compact Clifford-Klein form [25].
+Since the foundational works of Benoist, Labourie and T.
+Kobayashi [2],[3],[24] a lot of work has been done in the whole area.
+Since there are several excellent surveys of the whole topic (see [23],[25]) we prefer to refer to these works instead of a detailed exposition.
 In the context of this work, one should mention that there are many non-existence results for compact Clifford-Klein forms (not necessarily standard).
-For example, there are no standard Clifford-Klein forms of exceptional Lie groups [4].
+We have already mentioned [6].
 Topological obstructions to compact Clifford-Klein forms were found in [30],[31], [36] together with examples.
@@
 The existence of non-standard compact Clifford-Klein forms is a deep mathematical problem of a different nature.
-We refer to a very recent work [26].
-Acknowledgement.
-We thank T.
+For instance there are examples of (non-compact) Clifford-Klein forms which are in some sense strongly non-standard: in [9] one can find examples of homogeneous spaces $G/H$ of reductive type which admit proper actions of discrete subgroups of $G$ isomorphic to co-compact lattices in $O(n,1),$ $n=2,3,4$ but do not admit proper actions of any non-compact semisimple subgroups of $G.$ We also refer to a very recent work [29].
+Finally, let us mention that we always consider the problem of standard Clifford-Klein forms under the assumption that $\mathfrak{h}$ and $\mathfrak{l}$ are reductive.
+The problem of existence of compact Clifford-Klein forms is meaningful also for other types of subgroups of $G$, for example for the solvable ones.
+There are no such forms for nilpotent $N\subset G$ [2] and for solvable $\Gamma\subset G$ under rather strong restrictions on $H$ [5].
+1.2 Methods of proof of the main results
+The proofs in this article are rather long and sometimes technical, therefore we feel it might be useful for the reader to have some overview first.
+Note that by [24] when looking for a standard compact Clifford-Klein form determined by a triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ one encounters with two somewhat different tasks:
+  1. 1.
+    one needs to know when a subgroup $L$ acts properly on $G/H$, and
+  2. 2.
+    when the double coset space $\Gamma\backslash G/H$ is compact.
+The criteria of properness were found in [2] and [24].
+They are expressed in Lie algebra terms and depend on the knowledge of the embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$ and $\mathfrak{l}\hookrightarrow\mathfrak{g}$.
+This is one of the main difficulties if one wants to solve the general problem, since such knowledge may be not available.
+The second difficulty is to learn whether the given proper action is co-compact.
+According to [24] one needs to compare the non-compact dimensions of $\mathfrak{g},\mathfrak{h}$ and $\mathfrak{l}$ and prove the equality $d(\mathfrak{g})=d(\mathfrak{h})+d(\mathfrak{l})$ (still it is not an easy task, although the calculation does not depend on the embedding).
+Our methods fully use the results [21],[22], [24], but:
+  1. 1.
+    we derive a new property of the Iwasawa decompositions of $\mathfrak{g}$,$\mathfrak{h}$ and $\mathfrak{l}$ of any standard triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$.
+This property (called well-embeddedness ) is formulated in Lemma 1 and proved in [8] with the use of [1],
+  2. 2.
+    using Lemma 1 we derive some properties of the real root space decompositions of $\mathfrak{g},\mathfrak{h}$ and $\mathfrak{l}$ and their relations with root decompositions of the complexifications $\mathfrak{g}^{c}$,$\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ under the assumption that $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$ (this result was proved in [8], we repeat its formulation here, Theorem 4),
+  3. 3.
+    assuming now that the embedding $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is regular, we derive a contradiction between Theorem 4 and the regularity (Theorem 7),
+  4. 4.
+    in the non-regular case, we use Dynkin’s classification of maximal subalgebras of simple classical Lie algebras and under the assumptions that $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are both simple and non-regular, we are able to derive the inequality $d(\mathfrak{g})>d(\mathfrak{h})+d(\mathfrak{l})$ which completes the whole proof.
+The following remark seems to be in order.
+Our proofs essentially use the classification results from variuous sources: classical Dynkin’s papers [13],[14], tables of fundamental irreducible representations of simple Lie algebras [10],[11], tables of real forms of simple complex Lie algebras, of non-compact dimensions of simple Lie algebras and tables of maximal semisimple subalgebras in simple Lie algebras [33].
+Parts of these used in this article are reproduced as Tables 2-6
+Remark 1.
+We were informed by the reviewer of our paper [8] that S.
+Mehdi and M.
+Olbrich announced in [28] that there are no compact standard Clifford-Klein forms of simple absolutely simple Lie groups $G$ and non-compact $H,L$ except those whose Lie algebras $\mathfrak{g},\mathfrak{h}$ and $\mathfrak{l}$ are contained in Table 1.
+The proof has not been published yet.
+Acknowledgment.
+We thank Yves Benoist, Pralay Chatterjee, Toshiyuki Kobayashi, Yosuke Morita and T.
 N.
-Venkataramana for valuable discussions during the meeting "Zariski dense subgroups, number theory and geometric applications" held at ICTS, Bangalore in January 2024, and for pointing out an interpretation of our results described by Theorem 3.
-2 Notation and summary of [6]
-2.1 Properties of Lie algebra embeddings yielding compact Clifford-Klein forms
+Venkataramana for valuable discussions during the meeting "Zariski dense subgroups, number theory and geometric applications" held at ICTS, Bangalore in January 2024.
+Also, some parts of this work were completed during the second author research stay at the Institute of Mathematical Sciences in Chennai.
+His special thanks go to the Institute for a wonderful research atmosphere and hospitality.
+2 Preliminaries
+2.1 Notation
+Our basic references are [16],[17] and [33].
 Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple, which means that $\mathfrak{g}$ is absolutely simple, $\mathfrak{h},\mathfrak{l}$ are reductive subalgebras of $\mathfrak{g}$ and $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ induces a compact standard Clifford-Klein form.
@@
 Let $\mathfrak{a}\subset\mathfrak{p}$ be a maximal abelian subspace of $\mathfrak{p}$ and denote by $\mathfrak{m}_{0}$ the centralizer of $\mathfrak{a}$ in $\mathfrak{k}.$ Choose a Cartan subalgebra $\mathfrak{t}$ of $\mathfrak{m}_{0}.$ Then $\mathfrak{j}:=\mathfrak{t}+\mathfrak{a}$ is a Cartan subalgebra of $\mathfrak{g}$ and the complexification $\mathfrak{j}^{c}$ of $\mathfrak{j}$ is a Cartan subalgebra of $\mathfrak{g}^{c}.$ Let $\Delta$ be a root system of $\mathfrak{g}^{c}$ with respect to $\mathfrak{j}^{c}.$ Consider the root decomposition
-$\mathfrak{g}^{c}=\mathfrak{j}^{c}+\sum_{\alpha_{c}\in\Delta}\mathfrak{g}_{\alpha_{c}}.$
-The Lie algebra $\mathfrak{j}:=i\mathfrak{t}+\mathfrak{a}$ is a real form of $\mathfrak{j}^{c}$ with
-$\mathfrak{j}=\{A\in\mathfrak{j}^{c}\ |\ \alpha_{c}(A)\in\mathbb{R}\ \textrm{for any}\ \alpha_{c}\in\Delta\}.$
+$\mathfrak{g}^{c}=\mathfrak{j}^{c}+\sum_{\alpha\in\Delta}\mathfrak{g}_{\alpha}.$
+The Lie algebra $\hat{\mathfrak{j}}:=i\mathfrak{t}+\mathfrak{a}$ is a real form of $\mathfrak{j}^{c}$ and
+$\hat{\mathfrak{j}}=\{A\in\mathfrak{j}^{c}\ |\ \alpha(A)\in\mathbb{R}\ \textrm{for any}\ \alpha\in\Delta\}.$
 The weights of the adjoint representation of $\mathfrak{a}$ constitute a root system $\Sigma\subset\mathfrak{a}^{*}$ (not necessarily reduced).
-There is a natural map $pr:\Delta\rightarrow\Sigma$ given by $\alpha_{c}\rightarrow\alpha_{c}|_{\mathfrak{a}}$.
-It is easy to see that $pr(\Delta)=\Sigma\cup\{0\}.$ Therefore, we will call $pr(\alpha_{c})=\alpha_{c}|_{\mathfrak{a}}\in\mathfrak{a}^{*}$ (if non-zero) the restricted roots.
-Thus, $\Sigma=\{\alpha:=\alpha_{c}|_{\mathfrak{a}}\ |\ \alpha_{c}\in\Delta\}\setminus\{0\},$
-$\mathfrak{g}_{\alpha}:=\sum_{\begin{subarray}{c}\alpha_{c}\in\Delta\\ \alpha_{c}|_{\mathfrak{a}}=\alpha\end{subarray}}\mathfrak{g}_{\alpha_{c}}\cap\ \mathfrak{g}.$
+There is a natural map $pr:\Delta\rightarrow\mathfrak{a}^{*}$ given by $\alpha\rightarrow\alpha|_{\mathfrak{a}}$.
+It is easy to see that $pr(\Delta)=\Sigma\cup\{0\}.$ Therefore, we will call $pr(\alpha)=\alpha|_{\mathfrak{a}}\in\mathfrak{a}^{*}$ (if non-zero) the restricted roots.
+Thus, $\Sigma=\{\alpha|_{\mathfrak{a}}\ |\ \alpha\in\Delta\}\setminus\{0\},$
+$\mathfrak{g}_{\beta}:=\sum_{\begin{subarray}{c}\alpha\in\Delta\\ \alpha|_{\mathfrak{a}}=\beta\end{subarray}}\mathfrak{g}_{\alpha}\cap\ \mathfrak{g}.$
 The reader should keep in mind, that we use the following convention.
-Roots of complex Lie algebra $\mathfrak{g}^{c}$ are denoted by $\alpha_{c},\beta_{c},,...$, the corresponding restrictions are $\alpha,\beta,...$ and so on.
-Vectors from the root spaces $\mathfrak{g}_{\alpha_{c}}$ of $\mathfrak{g}^{c}$ and $\mathfrak{g}_{\alpha}$ from $\mathfrak{g}$ are always denoted by $x_{\alpha},x_{\beta},...$.
-The meaning of these will be always clear from the context.
-However, to avoid clumsy notation, in several cases we write $\alpha_{i}$ for complex roots.
-In each case we state explicitly, which root (complex or its restriction) is used.
+We use the same notation for complex roots in $\Delta$ and real (restricted) roots in $\Sigma$.
+The meaning of these will be always clear either from the context, or explicitly stated.
 We get the following decomposition of $\mathfrak{g}$
-$\mathfrak{g}=\mathfrak{m}_{0}+\mathfrak{a}+\sum_{\alpha\in\Sigma}\mathfrak{g}_{\alpha}.$
+$\mathfrak{g}=\mathfrak{m}_{0}+\mathfrak{a}+\sum_{\alpha\in\Sigma}\mathfrak{g}_{\alpha},$
+where $\mathfrak{m}_{0}=\mathfrak{z}_{\mathfrak{g}}(\mathfrak{a})$.
 Let $B$ be the Killing form of $\mathfrak{g}.$ It extends onto $\mathfrak{g}^{c}$ as a Killing form of $\mathfrak{g}^{c}$ and we denote it by the same letter.
@@
 It is well known ([34], p.
-228) that there is a Hermitian non-degenerate form on $\mathfrak{g}^{c}$ given by the formula $B_{\tau}(X,Y)=-B(X,\tau(Y)).$ Choose a set of positive restricted roots $\Sigma^{+}\subset\Sigma.$ There is a system of positive roots $\Delta^{+}\subset\Delta$ such that
-$\Sigma^{+}=\{\alpha_{c}|_{\mathfrak{a}}\ |\ \alpha_{c}\in\Delta^{+}\}-\{0\}.$
+228) that there is a Hermitian positive-definite form on $\mathfrak{g}^{c}$ given by the formula $B_{\tau}(X,Y)=-B(X,\tau(Y)).$ Choose a set of positive restricted roots $\Sigma^{+}\subset\Sigma.$ There is a system of positive roots $\Delta^{+}\subset\Delta$ such that
+$\Sigma^{+}=\{\alpha|_{\mathfrak{a}}\ |\ \alpha\in\Delta^{+}\}-\{0\}.$
 The subalgebra $\mathfrak{n}:=\sum_{\alpha\in\Sigma^{+}}\mathfrak{g}_{\alpha}$ is a nilpotent subalgebra of $\mathfrak{g}.$ The decomposition
@@
 as well as $\mathfrak{n}_{h}^{-},\mathfrak{n}_{l}^{-}$ for $\mathfrak{h}$ and $\mathfrak{l}.$ Since $\theta(\mathfrak{g}_{\alpha})=\mathfrak{g}_{-\alpha}$ for all $\alpha\in\Sigma$ ([16], Proposition 5.9) we get $\theta(\mathfrak{n})=\mathfrak{n}^{-}$.
+We will denote by $\Sigma_{h}$, $\Sigma_{l}$ the real root systems of $\mathfrak{h}$ and $\mathfrak{l}$, respectively, and consider $\mathfrak{n}_{h}$ and $\mathfrak{n}_{l}$ corresponding to $\Sigma_{h}$ and $\Sigma_{l}$.
+2.2 Properties of Lie algebra embeddings yielding compact Clifford-Klein forms
 Lemma 1 ([8]).
-Under the assumptions of Theorem 4, we may assume that
+Let $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ be a standard triple. Then we may assume that
 $\mathfrak{n}=\mathfrak{n}_{h}\oplus\mathfrak{n}_{l},\ \ \mathfrak{n}^{-}=\mathfrak{n}_{h}^{-}\oplus\mathfrak{n}_{l}^{-},\ \ \mathfrak{a}=\mathfrak{a}_{h}\oplus\mathfrak{a}_{l},$
 $\theta(\mathfrak{h})=\mathfrak{h},\ \ \theta(\mathfrak{l})=\mathfrak{l}.$
-We will denote by $\Sigma_{h}$, $\Sigma_{l}$ the real root systems of $\mathfrak{h}$ and $\mathfrak{l}$, respectively, corresponding to the above decompositions.
+We will denote by $\Sigma_{h}$, $\Sigma_{l}$ the restricted root systems of $\mathfrak{h}$ and $\mathfrak{l}$, respectively, corresponding to the above decompositions.
 Definition 2.
@@
 Assume that $0\not=X\in\mathfrak{t}$. Define
-$\Delta_{m}=\{\alpha_{c}\in\Delta\,|\,\alpha_{c}|_{\mathfrak{a}}=0\},$
+$\Delta_{m}=\{\alpha\in\Delta\,|\,\alpha|_{\mathfrak{a}}=0\},$
 $\Delta_{m}^{+}=\Delta^{+}\setminus\Delta_{m},\Delta_{m}^{-}=\Delta^{-}\setminus\Delta_{m},$
-$\Delta_{p}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)>0\},\,\Delta_{n}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)<0\},$
-$\Delta_{p}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)>0\},\,\Delta_{n}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)<0\},$
-$\Delta_{0}^{+}=\{\alpha_{c}\in\Delta_{m}^{+}\,|\,\alpha_{c}(iX)=0\},\ \Delta_{0}^{-}=\{\alpha_{c}\in\Delta_{m}^{-}\,|\,\alpha_{c}(iX)=0\},$
+$\Delta_{p}^{+}=\{\alpha\in\Delta_{m}^{+}\,|\,\alpha(iX)>0\},\,\Delta_{n}^{+}=\{\alpha\in\Delta_{m}^{+}\,|\,\alpha(iX)<0\},$
+$\Delta_{p}^{-}=\{\alpha\in\Delta_{m}^{-}\,|\,\alpha(iX)>0\},\,\Delta_{n}^{-}=\{\alpha\in\Delta_{m}^{-}\,|\,\alpha(iX)<0\},$
+$\Delta_{0}^{+}=\{\alpha\in\Delta_{m}^{+}\,|\,\alpha(iX)=0\},\ \Delta_{0}^{-}=\{\alpha\in\Delta_{m}^{-}\,|\,\alpha(iX)=0\},$
 $\Delta_{0}=\Delta_{0}^{+}\cup\Delta_{0}^{-}.$
+Let us now formulate the basic technical tool used in the proof of the main theorem.
+This tool is a restriction on the root system of $\mathfrak{g}^{c}$ derived under the assumption that the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ yields a standard compact Clifford-Klein form, but $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$.
 Theorem 4 ([8]).
+Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple of well embedded subalgebras and that $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$.
+We may assume that there exists a non-zero $X\in\mathfrak{t}$ such that $X$ is $B$-orthogonal to $\mathfrak{h}+\mathfrak{l}$.
 Let
-$Z=\sum_{\alpha_{c}\in\Delta_{p}^{+}\cup\Delta_{n}^{+}}\mathfrak{g}_{\alpha_{c}}\subset\mathfrak{n}^{c}.$
-Let $\pi:\mathfrak{n}^{c}=\sum_{\alpha_{c}\in\Delta_{m}^{+}}\mathfrak{g}_{\alpha_{c}}\rightarrow Z$ be the natural projection, $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$.
+$Z=\sum_{\alpha\in\Delta_{p}^{+}\cup\Delta_{n}^{+}}\mathfrak{g}_{\alpha}\subset\mathfrak{n}^{c}.$
+Let $\pi:\mathfrak{n}^{c}=\sum_{\alpha\in\Delta_{m}^{+}}\mathfrak{g}_{\alpha}\rightarrow Z$ be the natural projection, $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$.
 Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is a standard triple and that there exists a non-zero $X\in\mathfrak{t}$ such that $X\not\in(\mathfrak{h}+\mathfrak{l})$.
-We may assume that there exists a non-zero $X\in\mathfrak{t}$ such that $X$ is $B$-orthogonal to $\mathfrak{h}+\mathfrak{l}.$ Then:
+We may assume that $X\in\mathfrak{t}$ and that $X$ is $B$-orthogonal to $\mathfrak{h}+\mathfrak{l}.$ Then:
   1. 1.
@@
     where $\alpha_{i_{1}},...,\alpha_{i_{s}}$ are all roots from $\Delta_{p}^{+}$ whose restrictions onto $\mathfrak{a}_{h}$ coincide with $\gamma$, while all $Q_{j}$ satisfy the conditions
-    $Q_{j}\in\sum_{\alpha_{c}\in\Delta_{0},\ \alpha_{c}|_{\mathfrak{a}_{h}}=\gamma}\mathfrak{g}_{\alpha_{c}}.$
+    $Q_{j}\in\sum_{\alpha\in\Delta_{0},\ \alpha|_{\mathfrak{a}_{h}}=\gamma}\mathfrak{g}_{\alpha_{c}}.$
 Analogous conditions hold for $\mathfrak{l}_{\gamma}^{c}$.
-Corollary 1.
+Corollary 1 ([8]).
 If $(\mathfrak{g},\mathfrak{h},\mathfrak{l}),$ with $\mathfrak{g}$ absolutely simple and split, yields a standard compact Clifford-Klein form then $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}.$ Thus $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ is contained in Table 1.
-We consider triples $(G,H,L)$ which yield Clifford-Klein forms up to an equivalence which is given by the following well-known fact.
-Proposition 2.
-If a triple $(G,H,L)$ yields a standard compact Clifford-Klein form then so does $(G,g_{1}Hg_{1}^{-1},g_{2}Lg_{2}^{-1})$ for any $g_{1},g_{2}\in G.$
-This enables us to choose embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$, $\mathfrak{l}\hookrightarrow\mathfrak{g}$ up to conjugacy in $G$, and we will do this throughout without further notice.
-It is sufficient to consider the condition $G=HL$ on the Lie algebra level, because of the classical result of Onishchik.
-Theorem 5 ([29], Theorem 3.1).
-Let $\mathfrak{g}$ be a reductive real Lie algebra and let $\mathfrak{g}^{\prime},$ $\mathfrak{g}^{\prime\prime}$ be two reductive subalgebras.
-Let $G$ be a connected Lie subgroup with Lie algebra $\mathfrak{g}$ and let $G^{\prime}$ and $G^{\prime\prime}$ be its connected Lie subgroups corresponding to Lie algebras $\mathfrak{g}^{\prime}$ and $\mathfrak{g}^{\prime\prime}$.
-Then $G=G^{\prime}G^{\prime\prime}$ if and only if $\mathfrak{g}=\mathfrak{g}^{\prime}+\mathfrak{g}^{\prime\prime}.$
-Theorem 6 ([2], Corollary 3, see also [21]).
-If $G/H$ admits a compact Clifford-Klein form then the center of $H$ is compact.
-Let
+Theorem 4 is quite technical, so we think it will be instructive to support it by example which somehow explains the consequences of the relations between complex and real root systems caused by the assumption $\mathfrak{g}\not=\mathfrak{h}+\mathfrak{l}$.
+These relations together with Lemma 4 yield a contradiction.
+Example 1.
+We follow notation from Proposition 4, the reader should consult it first. Let $\mathfrak{g}=\mathfrak{so}(n,n+2)$. Looking at Table 4 in [33] one can look at the Satake diagram in this case
+[…]
+and notice that $\dim\mathfrak{m}_{0}=1$ and that $\Delta_{m}=\emptyset$. It follows easily that necessarily
+$X=H_{n+1},\Delta_{p}^{+}=\{\bar{\alpha}_{n}=\alpha_{n}+\alpha_{n+1}\},\Delta_{p}^{-}=\{\bar{\alpha}_{n}=\alpha_{n}-\alpha_{n+1}\}.$
+By Lemma 4 every root in $\Delta_{0}-\Delta_{m}$ is a sum of a root from $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and a root from $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}\cup\Delta_{m}$.
+Since $\Delta_{m}=\emptyset$ we get $\Delta=D_{n+1}$ and therefore it must be
+$\Delta=\{\pm\alpha_{n}\pm\alpha_{n+1}\},$
+but then the only possibility is $\Delta=D_{2}$, which implies $\mathfrak{g}=\mathfrak{so}(1,3)$ which never yields compact Clifford-Klein forms (with $\mathfrak{h}$ proper and of non-compact type) because the real rank of $\mathfrak{g}$ is 1.
+We consider triples $(G,H,L)$ which yield Clifford-Klein forms up to an equivalence which is given by the following well-known fact: if a triple $(G,H,L)$ yields a standard compact Clifford-Klein form then so does $(G,g_{1}Hg_{1}^{-1},g_{2}Lg_{2}^{-1})$ for any $g_{1},g_{2}\in G.$ This enables us to choose embeddings $\mathfrak{h}\hookrightarrow\mathfrak{g}$, $\mathfrak{l}\hookrightarrow\mathfrak{g}$ up to conjugacy in $G$, and we will do this throughout without further notice.
+By the general theory of compact Clifford-Klein forms and [32], Theorem 3.1 we can work on the Lie algebra level and study Lie algebra decompositions and their relation to the problem of standard compact Clifford-Klein forms.
+Also, using the well known result [3] we restrict ourselves to the case when $\mathfrak{h}$ and $\mathfrak{l}$ are semisimple and are direct sums of ideals of non-compact type (one can also consult [8]).
+Put
 $d(G):=\dim\,G/K,\,d(H):=\dim\,H/K_{h},\,d(L):=\dim\,L/K_{l}.$
@@
 $d(G)=d(H)+d(L),\,\text{or, equivalently}\,\,d(\mathfrak{g})=d(\mathfrak{h})+d(\mathfrak{l}).$ (1)
-By [2] we may assume that the centers of $H$ and $L$ are compact.
-Let $\mathfrak{h}^{\prime}\subset\mathfrak{h}$ be a semisimple Lie algebra given by the sum of all simple ideals of $\mathfrak{h}$ which are of non-compact type.
-Let $H^{\prime}\subset H$ be the closed connected subgroup corresponding to $\mathfrak{h}^{\prime}.$ We see that $L$ acts properly on $G/H^{\prime}$ and that $d(G)=d(H^{\prime})+d(L).$ This way the triple $(G,H^{\prime},L)$ induces a standard compact Clifford-Klein form.
-2.2 Chevalley bases compatible with non-compact real forms
-Given a root $\alpha_{c}$ and a root space $\mathfrak{g}_{\alpha_{c}}$ we denote by $t_{\alpha}\in\mathfrak{j}^{c}$ the vector determined by the equality $B(t_{\alpha},h)=\alpha_{c}(h)$ for all $h\in\mathfrak{j}^{c}$.
+Now we cite a theorem of T.
+Kobayashi which plays an important role in the proof of the main theorem of this article: it eliminates some (otherwise possible) triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ from the list of candidates.
+Theorem 6 ([21], Theorem 1.5).
+Let $G/H$ be a homogeneous space of reductive type. If $G/H$ has a compact Clifford-Klein form, then there does not exist a closed subgroup $G^{\prime}$ reductive in $G$ satisfying the conditions
+  • $\mathfrak{a}_{g^{\prime}}\subseteq\mathfrak{a}_{h}$,
+  • $d(G^{\prime})>d(H)$.
+In this work we consider embeddings of real Lie algebras $\mathfrak{h}\subset\mathfrak{g}$ and their complexifications $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$.
+Also, if for an embedding of complex Lie algebras $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$ there exist real forms $\mathfrak{h}$ of $\mathfrak{h}^{c}$ and $\mathfrak{g}\subset\mathfrak{g}^{c}$ such that $\mathfrak{h}\subset\mathfrak{g}$, we say that the embedding $\mathfrak{h}\subset\mathfrak{g}$ is the real form of the complex embedding $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$.
+Given a root $\alpha$ and a root space $\mathfrak{g}_{\alpha}$ we denote by $t_{\alpha}\in\mathfrak{j}^{c}$ the vector determined by the equality $B(t_{\alpha},h)=\alpha(h)$ for all $h\in\mathfrak{j}^{c}$.
 Set $h_{\alpha}=2t_{\alpha}/B(t_{\alpha},t_{\alpha}).$
 Definition 3.
-A Chevalley basis of $\mathfrak{g}^{c}$ with respect to $\mathfrak{j}^{c}$ is a basis of $\mathfrak{g}^{c}$ consisting of $x_{\alpha}\in\mathfrak{g}_{\alpha_{c}}$ and $h_{\alpha}$ with the following properties:
+A Chevalley basis of $\mathfrak{g}^{c}$ with respect to $\mathfrak{j}^{c}$ is a basis of $\mathfrak{g}^{c}$ consisting of $x_{\alpha}\in\mathfrak{g}_{\alpha}$ and $h_{\alpha}$ with the following properties:
   1. 1.
-    $[x_{\alpha},x_{-\alpha}]=-h_{\alpha},\,\forall\alpha_{c}\in\Delta$,
+    $[x_{\alpha},x_{-\alpha}]=-h_{\alpha},\,\forall\alpha\in\Delta$,
   2. 2.
-    for each pair $\alpha_{c},\beta_{c}\in\Delta$ such that $\alpha_{c}+\beta_{c}\in\Delta$ the constants $c_{\alpha,\beta}\in\mathbb{C}$ determined by $[x_{\alpha},x_{\beta}]=c_{\alpha,\beta}x_{\alpha+\beta}$ satisfy $c_{\alpha,\beta}=c_{-\alpha,-\beta}$.
+    for each pair $\alpha,\beta\in\Delta$ such that $\alpha+\beta\in\Delta$ the constants $c_{\alpha,\beta}\in\mathbb{C}$ determined by $[x_{\alpha},x_{\beta}]=c_{\alpha,\beta}x_{\alpha+\beta}$ satisfy $c_{\alpha,\beta}=c_{-\alpha,-\beta}$.
 Throughout, we work with a special choice of the Chevalley basis compatible with a non-compact real form $\mathfrak{g}$ of $\mathfrak{g}^{c}$.
-Lemma 2.
-For the chosen $\mathfrak{j}^{c}$ in $\mathfrak{g}^{c}$ there is a base of $\sum_{\alpha_{c}\in\Delta}\mathfrak{g}_{\alpha_{c}}$ of the form
-$\{\tilde{x}_{\alpha}\ |\ \tilde{x}_{\alpha}\in\mathfrak{g}_{\alpha_{c}},\ \alpha_{c}\in\Delta\},$
-such that for every $\alpha_{c}\in\Delta$
+Lemma 2 ([8]).
+For the chosen $\mathfrak{j}^{c}$ in $\mathfrak{g}^{c}$ there is a base of $\sum_{\alpha\in\Delta}\mathfrak{g}_{\alpha}$ of the form
+$\{\tilde{x}_{\alpha}\ |\ \tilde{x}_{\alpha}\in\mathfrak{g}_{\alpha},\ \alpha\in\Delta\},$
+such that for every $\alpha\in\Delta$
 $\tau(\tilde{x}_{\alpha})=\tilde{x}_{-\alpha},\ \ B(\tilde{x}_{\alpha},\tilde{x}_{-\alpha})=-1.$
-Proof.
-Note that $\tau(\mathfrak{j}^{c})=\mathfrak{j}^{c}$ and therefore $\tau$ permutes the root spaces.
-Since $x_{\alpha}+\tau(x_{\alpha})\in\mathfrak{g}_{u}$ we get $B(x_{\alpha}+\tau(x_{\alpha}),x_{\alpha}+\tau(x_{\alpha}))<0$.
-Note that $B(x_{\alpha},x_{\beta})\not=0$ if and only if $\alpha_{c}=-\beta_{c}$.
-Hence $B(x_{\alpha},\tau(x_{\alpha}))<0$ and therefore $\tau(x_{\alpha})=ax_{-\alpha}$ for some $a>0$.
-Since $\tau$ is an involution, $\tau(x_{-\alpha})=\frac{1}{a}x_{\alpha}$.
-Therefore, one can take $\tilde{x}_{\alpha}=\frac{1}{\sqrt{ab}}x_{\alpha}$ and $\tilde{x}_{-\alpha}=\frac{\sqrt{a}}{\sqrt{b}}x_{-\alpha}$.
-∎
-Remark 1.
-Note that in this subsection we consider only root systems of semisimple complex Lie algebras, and, therefore, follow the standard notation.
-However, in the subsequent sections we distinguish between real and complex roots.
 3 Some properties of root systems
-In this subsection we establish some properties of abstract root systems.
-Our aim is Lemma 5 which describes subsets of the root system consisting of roots simultaneously vanishing on a pair of vectors.
-This may be known, but we did not find any appropriate reference.
+In this subsection we present Lemma 3 (proved in [8]) which describes subsets of the root system consisting of roots simultaneously vanishing on a pair of vectors.
+Lemma 3 is used in the proof of Lemma 4 which will be essential in the proof of the main result.
 Let $\Delta\subset\mathbb{R}^{n}$ be an indecomposable root system in the Euclidean space $(\mathbb{R}^{n},(,))$ ([33], Chapter 3).
@@
 There is a unique maximal root $\beta\in\Delta^{+}$ such that for every $\alpha\in\Delta^{+}$ vector $\beta-\alpha$ is a combination of simple roots with non-negative coefficients.
-Lemma 3.
-Let $A$ be a vector in the interior of the positive Weyl chamber. Then $\beta(A)>\alpha(A)$ for any $\alpha\in\Delta^{+}-\{\beta\}.$
-Proof.
-Since $A$ is in the interior of the positive Weyl chamber, $\alpha_{i}(A)>0$ for any simple root $\alpha_{i}\in\Pi.$ Thus for $\alpha\in\Delta^{+},$ $\alpha\neq\beta$ we obtain $(\beta-\alpha)(A)>0.$
-Proposition 3 ([30], Chapter 3, Proposition 1.2).
-Let $\gamma_{1},\gamma_{2}\in\Delta.$ If $(\gamma_{1},\gamma_{2})<0$ then $\gamma_{1}+\gamma_{2}\in\Delta.$
-Lemma 4.
-If $\beta\in\Delta^{+}$ is such that for any $\alpha\in\Delta^{+}$ the vector $\beta+\alpha$ is not a root then $\beta$ is the highest root.
-Proof.
-Assume that $\beta,\gamma\in\Delta^{+}$ are distinct and have the property that $\beta+\alpha$ and $\gamma+\alpha$ are not roots for any $\alpha\in\Delta^{+}.$ By Proposition 3 for any simple root $\alpha_{i}$ we get $(\beta,\alpha_{i}),(\gamma,\alpha_{i})\geq 0.$ As $\beta\neq\gamma,$ for some simple root $\alpha_{i}$ we get $(\beta,\alpha_{i})\neq(\gamma,\alpha_{i})$.
-So $(\beta+\gamma,\alpha_{i})>0.$ By Proposition 3 (setting $\gamma_{1}=\alpha+\beta,$ $\gamma_{2}=-\alpha_{i}$) we see that $\theta:=\beta+\gamma-\alpha_{i}$ is a root.
-Since $\beta$ is the highest root, $\gamma$ is a positive root and $\alpha_{i}$ is simple (a lowest root) therefore $\gamma-\alpha_{i}=0$ (otherwise $\beta-\theta$ written as a combination of positive roots would have some negative coefficients).
-But $\gamma$ is a simple root and so does not have the assumed property that $\gamma+\alpha$ is not a root for every positive $\alpha$ (unless $\Delta=\{\alpha,-\alpha\}$ in which case $\gamma=\beta$).
-∎
-Lemma 5.
-Let $X,H\in\mathbb{R}^{n}$ be non-zero. Define
+Lemma 3 ([8]).
+Let $\Delta\subset\mathbb{R}^{n}$ be an indecomposable root system and $X,H\in\mathbb{R}^{n}-\{0\}$ be non-zero vectors. Define
 $C_{X}:=\{\alpha\in\Delta\ |\ \alpha(X)=0\},\ \ C_{H}:=\{\alpha\in\Delta\ |\ \alpha(H)=0\}.$
 Then $\Delta\neq C_{X}\cup C_{H}.$
+The following Lemma was proved in [8], but it is so essential, that we repeat the proof here.
+Lemma 4.
+None of the non-zero root vectors $x_{\beta}$ for $\beta\in\Delta_{0}-\Delta_{m}$ centralizes $Z+\tau(Z).$
 Proof.
-Assume that $\Delta=C_{X}\cup C_{H}.$ Consider sets $C_{H}\setminus C_{X}$ and $C_{X}\setminus C_{H}.$ If one of them is empty then all the roots in $\Delta$ are orthogonal to $X$ or to $H$, but this is impossible.
-Therefore there exist simple roots $\gamma_{1},\gamma_{2}$ (for any choice of a subset of simple roots) such that
-$\gamma_{1}\in C_{H}\setminus C_{X},\ \ \gamma_{2}\in C_{X}\setminus C_{H},$
-because any root can be represented as a linear combination of simple roots. Since $C_{H}\cap C_{X}$ is a root system, we can take the set of positive roots $C_{HX}^{+}\subset C_{H}\cap C_{X}.$ Let
-$C_{X}^{+}:=\{\alpha\in C_{H}\setminus C_{X}\ |\ \alpha(X)>0\},$
-$C_{H}^{+}:=\{\alpha\in C_{X}\setminus C_{H}\ |\ \alpha(H)>0\},$
-One can see that for $\alpha\in C_{X}^{+}$ and $\beta\in C_{H}^{+}$, vector $\alpha+\beta$ is not a root, because otherwise one would obtain a root which is non-zero on $H$ and on $X.$ Consider the set
-$\tilde{\Delta}^{+}:=C_{HX}^{+}\cup C_{X}^{+}\cup C_{H}^{+}.$
-Note that for any $\alpha\in\Delta$ exactly one of the roots $\alpha,$ $-\alpha$ is contained in $\tilde{\Delta}^{+}.$ Also if $\alpha,\gamma\in\tilde{\Delta}^{+}$ and $\alpha+\gamma$ is a root, then $\alpha+\gamma\in\tilde{\Delta}^{+}.$ Therefore $\tilde{\Delta}^{+}$ is the set of positive roots from $\Delta$ with the property
-$\forall_{\alpha\in\tilde{\Delta}^{+}}\alpha(X)\geq 0,\ \alpha(H)\geq 0.$
-Consider the highest root $\tilde{\beta}\in\tilde{\Delta}^{+}.$ Since $\Delta$ is indecomposable $\tilde{\beta}$ is a linear combination of all simple roots of $\tilde{\Delta}^{+}$ with positive coefficients.
-But we have seen before that there are simple roots $\tilde{\gamma}_{1},$ $\tilde{\gamma}_{2}\in\tilde{\Delta}^{+}$ such that
-$\tilde{\gamma}_{1}(X)>0,\ \ \tilde{\gamma}_{2}(H)>0.$
-Thus $\tilde{\beta}(X)>0$ and $\tilde{\beta}(H)>0.$ A contradiction. ∎
-Proposition 4 ([6]).
-Assume that $X\in\mathfrak{t}$ satisfies the assumptions of Theorem 4 and put $Z_{h}=\pi(\mathfrak{n}_{h}^{c}),Z_{l}=\pi(\mathfrak{n}_{l}^{c})$. Then
-$Z=Z_{h}\oplus[X,Z_{h}]=Z_{l}\oplus[X,Z_{l}]=Z_{h}\oplus Z_{l}=Z_{h}\oplus[X,Z_{l}].$
-Lemma 6 ([6]).
-Under the assumptions of Theorem 4, The involution $\sigma$ permutes the root spaces given by $\Delta^{+}_{m}$ so that if $\alpha_{c}\in\Delta_{p}^{+}$ then $\sigma(\mathfrak{g}_{\alpha_{c}})\subset\mathfrak{g}_{\alpha_{c}^{{}^{\prime}}}$ for some $\alpha_{c}^{{}^{\prime}}\in\Delta_{n}^{+}$ (by abuse of notation we write $\sigma(\Delta_{p})=\Delta_{n}$).
-Thus Analogous properties hold for $\Delta_{p}^{-},\Delta_{n}^{-}.$
-Lemma 7 ([6]).
-None of the non-zero root vectors $x_{\beta}$ for $\beta_{c}\in\Delta_{0}-\Delta_{m}$ centralizes $Z+\tau(Z).$
-Proof.
-Assume there exists $x_{\beta},\beta_{c}\in\Delta_{0}$ such that $[x_{\beta},Z+\tau(Z)]=0$.
+Assume there exists $x_{\beta},\beta\in\Delta_{0}$ such that $[x_{\beta},Z+\tau(Z)]=0$.
 Take $H_{\beta}\in i\mathfrak{t}+\mathfrak{a}$, defined, as usual by the condition $B(\tilde{H},H_{\beta})=\beta(\tilde{H})$ for any $\tilde{H}\in i\mathfrak{t}+\mathfrak{a}$.
@@
 Define
-$C_{X}=\{\alpha_{c}\in\Delta\,|\alpha_{c}(X)=0\},\,C_{H}=\{\alpha_{c}\in\Delta\,\,|\,\alpha_{c}(H)=0\}.$
+$C_{X}=\{\alpha\in\Delta\,|\alpha(X)=0\},\,C_{H}=\{\alpha\in\Delta\,\,|\,\alpha(H)=0\}.$
 Consider the decomposition
 $\Delta=\Delta_{m}\cup\Delta_{0}\cup\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$
-Note that for any $\mu_{c}\in\Delta_{m}$, $\mu_{c}(H)=0$. Assume that
+Note that for any $\mu\in\Delta_{m}$, $\mu(H)=0$. Assume that
 $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}\subset C_{H}.$ (2)
@@
 $\Delta=C_{X}\cup C_{H},$
-but this contradicts Lemma 5. Thus there exists $\alpha_{c},-\alpha_{c}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha_{c}(H)\not=0$. We have one of the following possibilities
+but this contradicts Lemma 3. Thus there exists $\alpha,-\alpha\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha(H)\not=0$. We have one of the following possibilities
   1. 1.
-    $\alpha_{c}(iT)\neq-\alpha_{c}(H).$ In this case (since one can choose between $\alpha_{c}$ and $-\alpha_{c}$) we may assume that $\alpha_{c}(H_{\beta})<0.$
+    $\alpha(iT)\neq-\alpha(H).$ In this case (since one can choose between $\alpha$ and $-\alpha$) we may assume that $\alpha(H_{\beta})<0.$
   2. 2.
-    $\alpha_{c}(iT)=-\alpha_{c}(H).$ In this case we see that $\sigma(\alpha_{c})(H_{\beta})=\alpha_{c}(-iT+H)=2\alpha_{c}(H)\neq 0$ and by Lemma 6, $\sigma(\alpha_{c})\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ In this case (since one can choose between $\sigma(\alpha_{c})$ and $-\sigma(\alpha_{c})$) we may assume that $\sigma(\alpha_{c})(H_{\beta})<0.$
-Therefore in any case there exists $\alpha_{c}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha_{c}(H_{\beta})<0.$ But for any root vector $x_{\alpha}\in Z+\tau(Z)$ we have $\operatorname{ad}_{H_{\beta}}(x_{\alpha})=\alpha_{c}(H_{\beta})x_{\alpha}$ and so for $x_{\alpha}\in\mathfrak{g}^{x_{\beta}},$ $j=\alpha_{c}(H_{\beta})$.
+    $\alpha(iT)=-\alpha(H).$ In this case we see that $\sigma(\alpha)(H_{\beta})=\alpha(-iT+H)=2\alpha(H)\neq 0$ and by Lemma LABEL:sigma, $\sigma(\alpha)\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ In this case (since one can choose between $\sigma(\alpha)$ and $-\sigma(\alpha)$) we may assume that $\sigma(\alpha)(H_{\beta})<0.$
+Therefore in any case there exists $\alpha\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that $\alpha(H_{\beta})<0.$ But for any root vector $x_{\alpha}\in Z+\tau(Z)$ we have $\operatorname{ad}_{H_{\beta}}(x_{\alpha})=\alpha(H_{\beta})x_{\alpha}$ and so for $x_{\alpha}\in\mathfrak{g}^{x_{\beta}},$ $j=\alpha(H_{\beta})$.
 This is a contradiction with the inequality $j\geq 0$.
 The proof of the Lemma is complete.
+Note that Lemma 2 is tacitly used here.
 ∎
 4 Proof of Theorem 2: eliminating the regular case
-We have the following setting
-  • $\mathfrak{g}$ is an absolutely simple Lie algebra
-  • $\mathfrak{h},\mathfrak{l}$ are semisimple (and consist of simple ideals of non-compact type) subalgebras of $\mathfrak{g}$ s.t.
-the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ gives a standard compact Clifford-Klein form
-  • The triple is not contained in Table 1 and thus there exists $0\neq X\in\mathfrak{m}_{0}$ orthogonal to $\mathfrak{h}\oplus\mathfrak{l}$.
+4.1 Assumptions
+In this section we consider the following triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$:
+  1. 1.
+    $\mathfrak{g}$ is an absolutely simple Lie algebra,
+  2. 2.
+    $\mathfrak{h},\mathfrak{l}$ are absolutely simple (of non-compact type) subalgebras of $\mathfrak{g}$ and the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ gives a standard compact Clifford-Klein form,
+  3. 3.
+    the triple is not contained in Table 1 and thus there exists $0\neq X\in\mathfrak{m}_{0}$ orthogonal to $\mathfrak{h}+\mathfrak{l}$,
+  4. 4.
     $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c},$ that is, some Cartan subalgebra of $\mathfrak{g}^{c}$ normalizes $\mathfrak{h}^{c}.$
-Denote by $\mathfrak{d}^{c}$ the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}^{c}.$ We have
-$\mathfrak{d}^{c}=\mathfrak{m}_{0}^{c}\oplus(\mathfrak{a}_{h}^{\perp})^{c}\oplus(\Sigma_{\alpha\in\Delta_{0}^{\pm},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha})\oplus\mathfrak{a}_{h}^{c},$
-where
-$\mathfrak{m}_{0}^{c}=\mathfrak{t}^{c}\oplus\Sigma_{\alpha\in\Delta_{m}}\mathfrak{g}_{\alpha}.$
-Notice that $\mathfrak{d}^{c}$ is a reductive subalgebra.
-Also if $\alpha\in\Delta_{0}^{\pm}$ and $\alpha|_{\mathfrak{a}_{h}}=0$ then (by Theorem 4) necessarily $\alpha(iX)=0.$ Consider the semisimple subalgebra
+In this section we prove the "regular" part of Theorem 2, that is, the following.
+Theorem 7.
+Assume that neither $G/H$ nor $H$ is compact. Under the assumptions of this section the pair $(\mathfrak{g},\mathfrak{h})$ cannot admit compact standard Clifford-Klein forms.
+4.2 Plan of the proof
+The plan of the proof is as follows.
+  1. 1.
+    The assumptions that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ induces a compact standard Clifford-Klein form and that there exists $0\neq X\not\in\mathfrak{h}+\mathfrak{l}$ yield the conclusions of Theorem 4, that is, some particular properties of the spaces $Z_{h}$ and $Z_{l}$ with bases $S_{h}^{i}$ and $S_{l}^{i}$ (statement 1), the relation between $S_{h}^{i}$ and $\Sigma_{h}$ ( statement 2), and the form of the base of the complexified real root spaces $\mathfrak{h}_{\gamma}^{c}$ (statement 3).
+  2. 2.
+    The assumption of the regularity of $\mathfrak{h}^{c}$ together with the assumption $X\not\in\mathfrak{h}+\mathfrak{l}$ yields the existence of a Cartan subalgebra $\mathfrak{c}\subset\mathfrak{g}^{c}$ such that $\mathfrak{a}_{h}\subset\mathfrak{c}$ and an element $0\not=Y\in\mathfrak{c}$ such that $Y\not\in\mathfrak{m}_{0}$, but $Y\in\mathfrak{z}_{\mathfrak{g}}(\mathfrak{h})$ and $[Y,iX]=0$.
+  3. 3.
+    Statements 1) and 2) are contradictory, and, therefore, $\mathfrak{g}=\mathfrak{h}+\mathfrak{l}$, as required.
+4.3 Deriving the contradiction between 1) and 2)
+Let us show first, how to derive the required contradiction. Put
+$Z_{p}=\sum_{\alpha\in\Delta_{p}^{\pm}}\mathfrak{g}_{\alpha},Z_{n}=\sum_{\alpha\in\Delta_{n}^{\pm}}\mathfrak{g}_{\alpha},O=\sum_{\alpha\Delta_{0}}\mathfrak{g}_{\alpha}.$
+By 2),
+$[Y,iX]=0,[Y,\mathfrak{a}_{h}]=0.$ (3)
+This implies
+$[Y,Z_{n}]\subset Z_{n},[Y,Z_{p}]\subset Z_{p},[Y,O]\not\subset(Z_{p}+Z_{n})\setminus\{0\}.$
+Indeed, $\mathfrak{g}_{\alpha}\subset Z_{p}$ if and only if $\alpha(iX)>0$ and $\alpha(A^{+})>0$ for any $A^{+}$ from the interior of the positive Weyl chamber in $\mathfrak{a}_{h}$ (we consider Weyl chambers in $\mathfrak{a}_{h}$).
+Let $U\in\mathfrak{g}_{\alpha}\subset Z_{p}$.
+The Jacobi identity and (3) imply
+$[iX,[Y,U]]=\alpha(iX)[Y,U],\,\text{and}\,\,[iX,[Y,A^{+}]]=\alpha(A^{+})[Y,U].$
+Thus, $[Y,U]$ is contained in a sum of root spaces determined by roots which are positive on $iX$ and on $A^{+}$, as required. Similar arguments work for $Z_{n}$ and $O$.
+Since $[Y,\mathfrak{h}^{c}]=0$, by Theorem 4 (statements 2 and 3), $[Y,S_{h}^{i}+Q_{i}]=0$ for all $i=1,...,k$. Theorem 4 (statement 1) implies
+$[Y,x_{\alpha_{i}}]=0,\,\,\text{for all}\,\,\alpha_{i}\in\Delta_{p}^{\pm}.$
+It follows that $[Y,\sigma(x_{\alpha_{i}})]=0$ as well, and, therefore, $[Y,x_{\alpha_{j}}]=0$ for any $\alpha_{j}\in\Delta^{\pm}_{n}$.
+Lemma 4 yields $[Y,\mathfrak{a}]=0$ which contradicts 2), that is, the assumption $Y\not\in\mathfrak{m}_{0}$.
+4.4 Constructing $Y\in\mathfrak{c},Y\not\in\mathfrak{m}_{0},Y\in\mathfrak{z}_{\mathfrak{g}}(\mathfrak{h}),[Y,iX]=0$
+We will construct the required $Y$ in the centralizer $\mathfrak{z}_{\mathfrak{g}^{c}}(\mathfrak{a}_{h})$.
+4.4.1 The structure of $\mathfrak{z}_{\mathfrak{g}^{c}}(\mathfrak{a}_{h})$
+One easily sees that
+$\mathfrak{z}_{\mathfrak{g}^{c}}(\mathfrak{a}_{h})=\mathfrak{d}^{c}=\mathfrak{m}_{0}+(\mathfrak{a}_{h}^{\perp})^{c}+\mathfrak{a}_{h}+\sum_{\alpha\in\Delta_{0},\alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha},$
+where $(\mathfrak{a}_{h}^{\perp})^{c}$ is the complexification of the orthogonal complement to $\mathfrak{a}_{h}$ in $\mathfrak{a}$.
+This is a reductive subalgebra in $\mathfrak{g}^{c}$.
+Since $\mathfrak{h}^{c}$ is regular, there exists a Cartan subalgebra $\mathfrak{c}\subset\mathfrak{g}^{c}$ such that $[\mathfrak{c},\mathfrak{h}^{c}]\subset\mathfrak{h}^{c}.$ Clearly, $\mathfrak{h}^{c}\cap\mathfrak{c}$ is a Cartan subalgebra in $\mathfrak{h}^{c}$.
+Since any two Cartan subalgebras in $\mathfrak{h}^{c}$ are conjugate by elements $\operatorname{Ad}(h)$ (where $h$ is an element in the connected subgroup $H^{c}\subset G^{c}$ corresponding to $\mathfrak{h}^{c}$), any Cartan subalgebra of $\mathfrak{h}^{c}$ is of the form $\operatorname{Ad}(h)(\mathfrak{c})\cap\mathfrak{h}^{c}$.
+It easily follows that without loss of generality one may assume that $\mathfrak{a}_{h}^{c}\subset\mathfrak{c}$.
+Since $\mathfrak{d}^{c}$ centralizes $\mathfrak{a}_{h}$ one finally obtains
+$\mathfrak{a}_{h}^{c}\subset\mathfrak{c}\subset\mathfrak{d}^{c}.$ (4)
+Lemma 5.
+The Cartan subalgebra $\mathfrak{c}$ can be chosen in a way to satisfy $(4)$ and to be $\sigma$- and $\theta$-invariant.
+Proof.
+Clearly, $\sigma(\mathfrak{a}_{h})=\mathfrak{a}_{h}$.
+It follows that $\sigma(\mathfrak{d}^{c})=\mathfrak{d}^{c}.$ Denote by $\mathfrak{d}^{\sigma}$ the real form of $\mathfrak{d}^{c}$ given by $\sigma.$ We see that $\mathfrak{d}^{\sigma}$ is the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}.$ Put $\mathfrak{c}_{h}:=\mathfrak{c}\cap\mathfrak{h}^{c}.$ After conjugating $\mathfrak{c}$ by an element of the centralizer of $\mathfrak{a}_{h}$ in $H^{c}$ we get $\sigma(\mathfrak{c}_{h})=\mathfrak{c}_{h}.$ Therefore $\mathfrak{c}_{h}^{\sigma}$ (the real form of $\mathfrak{c}_{h}$ with respect to $\sigma$) is in $\mathfrak{d}^{\sigma}.$ Take a $\sigma$-stable Cartan subalgebra $\mathfrak{c}^{\prime}$ of the normalizer of $\mathfrak{h}^{c}$ in $\mathfrak{d}^{c}$ containing $\mathfrak{c}_{h}^{\sigma}.$ One checks that $[\mathfrak{c}^{\prime},\mathfrak{h}^{c}]\subset\mathfrak{h}^{c}$ and $\sigma(\mathfrak{c}^{\prime})\subset\mathfrak{c}^{\prime},$ as desired.
+By a similar argument (conjugating $(\mathfrak{c}^{\prime})^{\sigma}$ inside $G$) we get $\theta(\mathfrak{c}^{\prime})=\mathfrak{c}^{\prime}.$
+Lemma 6.
+The following equality is satisfied:
+$\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}=\{0\}.$
+Proof.
+Let $Y\in\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}.$ For every $1\leq i\leq k$ we get the following:
+$[Y,S_{h}^{i}+Q_{i}+\sigma(S_{h}^{i}+Q_{i})]=0,$
+$[Y,i(S_{h}^{i}+Q_{i}-\sigma(S_{h}^{i}+Q_{i}))]=0,$
+This easily follows from the fact that for any $i=1,...,k$, vectors
+$S_{h}^{i}+Q_{i}+\sigma(S_{h}^{i}+Q_{i})\,\,\text{and}\,\,i(S_{h}^{i}+Q_{i}-\sigma(S_{h}^{i}+Q_{i})$
+belong to $\mathfrak{h}$, and the properties of the Killing form.
+Indeed, one can notice that $[Y,h]\perp\mathfrak{h}$ for any $h\in\mathfrak{h}$.
+On the other hand, $[Y,h]\in\mathfrak{h}$ since $Y$ normalizes $\mathfrak{h}$.
+Now observe that since $i$ takes all values from $1$ to $k$, $Y$ normalizes all root vectors of $\Delta_{p}^{\pm}$ and $\Delta_{n}^{\pm}.$ By Lemma 4 every root of $\Delta_{0}^{\pm}$ is of the form $\alpha_{1}+\alpha_{2}$ for some $\alpha_{1}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and $\alpha_{2}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}\cup\Delta_{m}.$ This implies that $Y$ centralizes $\mathfrak{n}$ and so $Y=0.$
+Consider the semisimple part of the reductive subalgebra $\mathfrak{d}^{c}$, that is
 $\mathfrak{d}_{s}^{c}:=[\mathfrak{d}^{c},\mathfrak{d}^{c}].$
+Define
+$\Phi=\{\alpha\in\Delta_{0}\,|\,\alpha|_{\mathfrak{a}_{h}}=0\}.$
 Lemma 7.
-The subalgebra $\mathfrak{d}^{c}_{s}$ is not simple or $\Phi:=\{\alpha\in\Delta_{0}^{\pm}\ |\ \alpha|_{\mathfrak{a}_{h}}=0\}=\emptyset.$
+At least one of the following is valid:
+  1. 1.
+    $\mathfrak{d}^{c}_{s}$ is not simple,
+  2. 2.
+    $iX$ is in the center of $\mathfrak{d}^{c}.$
 Proof.
-Assume that $\Phi\neq\emptyset.$ If $\alpha\in\Delta_{m}$ is s.t.
-$\alpha(iX)\neq 0$ and $\beta\in\Phi$ then $\alpha\pm\beta$ is not a root since otherwise $\alpha\pm\beta\in\Phi$ and $(\alpha\pm\beta)(iX)\neq 0,$ which is impossible.
-Take any $\alpha_{1}\in\Phi.$ It follows that for any $\alpha\in\Delta_{m},$ $\alpha(iX)\neq 0$ we have $\alpha\perp\alpha_{1}.$ But now every root of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ is orthogonal to $iX$ or to $\alpha_{1}$ and so all roots are contained in two hyperplanes.
+If $iX$ is in the center of $\mathfrak{d}^{c}$ then the proof is complete.
+So assume that $\alpha(iX)\neq 0$ for some $\alpha\in\Delta_{m}.$ By Lemma 6, $\Phi\not=\emptyset$.
+If $\beta\in\Phi$ then $\alpha\pm\beta$ is not a root since otherwise $\alpha\pm\beta\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and $(\alpha\pm\beta)|_{\mathfrak{a}_{h}}=0,$ which is impossible (by Theorem 4).
+Take any $\alpha_{1}\in\Phi.$ It follows that for any $\gamma\in\Delta_{m},$ $\gamma(iX)\neq 0$ we have $\gamma\perp\alpha_{1}.$ But now every root of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ is orthogonal to $iX$ or to $\alpha_{1}$ and so all roots are contained in two hyperplanes (but not all roots are in the first one, nor in the second one).
 Lemma 3 implies that $\mathfrak{d}_{s}^{c}$ is not simple.
@@
 Corollary 2.
-We can decompose $\mathfrak{d}^{c}$ into a sum of three ideals $\mathfrak{d}^{c}=\mathfrak{d}_{1}\oplus\mathfrak{d}_{2}\oplus\mathfrak{a}_{h}^{c}$ such that $iX\in\mathfrak{d}_{1}$ and $\mathfrak{a}_{h}^{\perp},\ (\Sigma_{\alpha\in\Delta_{0}^{\pm},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha})\subset\mathfrak{d}_{2}.$
+Lie algebra $\mathfrak{d}^{c}$ decomposes into a direct sum of Lie algebras
+$\mathfrak{d}^{c}=\mathfrak{d}_{1}\oplus\mathfrak{d}_{2}\oplus\mathfrak{a}_{h}^{c}$
+such that
+  1. 1.
+    $iX\in\mathfrak{d}_{1}$
+  2. 2.
+    $\mathfrak{a}_{h}^{\perp}\oplus\Sigma_{\alpha\in\Delta_{0},\ \alpha|_{\mathfrak{a}_{h}}=0}\mathfrak{g}_{\alpha}\subset\mathfrak{d}_{2}.$
+  3. 3.
+    $\sigma(\mathfrak{d}_{2})\subset\mathfrak{d}_{2},\theta(\mathfrak{d}_{2})\subset\mathfrak{d}_{2}$.
 Proof.
-First notice that $\mathfrak{a}_{h}^{c}$ is in the center of $\mathfrak{d}^{c}.$ If $\Phi=\emptyset$ then take $\mathfrak{d}_{1}:=\mathfrak{m}_{0}^{c},$ $\mathfrak{d}_{2}:=(\mathfrak{a}_{h}^{\perp})^{c}.$
-Assume that $\alpha_{1}\in\Phi\neq\emptyset.$ All roots of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ lie on two hypersurfaces in $(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c},$ $s_{iX}$ (orthogonal to $iX$) and $s_{\alpha_{1}}$ (orthogonal to $\alpha_{1}$).
-We see that if $\tilde{\mathfrak{d}}$ is a simple ideal of $\mathfrak{d}_{s}^{c}$ then (by Lemma 5) all roots of $(\tilde{\mathfrak{d}},\tilde{\mathfrak{d}}\cap(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c}))$ lie either in $s_{iX},$ or in $s_{\alpha_{1}}.$ Therefore we may take $\mathfrak{d}_{2}$ the be the sum of all simple ideals which roots lie in $s_{iX}$ plus $(\mathfrak{a}_{h}^{\perp})^{c}.$ Then $\mathfrak{d}_{1}$ is the ideal of $\mathfrak{m}_{0}^{c}$ complementary to $\mathfrak{d}_{2}\cap\mathfrak{m}_{0}^{c}.$
-As $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c}$ then some Cartan subalgebra $\mathfrak{c}$ of $\mathfrak{g}^{c}$ normalizes $\mathfrak{h}^{c}.$ We may assume that $\mathfrak{a}_{h}^{c}\subset\mathfrak{c}.$ Indeed, notice that $\mathfrak{h}^{c}\cap\mathfrak{c}$ is a Cartan subalgebra of $\mathfrak{h}^{c}.$ Since any two Cartan subalgebras of $\mathfrak{h}^{c}$ are conjugate by an element of $H^{c}$ thus any Cartan subalgebra of $\mathfrak{h}^{c}$ is of the form $Ad_{h}(\mathfrak{c})\cap\mathfrak{h}^{c}$ for some $h\in H^{c}$ (where $H^{c}$ denotes a connected subgroup of $G^{c}$ corresponding to $\mathfrak{h}^{c}$).
-Obviously $Ad_{h}(\mathfrak{c})$ normalizes $\mathfrak{h}^{c}.$ As $\mathfrak{a}_{h}^{c}$ is in some Cartan subalgebra of $\mathfrak{h}^{c}$ therefore we may assume that
-$\mathfrak{a}_{h}^{c}\subset\mathfrak{c}\subset\mathfrak{d}^{c}.$
-Lemma 9.
-We may assume that
-$\sigma(\mathfrak{c})=\mathfrak{c},\ \ \theta(\mathfrak{c})=\mathfrak{c},\ \ \sigma(\mathfrak{d}_{2})=\mathfrak{d}_{2},\ \ \theta(\mathfrak{d}_{2})=\mathfrak{d}_{2}.$
-Proof.
-First notice that $\sigma(\mathfrak{a}_{h})=\mathfrak{a}_{h}$ and so $\sigma(\mathfrak{d}^{c})=\mathfrak{d}^{c}.$ Denote by $\mathfrak{d}^{\sigma}$ the real form of $\mathfrak{d}^{c}$ given by $\sigma.$ We see that $\mathfrak{d}^{\sigma}$ is the centralizer of $\mathfrak{a}_{h}$ in $\mathfrak{g}.$ Put $\mathfrak{c}_{h}:=\mathfrak{c}\cap\mathfrak{h}^{c}.$ After conjugating $\mathfrak{c}$ by an element of the centralizer of $\mathfrak{a}_{h}$ in $H^{c}$ we have $\sigma(\mathfrak{c}_{h})=\mathfrak{c}_{h}.$ Therefore $\mathfrak{c}_{h}^{\sigma}$ (the real form of $\mathfrak{c}_{h}$ w.r.t.
-$\sigma$) is in $\mathfrak{d}^{\sigma}.$ We take a $\sigma$-stable Cartan subalgebra $\mathfrak{c}^{\prime}$ of the normalizer of $\mathfrak{h}^{c}$ in $\mathfrak{d}^{c}$ containing $\mathfrak{c}_{h}^{\sigma}.$ We have $[\mathfrak{c}^{\prime},\mathfrak{h}^{c}]\subset\mathfrak{h}^{c}$ and $\sigma(\mathfrak{c}^{\prime})\subset\mathfrak{c}^{\prime},$ as desired.
-By the similar argument (by conjugating $(\mathfrak{c}^{\prime})^{\sigma}$ inside $G$) we have $\theta(\mathfrak{c}^{\prime})=\mathfrak{c}^{\prime}.$
-Notice that as $\sigma(X)=\theta(X)=X$ we have $\sigma(\mathfrak{d}_{2})=\theta(\mathfrak{d}_{2})=\mathfrak{d}_{2}.$
-Lemma 10.
-We have $\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}=\{0\}$ and so $\Phi\neq\emptyset.$
-Proof.
-Let $Y\in\mathfrak{a}_{h}^{\perp}\cap\mathfrak{c}^{\sigma}.$ For every $1\leq i\leq k$ we have
-$[Y,S_{h}^{i}+Q_{i}+\sigma(S_{h}^{i}+Q_{i})]\subset\mathfrak{n}_{h}\cap\mathfrak{n}_{h}^{\perp}=\{0\},$
-$[Y,i(S_{h}^{i}+Q_{i}-\sigma(S_{h}^{i}+Q_{i}))]\subset\mathfrak{n}_{h}\cap\mathfrak{n}_{h}^{\perp}=\{0\},$
-where $\mathfrak{n}_{h}^{\perp}$ is the orthogonal complement of $\mathfrak{n}_{h}$ in $\mathfrak{n}$ w.r.t.
-$B_{\theta}(\bullet,\bullet)=B(\bullet,\theta(\bullet)).$ It follows that $Y$ centralizes all roots of $\Delta_{p}^{\pm}$ and $\Delta_{n}^{\pm}.$ By Lemma 7 every root of $\Delta_{0}^{\pm}$ is of the form $\alpha_{1}+\alpha_{2}$ for some $\alpha_{1},\alpha_{2}\in\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ But now $Y$ centralizes $\mathfrak{n}$ and so $Y=0.$
-Notice that
-$\mathfrak{d}^{\sigma}\subset\mathfrak{k}_{d}\oplus\mathfrak{m}_{0}\oplus\mathfrak{p}_{d}\oplus\mathfrak{a}_{h},$
-$\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma},$
-$\mathfrak{p}_{d}\subset\mathfrak{d}_{2}^{\sigma},$
-where
-$\mathfrak{p}_{d}=\mathfrak{a}_{h}^{\perp}\oplus\textrm{Lin}_{\mathbb{R}}\{x_{\alpha}+\sigma(x_{\alpha})-\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))-\theta(i(x_{\alpha}+\sigma(x_{\alpha})))\ |\ \alpha\in\Phi\},$
-$\mathfrak{k}_{d}=\textrm{Lin}_{\mathbb{R}}\{x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))+\theta(i(x_{\alpha}+\sigma(x_{\alpha})))\ |\ \alpha\in\Phi\}.$
-Previous Lemmas and the fact that $\textrm{rank}(\mathfrak{g})>\textrm{rank}(\mathfrak{m}_{0}\oplus\mathfrak{a}_{h})$ imply the following corollary.
+Note that $\mathfrak{a}_{h}^{c}$ is in the center of $\mathfrak{d}^{c}.$ If $iX\in\mathfrak{z}(\mathfrak{d}^{c})$, then put $\mathfrak{d}_{1}:=\textrm{Span}_{\mathbb{C}}(iX)$ and take the ideal $\mathfrak{d}_{2}$ complementary to $\textrm{Span}_{\mathbb{C}}(iX)\oplus\mathfrak{a}_{h}^{c}.$ Otherwise, assume that $\alpha_{1}\in\Phi\neq\emptyset$ and $\alpha(iX)\neq 0$ for some $\alpha\in\Delta_{m}.$ Define hyperplanes $s_{iX}$ and $s_{\alpha_{1}}$ in $(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c},$ orthogonal to $iX$ and $\alpha_{1},$ respectively.
+All roots of $(\mathfrak{d}_{s}^{c},(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c})\cap\mathfrak{d}_{s}^{c})$ lie in these two hyperplanes.
+We see that if $\tilde{\mathfrak{d}}$ is a simple ideal of $\mathfrak{d}_{s}^{c}$, then (by Lemma 3) all roots of $(\tilde{\mathfrak{d}},\tilde{\mathfrak{d}}\cap(\mathfrak{t}^{c}\oplus\mathfrak{a}^{c}))$ lie either in $s_{iX},$ or in $s_{\alpha_{1}}.$ Therefore we may take $\mathfrak{d}_{2}$ to be the sum of all simple ideals whose roots lie in $s_{iX}+(\mathfrak{a}_{h}^{\perp})^{c}.$ Then $\mathfrak{d}_{1}$ is the ideal of $\mathfrak{m}_{0}^{c}$ complementary to $\mathfrak{d}_{2}\cap\mathfrak{m}_{0}^{c}.$
+Finally, the third part of the Corollary follows, since $\sigma(X)=X$ and $\theta(X)=X$. ∎
+Let $\mathfrak{k}_{d}$ denote the subalgebra of $\mathfrak{g}$ generated by vectors
+$x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))+\theta(i(x_{\alpha}-\sigma(x_{\alpha}))),\forall\alpha\in\Phi,$
+and $\mathfrak{p}_{d}$ denote the subspace in $\mathfrak{g}$ generated by $\mathfrak{a}_{h}^{\perp}$ and all vectors of the form
+$x_{\alpha}+\sigma(x_{\alpha})-\theta(x_{\alpha}+\sigma(x_{\alpha})),\ i(x_{\alpha}-\sigma(x_{\alpha}))-\theta(i(x_{\alpha}-\sigma(x_{\alpha}))),\forall\alpha\in\Phi.$
+From the very definition of $\mathfrak{k}_{d}$ and $\mathfrak{p}_{d}$, as well as the definition of $\mathfrak{d}^{c}$ and taking into consideration Corollary 2 one obtains the inclusions
+$\mathfrak{k}_{d}\oplus\mathfrak{p}_{d}\subset\mathfrak{d}_{2}^{\sigma}\subset\mathfrak{d}^{\sigma}\subset\mathfrak{k}_{d}\oplus\mathfrak{p}_{d}\oplus\mathfrak{m}_{0}\oplus\mathfrak{a}_{h}.$ (5)
 Corollary 3.
-We have
+The following is valid
 $\mathfrak{c}^{\sigma}\cap(\mathfrak{p}_{d}\oplus\mathfrak{a}_{h})=\mathfrak{a}_{h}.$
 Proof.
-Obviously $\mathfrak{a}_{h}\subset\mathfrak{c}^{\sigma}.$ Take any $Y\in\mathfrak{p}_{d}\cap\mathfrak{c}_{\sigma}.$ Then $Y$ is conjugate by an element $k^{-1}$ of $K_{d}$ (connected closed subgroup of $K$ corresponding to reductive subalgebra $\mathfrak{k}_{d}$) to $\mathfrak{a}_{h}^{\perp}.$ As $\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma}$ thus it centralizes $iX$ and $\mathfrak{a}_{h}$ and so for any $g\in K_{d}$ $Ad_{g}\mathfrak{h}$ is well embedded (as $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$).
-Also $\mathfrak{a}_{h}^{\perp}$ is orthogonal to $\textrm{Ad}_{g}\mathfrak{h}.$ We can repeat the argument of the proof of Lemma 10 to $\textrm{Ad}_{k}Y$ and $\textrm{Ad}_{k}\mathfrak{h}.$
-Because $\theta(\mathfrak{d}_{2}^{\sigma})=\mathfrak{d}_{2}^{\sigma}$ therefore $\mathfrak{d}_{2}^{\sigma}\subset\mathfrak{k}.$ As every element of a Lie algebra of compact type is in some Cartan subalgebra we may conjugate $\mathfrak{c}^{\sigma}$ by an element of $\mathfrak{d}_{2}^{\sigma}$ so that
-$0\neq\tilde{Y}:=x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha}))\in\mathfrak{c}^{\sigma},$
-where $\alpha\in\Phi.$ As $\mathfrak{d}_{2}^{\sigma}$ centralizes $iX$ and $\mathfrak{a}_{h}$ thus for any $g\in D_{2}^{\sigma}$ (connected closed subgroup of $G$ corresponding to reductive subalgebra $\mathfrak{d}_{2}^{\sigma}$) $Ad_{g}\mathfrak{h}$ is well embedded (as $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$).
-Thus Theorem 4 holds for $\textrm{Ad}_{g}\mathfrak{h}.$
+It is straightforward to see that $\mathfrak{a}_{h}\subset\mathfrak{c}^{\sigma}.$ Take any $Y\in\mathfrak{p}_{d}\cap\mathfrak{c}^{\sigma}.$ Let $K_{d}$ be a connected closed (reductive) subgroup of $K$ corresponding to $\mathfrak{k}_{d}$.
+Since $\mathfrak{a}_{h}^{\perp}$ is maximal abelian in $\mathfrak{p}_{d}$, there is an element $k\in K_{d}$ such that $\operatorname{Ad}_{k}(Y)\in\mathfrak{a}^{\perp}_{h}$.
+Since $\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma}$, it centralizes $iX$ and $\mathfrak{a}_{h}$ and so for any $g\in K_{d}$ the subalgebra $\operatorname{Ad}_{g}\mathfrak{h}$ is well embedded (since $\operatorname{Ad}_{g}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\operatorname{Ad}_{g}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\operatorname{Ad}_{g}\mathfrak{h})=\operatorname{Ad}_{g}\mathfrak{h}$).
+Also $\mathfrak{a}_{h}^{\perp}$ is orthogonal to $\operatorname{Ad}_{g}\mathfrak{h}.$ Now, one can repeat the argument of the proof of Lemma 6 (taking the appropriate Lie brackets) applying it to $\operatorname{Ad}_{k}Y$ and $\operatorname{Ad}_{k}\mathfrak{h}.$
+4.4.2 Construction of $Y$
+Let $D_{2}^{\sigma}$ denote the connected closed subgroup of $G$ corresponding to $\mathfrak{d}_{2}^{\sigma}$.
 Lemma 8.
-We have $\tilde{Y}\notin\mathfrak{h}.$
+For any $g\in D_{2}^{\sigma}$ the embedding $\operatorname{Ad}_{g}(\mathfrak{h})\hookrightarrow\mathfrak{g}$ is good. Thus, Theorem 4 can be applied to this embedding.
 Proof.
-Since $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}$ therefore there exist $H\in\mathfrak{n}_{h}$ and $L\in\mathfrak{n}_{l}$ s.t.
-$x_{\alpha}+\sigma(x_{\alpha})=H+L$ and so $\tilde{Y}=H+\theta(H)+L+\theta(L).$ If $\tilde{Y}\in\mathfrak{h}$ then $L+\theta(L)\in\mathfrak{h}.$ But
+Since $\mathfrak{d}_{2}^{\sigma}$ centralizes $iX$ and $\mathfrak{a}_{h}$ one derives that for any $g$ in the maximal compact subgroup of $D_{2}^{\sigma}$ the subalgebra $Ad_{g}\mathfrak{h}$ is well embedded (because $\textrm{Ad}\mathfrak{n}_{h}^{\pm}\subset\mathfrak{n}^{\pm},$ $\textrm{Ad}\mathfrak{a}_{h}=\mathfrak{a}_{h}$ and $\theta(\textrm{Ad}_{g}\mathfrak{h})=\textrm{Ad}_{g}\mathfrak{h}$).
+∎
+Lemma 9.
+For any $\alpha\in\Phi$ vector
+$0\neq\tilde{Y}:=x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha}))$
+does not belong to $\mathfrak{h}$, and for any $g\in D_{2}^{\sigma}$ vector $Y=\operatorname{Ad}_{g}(\tilde{Y})$ does not belong to $\operatorname{Ad}_{g}(\mathfrak{h})$.
+Proof.
+Since $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}$, there exist $H\in\mathfrak{n}_{h}$ and $L\in\mathfrak{n}_{l}$ such that $x_{\alpha}+\sigma(x_{\alpha})=H+L$ and so $\tilde{Y}=H+\theta(H)+L+\theta(L).$ If $\tilde{Y}\in\mathfrak{h}$ then $L+\theta(L)\in\mathfrak{h}.$ But
 $\{W+\theta(W)\ |\ W\in\mathfrak{n}\}=\{W+\theta(W)\ |\ W\in\mathfrak{n}_{h}\}\oplus\{W+\theta(W)\ |\ W\in\mathfrak{n}_{l}\}$
 and thus $L+\theta(L)=H_{1}+\theta(H_{1})$ for some $H_{1}\in\mathfrak{n}_{h}.$ As $\mathfrak{n}\cap\theta(\mathfrak{n})=\{0\}$ therefore $L\in\mathfrak{n}_{h}$ and so $L=0.$ Thus $x_{\alpha}+\sigma(x_{\alpha})\in\mathfrak{n}_{h}$ and $\alpha|_{\mathfrak{a}_{h}}=0,$ a contradiction.
+The second claim is clear.
 ∎
-It follows that there exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$ which is orthogonal (with respect to $B$) to $\mathfrak{h}.$ Since $\mathfrak{h}^{c}$ is regular we obtain the following
-Corollary 4.
-There exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$ which centralizes $\mathfrak{h}.$
-Theorem 8.
-Assume that neither $G/H$ nor $H$ is compact.
-In the setting of this section (there exists $iX\neq 0$ and $\mathfrak{h}^{c}$ is regular in $\mathfrak{g}^{c}$) the pair $(\mathfrak{g},\mathfrak{h})$ cannot induce standard Clifford-Klein form.
+Now we prove the main proposition which yields $Y$ as in the title of Subsection 4.4.
+Proposition 2.
+There exists a non-zero $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma},$ $Y\notin\mathfrak{m}_{0}$ which centralizes $\mathfrak{h}.$
 Proof.
-Put
-$Z_{p}:=\sum_{\alpha\in\Delta_{p}^{\pm}}\mathfrak{g}_{\alpha},\ \ Z_{n}:=\sum_{\alpha\in\Delta_{n}^{\pm}}\mathfrak{g}_{\alpha},\ \ O:=\sum_{\alpha\in\Delta_{0}^{\pm}}\mathfrak{g}_{\alpha}.$
-As $[Y,iX]=[Y,\mathfrak{a}_{h}]=0$ therefore
-$[Y,Z_{n}]\subset Z_{n},\ \ [Y,Z_{p}]\subset Z_{p},\ \ [Y,O]\subset O.$
-But $Y$ centralizes $\mathfrak{h}^{c}$ and so for $1\leq i\leq k$ we have $[Y,S_{h}^{i}+O_{i}]=0.$ By Theorem 4 (1.) we have
-$[Y,x_{\alpha_{i}}]=0,\ \ \alpha_{i}\in\Delta_{p}^{\pm}$
-and thus (as $[Y,\sigma(x_{\alpha_{i}})]=0$) $[Y,x_{\alpha_{j}}]=0$ for any $\alpha_{j}\in\Delta_{n}^{\pm}.$ By Lemma 7 this implies $[Y,x_{\beta_{i}}]=0$ for every $\beta_{i}\in\Delta_{0}^{\pm}.$ Thus
-$[Y,\mathfrak{n}]=[Y,\theta(\mathfrak{n})]=0$
-and so $[Y,\mathfrak{a}]=0$ as $\mathfrak{a}\subset[\mathfrak{n},\theta(\mathfrak{n})].$ But now $[Y,\mathfrak{p}]=0$ and so $[Y,\mathfrak{k}]=0$ since $\mathfrak{k}=[\mathfrak{p},\mathfrak{p}].$ Therefore $[Y,\mathfrak{g}]=0$ and $Y\neq 0.$ But now $\mathfrak{g}$ cannot be simple, a contradiction.
-∎
+For the convenience of the reader we first summarize facts needed in the course of the proof.
+  1. 1.
+    For any semisimple $\mathfrak{g}$ with a Cartan involution $\theta$ and any $\theta$-invariant Cartan subalgebra $\mathfrak{b}$ one has the decomposition
+    $\mathfrak{b}=\mathfrak{b}_{+}\oplus\mathfrak{b}_{-}.$
+    By an abuse of language we will call $\mathfrak{b}_{+}$ the compact part of $\mathfrak{b}$ and $\mathfrak{b}_{-}$ non-compact part.
+  2. 2.
+    By Lemma 5 $\mathfrak{c}^{\sigma}$ is a $\theta$-invariant subalgebra of $\mathfrak{d}^{\sigma}$ and, hence,
+    $\mathfrak{c}^{\sigma}=\mathfrak{c}^{\sigma}_{+}\oplus\mathfrak{c}^{\sigma}_{-}$
+    and by Corollary 2
+    $\mathfrak{c}_{-}^{\sigma}=\mathfrak{a}_{h},\mathfrak{c}^{\sigma}=\mathfrak{c}_{+}^{\sigma}\oplus\mathfrak{a}_{h}.$
+  3. 3.
+    By Corollary 3
+    $\mathfrak{c}^{\sigma}_{+}=\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}.$
+  4. 4.
+    It is easy to see that any two compact parts of any two $\theta$-invariant Cartan subalgebras (with compact parts of the same dimension) in any semisimple real Lie algebra $\mathfrak{g}$ are conjugate, and, in particular, any two Cartan subalgebras in $\mathfrak{d}_{2}^{\sigma}$ are conjugate by an element $\operatorname{Ad}_{g},g\in D_{2}^{\sigma}$.
+Therefore, any compact part of any Cartan subalgebra in $\mathfrak{d}_{2}^{\sigma}$ is conjugate to $\mathfrak{d}_{2}^{\sigma}\cap\mathfrak{c}^{\sigma}$.
+Now we complete the proof using 1)-4) above. Take
+$\tilde{Y}=x_{\alpha}+\sigma(x_{\alpha})+\theta(x_{\alpha}+\sigma(x_{\alpha}),\alpha\in\Phi.$
+By formula $(5)$, $\tilde{Y}\in\mathfrak{k}_{d}\subset\mathfrak{d}_{2}^{\sigma}$.
+Since $\mathfrak{k}_{d}$ is compact, $\tilde{Y}$ belongs to some compact part of some Cartan subalgebra in $\mathfrak{d}_{2}^{\sigma}$.
+It follows that $\tilde{Y}$ is conjugate to element $Y\in\mathfrak{c}^{\sigma}_{+}=\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$.
+Thus, $Y=\operatorname{Ad}_{g}(\tilde{Y})$ for some $g\in D_{2}^{\sigma}$.
+By Lemma 8 we can replace the pair $(\mathfrak{g},\mathfrak{h})$ by $(\mathfrak{g},\operatorname{Ad}_{g}(\mathfrak{h})$ and $\tilde{Y}$ by $Y\in\mathfrak{c}^{\sigma}\cap\mathfrak{d}_{2}^{\sigma}$.
+By construction, $Y\not\in\mathfrak{m}_{0}$ and by Lemma 9 $Y\not\in\operatorname{Ad}_{g}(\mathfrak{h})$, hence one can also choose $Y\perp\operatorname{Ad}_{g}(\mathfrak{h})$.
+Since $\mathfrak{h}^{c}$ is regular, this is the same as $Y\in\mathfrak{z}_{\mathfrak{g}}(\operatorname{Ad}_{g}(\mathfrak{h}))$.
+Now, by some abuse of notation we write this as in the title.
 5 Completion of proof of Theorem 2: non-regular case
-In view of Theorem 8 we assume that the embeddings $\mathfrak{h}^{c},\mathfrak{l}^{c}\hookrightarrow\mathfrak{g}^{c}$ are not regular. The proof is divided into two parts:
+In view of Theorem 7 we do not assume that the embeddings $\mathfrak{h}^{c},\mathfrak{l}^{c}\hookrightarrow\mathfrak{g}^{c}$ are regular.
+Instead, we assume that $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple.
+The proof is based on a careful consideration of possibilities for the real forms of the embeddings $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$, therefore, we postpone the calculations to the very end of the article.
+In this section we explain the main steps of the proof.
+5.1 Terminology and notation
+We begin by introducing some convenient (for us) terminology.
+Since $\mathfrak{g}^{c}$ is a classical Lie algebra, that is, $\mathfrak{g}^{c}$ can be $\mathfrak{sl}(V),\mathfrak{so}(V)$, or $\mathfrak{sp}(V)$, we get linear representations $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ in the complex vector space $V$ and call them linear if $\mathfrak{g}^{c}=\mathfrak{sl}(V)$, orthogonal, if $\mathfrak{g}^{c}=\mathfrak{so}(V)$ and symplectic, if $\mathfrak{g}^{c}=\mathfrak{sp}(V).$
+We single out tensor representations, that is, the representations of the form
   1. 1.
-    $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple,
+    $\mathfrak{h}^{c}=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{sl}(V_{1}\otimes V_{2})$,
   2. 2.
-    at least one of $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ is semisimple but not simple.
-The proof of the main theorem will follow from Propositions 6, 8, 9 proved in this section.
-We begin with case 1) and consider embeddings of classical Lie algebras $\mathfrak{h}^{c}$ as linear representations, and, following Dynkin [14], divide these into two classes: irreducible and reducible.
-In what follows we use the notation from the paper [14].
-We use tables of inclusions of simple classical Lie algebras from [14], however, we also reproduce parts of them as Tables 3,4,5 in Section 6.
-5.1 Proof in the case when $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple
-5.1.1 Irreducible representations and simple classical subalgebras
-We begin with the classification of maximal simple classical subalgebras in simple Lie algebras (see, [30], Chapter 6, Theorem 3.1).
-Following Dynkin [14] we introduce the following terminology.
-Let $\mathfrak{g}^{c}$ be a classical Lie algebra, that is $\mathfrak{g}^{c}=\mathfrak{sl}(V),\mathfrak{so}(V),\mathfrak{sp}(V)$ where $V$ is a finite-dimensional complex linear space.
-Note that by assumption, $\mathfrak{h}^{c}$ is also classical, so, $\mathfrak{h}^{c}=\mathfrak{sl}(U),\mathfrak{so}(U),\mathfrak{sp}(U)$ for some finite-dimensional complex linear space.
+    $\mathfrak{h}^{c}=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})\hookrightarrow\mathfrak{g^{c}}=\mathfrak{so}(V_{1}\otimes V_{2}$,
+  3. 3.
+    $\mathfrak{h}^{c}=\mathfrak{sp}(V_{1})\oplus\mathfrak{sp}(V_{2})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{sp}(V_{1}\otimes V_{2}).$
+The third type of subalgebras we need has the form
+  1. 1.
+    $\mathfrak{h}^{c}=\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$,
+  2. 2.
+    $\mathfrak{h}^{c}=\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})\hookrightarrow\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C}).$
+We will call such subalgebras symmetric (they are fixed points of involutive automorphisms of $\mathfrak{g}^{c}$, because they are maximal in $\mathfrak{g}^{c}$ and Dynkin’s classification of such subalgebras is up to conjugation).
+Note that we do not consider all possible tensor representations or symmetric subalgebras, but only the listed ones.
+Alternatively we refer to Table 6, where maximal subalgebras of maximal rank are presented.
 Definition 4.
 We say that $\mathfrak{h}^{c}$ is an irreducible/reducible subalgebra in $\mathfrak{g}^{c}$ if the linear representation $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is irreducible/reducible.
-Theorem 9 (Dynkin, [30]).
-The classification of maximal semisimple Lie subalgebras $\mathfrak{h}^{c}$ in simple complex Lie algebras $\mathfrak{g}^{c}$ of classical type falls into the classes:
+5.2 Maximal subalgebras of classical complex simple Lie algebras
+In this subsection we summarize Dynkin’s results on maximal subalgebras in classical Lie algebras in [13] and [33] which we formulate as Theorem 8 and Theorem 9.
+Theorem 8 ([33]).
+All maximal proper subalgebras in classical complex Lie algebras fall into the classes below, up to conjugation in $\mathfrak{g}^{c}$:
+I: Reducible
+  • maximal parabolic (I.1),
+  • symmetric (I.2)
+II: Irreducible
+  • simple (II.1)
+  • tensor representations (II.2)
+Not all irreducible simple subalgebras are maximal, however, we have the following ( $\omega_{r}$ is 00 if the representation is linear, $1$ if it is orthogonal, and $(-1)$ if symplectic).
+Theorem 9 ([13]).
+Irreducible simple subalgebras in classical Lie algebras are maximal, with exceptions given by the following two tables.
+Irreducible triples
+Inclusion 1 Inclusion 2 $\dim$ of the representation
+$C_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n-1+k}{k}$
+$B_{n}\subset A_{2n}$ $A_{2n}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n+1}{k}$
+$D_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n}{k}$
+$A_{n}\subset A_{\frac{(n-1)(n+2)}{2}}$ $A_{\frac{(n-1)(n+2)}{2}}\subset\tilde{\mathfrak{g}}^{c}$ $\frac{1}{8}(n-1)(n+1)(n+2)$
+$A_{n}\subset A_{\frac{n(n+3)}{2}}$ $A_{\frac{n(n+3)}{2}}\subset\tilde{\mathfrak{g}}^{c}$ $\frac{1}{8}(n+1)(n+2)(n+3)$
+$B_{n}\subset D_{n+1}$ $D_{n+1}\subset\tilde{\mathfrak{g}}^{c}$ $\prod_{s=1}^{n-1}\binom{k+2s-1}{k}/\binom{k+s}{k}$
+Table 2:
+Non-maximal irreducible subalgebras
+Type of inclusion dimension $\omega_{r}$
+$A_{n}\subset A_{l}$ $3\binom{n+2}{4}$ 00
+$A_{n}\subset A_{l}$ $3\binom{n+3}{4}$ 00
+$B_{2n+1}\subset\mathfrak{g}^{c}$ $\prod_{s=1}^{2n}{\binom{k+s-1}{k}/\binom{k+s}{k}}$ $(-1)^{n+1)k}$
+Table 3:
+5.3 Steps of proof of Theorem 2
+Proposition 3.
+If $\mathfrak{h}^{c}$ is a simple non-regular subalgebra in $\mathfrak{g}^{c}$ then the inequality
+$\operatorname{rank}\mathfrak{h}^{c}\leq\frac{1}{2}(\operatorname{rank}\mathfrak{g}^{c}+1)$ (6)
+holds for all pairs $(\mathfrak{g}^{c},\mathfrak{h}^{c})$ except the following:
   1. 1.
-    $\mathfrak{h}^{c}$ is reducible, and then a) $\mathfrak{h}^{c}$ is maximal parabolic, b) if $\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C})$; if $\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})$, $0<k<n$;
+    $\mathfrak{h}^{c}=B_{p-1}\subset D_{p}$,
   2. 2.
-    $\mathfrak{h}$ is irreducible and non-simple, and then the following possibilities may occur: $\mathfrak{g}^{c}=\mathfrak{sl}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C}),$ $n=st,s,t\geq 2$, if $\mathfrak{g}^{c}=\mathfrak{so}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate either to $\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})$, where $n=st,s,t\geq 3$, $s,t\not=4$, or to $\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})$, where $n=st,s,t\geq 2$, if $\mathfrak{g}^{c}=\mathfrak{sp}(n,\mathbb{C})$, then $\mathfrak{h}^{c}$ is conjugate to $\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{sp}(t,\mathbb{C})$, where $n=st,s\geq 2,t\geq 3$, or $s=2,t=4$.
+    $\mathfrak{h}^{c}\subset B_{p-1}\subset D_{p}\subset B_{n}$, where $D_{p}\subset B_{n}$ is regular,
   3. 3.
-    $\mathfrak{h}^{c}$ is irreducible and simple.
-We begin with Dynkin’s classification of simple complex $\mathfrak{h}^{c}$ such that the embedding $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is an irreducible representation of $\mathfrak{h}^{c}$ in one of the classical Lie algebras $\mathfrak{g}^{c}=\mathfrak{sl}(V)$,$\mathfrak{so}(V),\mathfrak{sp}(V)$ for some complex vector space $V$.
-It says that essentially, maximal and irreducible classes coincide with several exceptions.
-Theorem 10 (Dynkin, [30]).
-Let $R:\mathfrak{h}\rightarrow\mathfrak{gl}(V)$ be a nontrivial irreducible representation of a simple complex Lie algebra $\mathfrak{h}$.
-If there are no $R(\mathfrak{h})$-invariant bilinear form on $V$, then $R(\mathfrak{h})$ is maximal in $\mathfrak{sl}(V)$.
-If $R(\mathfrak{h})$ preserves some orthogonal or symplectic form, then $R(\mathfrak{h})$ is maximal in $\mathfrak{so}(V)$ or $\mathfrak{sp}(V)$.
-The only exceptions are given in Table 1 of [14] (or Table 7 of [30]).
-Remark 2.
-Note that exception means "non-maximal" in this theorem.
-Theorem 11 (Dynkin, [14], Theorem 2.3).
-The list of all triples of irreducible inclusions $\mathfrak{h}^{c}\subset\tilde{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ of simple Lie algebras $\mathfrak{h}^{c},\tilde{\mathfrak{h}}^{c},\mathfrak{g}^{c}$ is given by Table 5 in [14].
-Comparing Theorems 9 and 10 we get the following corollary.
-Corollary 5.
-Any irreducible simple classical subalgebra $\mathfrak{h}^{c}$ of classical Lie algebra $\mathfrak{g}^{c}$ is either maximal, or is contained in Table 1 of [14] (or, equivalently, in Table 7 of [30]), or in Table 5 in [14].
-If the representation of $\mathfrak{h}^{c}$ in $\mathfrak{g}^{c}$ is fundamental, then it is contained in Tables 1 and 2 of [7].
-Remark 3.
-Reading the formulations in [14] one may get an impression that Table 1 contains all the possibilities for irreducible simple classical subalgebras.
-However Table 5 [14] is necessary, since Table 1 of [14] and tables of fundamental representations are organized in the way which does not cover several possible triples.
-This is because, for example, $\mathfrak{sp}(n,\mathbb{C})$ is an irreducible representation of minimal dimension $2n$, which belongs to the table of symplectic irreducible representations, but also $\mathfrak{sp}(n,\mathbb{C})$ is contained in $\mathfrak{sl}(2n,\mathbb{C})$.
-Such cases are covered by Theorem 11.
-5.1.2 Using Dynkin’s classification
-We summarize the procedure for classification of standard Clifford-Klein forms based on the previous results as follows.
+    $\mathfrak{h}^{c}\subset B_{p-1}\subset D_{p}\subset D_{n}$, where $D_{p}\subset D_{n}$ is regular.
+Proof.
+Consider first the case when $\mathfrak{h}^{c}$ is irreducible.
+If $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ is a fundamental representation, Table 5 shows that (6) does hold (with exceptions 1)-3)).
+Weyl dimension formula shows that the same is valid for any irreducible representation.
+If $\mathfrak{h}^{c}$ is reducible, one uses Theorem 8 and obtains an inclusion $\mathfrak{h}^{c}\hookrightarrow\tilde{\mathfrak{h}}^{c}_{1}$, where $\tilde{\mathfrak{h}}^{c}_{1}$ denotes the simple summand of some maximal subalgebra $\tilde{\mathfrak{h}}^{c}$ (one can also consult Table 6).
+Since $\operatorname{rank}\tilde{\mathfrak{h}}^{c}_{1}<\operatorname{rank}\mathfrak{g}^{c}$, one can argue as follows.
+If $\mathfrak{h}^{c}\subset\tilde{\mathfrak{h}}^{c}$ where $\tilde{\mathfrak{h}}^{c}$ is parabolic, then $\mathfrak{h}^{c}$ is not regular in $\tilde{\mathfrak{h}}^{c}$ and one can argue by induction.
+If $\tilde{\mathfrak{h}}^{c}$ is symmetric, then the following two possibilities may occur: either $\mathfrak{h}^{c}$ is still regular in $\mathfrak{g}^{c}$, and we apply already proved part of Theorem 2, or it is one of the exceptions 1)-3).
+If $\mathfrak{h}^{c}\subset\tilde{\mathfrak{h}}^{c}$ is a tensor representation, then the difference between $\operatorname{rank}\mathfrak{h}^{c}$ and $\mathfrak{g}^{c}$ is even bigger.
+∎
 Proposition 4.
-The classification of possible standard Clifford-Klein forms of classical simple Lie algebras amounts to finding these as real forms of the following classes of triples $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c}$, where $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ are simple subalgebras (not necessarily proper) of the following lists of subalgebras:
-  1. 1.
-    the class of Levi subalgebras of parabolic subalgebras,
-  2. 2.
-    the class of all maximal irreducible simple subalgebras in classical $\mathfrak{g}^{c}$, or belong to Tables 1 and 5 of [14],
-  3. 3.
-    $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ is simple, reducible and is contained in one of the following inclusions:
-    $\mathfrak{h}^{c}\subset\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k)\subset\mathfrak{so}(n,\mathbb{C}),$
-    $\mathfrak{h}^{c}\subset\mathfrak{sp}(k,\mathbb{C})\oplus\mathfrak{sp}(n-k,\mathbb{C})\subset\mathfrak{sp}(n,\mathbb{C}),0<k<n,$
-    $\mathfrak{h}^{c}\subset\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{sl}(st,\mathbb{C}),\ s,t\geq 2,$
-    $\mathfrak{h}^{c}\subset\mathfrak{so}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{so}(st,\mathbb{C}),\ n=st,s,t\geq 3,s,t\neq 4,$
-    $\mathfrak{h}^{c}\subset\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{so}(t,\mathbb{C})\subset\mathfrak{so}(st,\mathbb{C}),\ s,t\geq 2,$
-    $\mathfrak{h}^{c}\subset\mathfrak{sp}(s,\mathbb{C})\otimes\mathfrak{sp}(t,\mathbb{C})\subset\mathfrak{sp}(st,\mathbb{C}),\ s\geq 2,t\geq 3,\ or\ s=2,t=4,$
-    (and the same inclusions may hold for $\mathfrak{l}^{c}$),
-  4. 4.
-    a class of simple reducible subalgebras such that they are contained in some simple irreducible subalgebra $\tilde{\mathfrak{h}}^{c}$ of $\mathfrak{g}^{c}$ (with $\tilde{\mathfrak{h}}^{c}$ necessarily satisfying assumptions of 1)).
-  5. 5.
-    exceptional simple Lie subalgebras with the following simplest representations $\mathfrak{g}_{2}\subset\mathfrak{so}(7,\mathbb{C}),$ $\mathfrak{f}_{4}\subset\mathfrak{so}(26,\mathbb{C}),$ $\mathfrak{e}_{6}\subset\mathfrak{sl}(27,\mathbb{C}),$ $\mathfrak{e}_{7}\subset\mathfrak{sp}(56,\mathbb{C}),$ $\mathfrak{e}_{8}\subset\mathfrak{so}(248,\mathbb{C}).$
-5.1.3 Non-generic case $\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1}$
-This case is in a sense "non-generic" and we consider it first. Note that it will be used in other considerations as well. We will essentially use the following.
-Theorem 12 ([21], Theorem 1.5).
-Let $G/H$ be a homogeneous space of reductive type. If $G/H$ has a compact Clifford-Klein form, then there does not exist a closed subgroup $G^{\prime}$ reductive in $G$ satisfying the conditions
-  • $\mathfrak{a}_{g^{\prime}}\subseteq\mathfrak{a}_{h}$,
-  • $d(G^{\prime})>d(H)$.
-Proposition 6.
-Let
-$\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1}$
+If $\mathfrak{h}^{c}\hookrightarrow\mathfrak{g}^{c}$ and $\mathfrak{l}^{c}\hookrightarrow\mathfrak{g}^{c}$ are embeddings of simple and non-regular $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$, then for any real forms $\mathfrak{h}\hookrightarrow\mathfrak{g}$ and $\mathfrak{l}\hookrightarrow\mathfrak{g}$ either the inequality $d(\mathfrak{g})>d(\mathfrak{h})+d(\mathfrak{l})$ holds (and hence the triple $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ cannot yield compact Clifford-Klein forms), or $\mathfrak{h}^{c}=B_{p-1}$ is embedded into $\mathfrak{g}$ as in cases 1)-3) in Proposition 3.
+However, none of the embeddings 1)-3) above can yield compact Clifford-Klein forms.
+Proof.
+The proof of the first part of this Proposition is a rather complicated case-by case calculation of non-compact dimensions and we postpone it until subsection 5.3.1.
+Here we settle the cases of embeddings 1)-3) and explain the method of computing $d(\mathfrak{h})$ and $d(\mathfrak{l})$ for the real forms of simple irreducible subalgebras which are not of types 1)-3).
+Cases 1)-3).
+First, consider the embedding
+$\mathfrak{h}^{c}\subset B_{n}\subset D_{n+1},$
 where $\mathfrak{h}^{c}$ is a regular simple subalgebra of $B_{n}$, and $B_{n}$ is a subalgebra of all fixed points of an outer automorphism of $D_{n+1}$.
 The pair $(\mathfrak{g},\mathfrak{h})$ cannot generate a standard compact Clifford-Klein form such that $\mathfrak{g}\neq\mathfrak{h}+\mathfrak{l}$ (for some $\mathfrak{l}\subset\mathfrak{g}$).
-Proof.
 Choose a Cartan subalgebra $\mathfrak{c}\subset D_{n+1}$ such that $\mathfrak{c}\cap\mathfrak{h}^{c}$ is a Cartan subalgebra of $\mathfrak{h}^{c}$.
@@
 Following [15], p.
-507, we write $H_{1},...,H_{n+1}$ for the simple co-roots and the root vectors corresponding to them as $X_{1},...,X_{n+1},$.
+507, we write $H_{1},...,H_{n+1}$ for the simple coroots and the root vectors corresponding to them as $X_{1},...,X_{n+1},$.
 The opposite root vectors are denoted by $Y_{1},...,Y_{n+1}$.
@@
 $X_{1},...,X_{n-1},X_{n}+X_{n+1},Y_{1},...,Y_{n-1},Y_{n}+Y_{n+1}.$
-Since any subalgebra $\mathfrak{h}^{c}$ which is regular in $B_{n}$ is obtained from the extended Dynkin diagram of $B_{n}$ by deleting some vertices, we can write all the possibilities for the simple roots and root vectors of $\mathfrak{h}^{c}$ in terms of the simple roots, or corresponding co-roots and root vectors expressed via the data for the type $D_{n+1}$.
+Since any subalgebra $\mathfrak{h}^{c}$ which is regular in $B_{n}$ is obtained from the extended Dynkin diagram of $B_{n}$ by deleting some vertices, we can write all the possibilities for the simple roots and root vectors of $\mathfrak{h}^{c}$ in terms of the simple roots, or corresponding coroots and root vectors expressed via the data for the type $D_{n+1}$.
 The extended Dynkin diagram for $B_{n}$ is given by vertices
 $\alpha_{0},\alpha_{1},...,\alpha_{n-1},\bar{\alpha}_{n}:=\frac{1}{2}(\alpha_{n}+\alpha_{n+1})$
-where $\alpha_{0}$ is the minimal root of $B_{n}$.
-Looking at the explicit expressions for maximal roots (for example, in [9]) one obtains the expression for each root of this diagram expressed via roots in of $D_{n+1}$:
 $\alpha_{0}=-(\alpha_{1}+2\alpha_{2}+\cdots+2\alpha_{n-1}+(\alpha_{n}+\alpha_{n+1})).$ (7)
@@
 Therefore, [35] yields the desired result.
-5.1.4 Completion of proof for $\mathfrak{h}^{c}$ and $\mathfrak{l}^{c}$ simple
-Proposition 7.
-If $\mathfrak{h}^{c}$ is a simple non-regular subalgebra in a simple Lie algebra $\mathfrak{g}^{c}$ of classical type, then
-$\operatorname{rank}\mathfrak{h}^{c}\leq\frac{1}{2}(\operatorname{rank}\mathfrak{g}^{c}+1)$
-with the following exceptions:
+Now we are ready to explain the method of calculating the non-compact dimensions for all embeddings of simple non-regular Lie subalgebras.
   1. 1.
-    $\mathfrak{h}^{c}:=B_{p-1}\subset D_{p}$
+    By Proposition 3 the inequality (6) holds.
   2. 2.
-[…]
-  3. 3.
-[…]
-Proof.
-Recall that by Proposition 5 the inclusions of a simple Lie algebra $\mathfrak{h}^{c}$ into classical $\mathfrak{g}^{c}$ fall into classes:
-  1. 1.
-    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$, where $\hat{\mathfrak{h}}^{c}$ is maximal parabolic,
-  2. 2.
-    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$, where
-    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}=\mathfrak{so}(k,\mathbb{C})\oplus\mathfrak{so}(n-k,\mathbb{C}),$
-    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}=\mathfrak{sp}(k,\mathbb{C}\oplus\mathfrak{sp}(n-k,\mathbb{C}).$
-  3. 3.
-    $\mathfrak{h}^{c}\subset\hat{\mathfrak{h}}^{c}$, where $\hat{\mathfrak{h}}^{c}$ is of tensor type,
-  4. 4.
-    $\mathfrak{h}^{c}$ is irreducible and simple.
-We begin the proof with case (4) which requires the longest calculations.
-Since $\hat{\mathfrak{h}}^{c}$ is simple and irreducible, we consider first the case when the inclusion $\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ is a fundamental representation.
-Looking at Tables 4,5,6 we calculate the possible ranks of the inclusions, using formulas from Table 4.
-In any case we obtain the required inequality.
-If the representation $\hat{\mathfrak{h}}^{c}\subset\mathfrak{g}^{c}$ is not fundamental, we use the Weyl dimension formula [18], Theorem 5.84.
-The proof of case $(2)$ can be obtained from the following observation.
-If $\mathfrak{h}\hookrightarrow\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$, then necessarily there is an embedding of $\mathfrak{h}$ to at least one of $\mathfrak{g}_{i}$ ($i=1,2$).
-Otherwise, the kernels of the projections onto $\mathfrak{g}_{i}$ composed with the embedding would yield non-trivial ideals in $\mathfrak{h}$.
-Thus, $\mathfrak{h}^{c}\subset\mathfrak{g}_{1}$ or $\mathfrak{h}^{c}\subset\mathfrak{g}_{2}$.
-In any case, one obtains an embedding of simple Lie algebras with $\mathfrak{g}_{i}$ of lower ranks.
-One proceeds by induction using $(4)$.
-This argument also settles case $(1)$.
-Since $\hat{\mathfrak{h}}^{c}$ is maximal parabolic, its reductive Levi subalgebra has the form
-$\mathfrak{f}(\mathfrak{c},[M]),\,M\subset\Pi,$
-where $M$ is a subset of $\Pi$, and $[M]$ is a subsystem of the root system $\Delta$.
-Thus, $\mathfrak{h}^{c}$ is a simple subalgebra contained in a semisimple subalgebra of a lower rank.
-In the same manner one settlles (3), since $\hat{\mathfrak{h}}^{c}$ is a direct sum of two simple Lie subalgebras of $\mathfrak{sl}(V_{1})\oplus\mathfrak{sl}(V_{2})$ embedded into $\mathfrak{sl}(V_{1})\otimes\mathfrak{sl}(V_{2})$.
-Thus, the same argument works.
-Corollary 6.
-If $\mathfrak{h}^{c}$ belongs to any of the inclusions in Proposition 7 then no real form of the pair $(\mathfrak{g}^{c},\mathfrak{h}^{c})$ cannot generate compact standard Clifford-Klein form.
-Proof.
-The proof follows from the proof of Proposition 6 with small modifications.
-One can notice that if $\mathfrak{g}^{c}=D_{n}$, or $\mathfrak{g}^{c}=B_{n}$ then the same argument works, this time applied to $\mathfrak{h}^{c}$ and the ambient Lie algebra $\mathfrak{g}^{c}$.
-Still, one obtains the embeddings
-$\mathfrak{so}(q,2p-1-q)\subset\mathfrak{so}(m,2n+1-m),\,\,\text{or}\,\,\mathfrak{so}(q,2p-1-q)\subset\mathfrak{so}(m,2n-m)$
-and the same argument using Theorem 12 completes the proof.
-The case $\mathfrak{g}^{c}=C_{n}$ need not be considered for the dimensional reasons: the smallest possible dimension of a symplectic representation of $D_{p}$ is $\frac{n}{2}$.
-∎
-Let us now list all simple real forms of non-compact type.
+    looking at Tables 4,5,3, 2 and using relations between real and complex ranks together with the constraint given by inequality (6), one obtains the desired inequality for the non-compact dimensions.
+As we have mentioned, the details of the calculation are postponed until Subsection 5.3.1.
+Clearly, Proposition 4 exhaust all the possibilities for non-regular triples $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c})$ of simple complex Lie algebras and their real forms, which completes the proof of Theorem 2.
+5.3.1 Proof of Proposition 4
+We need to list all simple real forms of non-compact type from [33], Table 4, in the form suitable for the calculations of the non-compact dimensions.
 By a slight abuse of notation we will treat simple but not absolutely simple real Lie algebras of rank 2n as Lie algebras of type $n.$ For instance $\mathfrak{g}=\mathfrak{sl}(n+1,\mathbb{C})$ is classified as type $A_{n}.$ Notice that in this case $\mathfrak{g}^{c}=\mathfrak{sl}(n+1,\mathbb{C})\oplus\mathfrak{sl}(n+1,\mathbb{C}).$
@@
   • $\mathfrak{e}_{8}(\mathbb{C}),$ $d=248,$ $r=8,$
-Remarks:
+Now we make some extra analysis of this list.
+Observations based on the list above
   1. 1.
-    For all classical types with $\textrm{rank}=t$ and $\textrm{rank}_{\mathbb{R}}=a$ one of the following three algebras has the highest $d$
+    For all classical types with $\textrm{rank}=t$ and $\textrm{rank}_{\mathbb{R}}=a$ one of the following three algebras has the highest number $d:=d(\mathfrak{g})$
     • I$:=\mathfrak{so}(a,2t+1-a),$ $d=2at+a-a^{2},$
@@
   3. 3.
-    For $\mathfrak{g}_{2}(\mathbb{C})$ one can find a real form of type I ($\mathfrak{so}(2,7)$) with the same real rank, with greater or equal $d$ and with rank at most half of the rank of the representation into algebra of rank 8 or greater.
-As the simplest representation of $\mathfrak{g}_{2}(\mathbb{C})$ is $B_{6}.$ Thus in each case one has to verify separately the embeddings of $\mathfrak{g}_{2}(\mathbb{C})$ into lie algebras of rank $6$ and $7$
+    For $\mathfrak{g}_{2}(\mathbb{C})$ one can find a real form of type I ($\mathfrak{so}(2,7)$) with the same real rank, and $d$ the same or greater, and with rank at most half of the rank of the representation into an algebra of rank $8$ or greater.
+As the simplest representation of $\mathfrak{g}_{2}(\mathbb{C})$ is $B_{6},$ in each case one has to verify separately the embeddings of $\mathfrak{g}_{2}(\mathbb{C})$ into some Lie algebras of rank $6$ and $7$
   4. 4.
-    For $\mathfrak{g}_{2(2)}$ one can find a real form of type I ($\mathfrak{so}(2,5)$) with the same real rank, with greater or equal $d$ and with rank at most half of the rank of the representation into algebra of rank 6 or greater.
-As the simplest representation of $\mathfrak{g}_{2(2)}$ is $B_{3}$ then one should in each case verify the embeddings of $\mathfrak{g}_{2(2)}$ into lie algebras of rank $4$ and $5$ This is because the real rank of $\mathfrak{g}_{2(2)}$ equals 2 and so if $(\mathfrak{g},\mathfrak{g}_{2(2)},\mathfrak{l})$ gives a standard compact Clifford-Klein form then necessarily the real rank of $\mathfrak{g}$ is at lest 3.
-If the $\textrm{rank}(\mathfrak{g})=3$ then $\mathfrak{g}$ is split and we can use Theorem “split”.
+    For $\mathfrak{g}_{2(2)}$ one can find a real form of type I ($\mathfrak{so}(2,5)$) with the same real rank, with $d$ the same or greater, and with rank at most half of the rank of the representation into a Lie algebra of rank 6 or greater.
+As the simplest representation of $\mathfrak{g}_{2(2)}$ is $B_{3}$, one should verify the embeddings of $\mathfrak{g}_{2(2)}$ into Lie algebras of rank $4$ and $5$ This is because the real rank of $\mathfrak{g}_{2(2)}$ equals 2 and so if $(\mathfrak{g},\mathfrak{g}_{2(2)},\mathfrak{l})$ gives a standard compact Clifford-Klein form then necessarily the real rank of $\mathfrak{g}$ is at least 3.
+If the $\textrm{rank}(\mathfrak{g})=3$ then $\mathfrak{g}$ is split and we can refer to Corollary 1.
 We only have to check the triples $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ such that $\mathfrak{h}$ and $\mathfrak{l}$ are one of the following
@@
 and $\textrm{rank}_{\mathbb{R}}(\mathfrak{h})=:a>0,$ $\textrm{rank}_{\mathbb{R}}(\mathfrak{l})=:b>0,$ $\textrm{rank}_{\mathbb{R}}(\mathfrak{g})=a+b.$
-Proposition 8.
-For any classical absolutely simple $\mathfrak{g}$ in every case described above
-$D:=d(\mathfrak{g})-d(\mathfrak{h})-d(\mathfrak{l})>0$
-and so one cannot obtain compact standard Clifford-Klein forms.
-Proof.
-By the remarks in the previous Section we only have to check the inequality for types I,II,III (and $G_{2}$ for low ranks).
+Final check up of the inequality $d(\mathfrak{h})+d(\mathfrak{l})<d(\mathfrak{g})$
+By the observations above we only have to check the inequality for types I,II,III (and $G_{2}$ for low ranks).
 By $X_{t}$ we will denote an arbitrary type of rank $t.$ By $(i,j),$ $i,j\in\{I,II,III\}$ we denote a pair of subalgebras of type i and j, respectively.
-We exclude the cases in which $\mathfrak{g}$ is split, becuse of [6].
-Type: $A_{2n}$ Case $\mathfrak{g}=\mathfrak{su}(a+b,2n+1-a-b)$ We start with $\mathfrak{h}$ of type $G_{2}.$ First we have to check $\mathfrak{g}_{2(2)}$ in $A_{4}.$ But in this case real rank of $\mathfrak{g}$ is at most 2.
+We exclude the cases in which $\mathfrak{g}$ is split, because of [8].
+Type $A_{2n}$ Case $\mathfrak{g}=\mathfrak{su}(a+b,2n+1-a-b)$ We begin with $\mathfrak{h}$ of type $G_{2}.$ First we have to check $\mathfrak{g}_{2(2)}$ in $A_{4}.$ But in this case the real rank of $\mathfrak{g}$ is at most 2.
 Secondly we have $\mathfrak{g}_{2}(\mathbb{C})$ in $A_{6}.$ So necessarily $\mathfrak{g}=\mathfrak{su}(3,4).$ One easily checks that there is no simple $\mathfrak{l}$ of real rank one with $d(\mathfrak{l})>8.$
 Clearly, the only proper regular subalgebras in $A_{2n}$ are of type $A_{p},$ $p<2n.$ Note that any non-regular simple Lie subalgebra of $A_{2n}$ is of the rank at most $n,$ by Proposition 4.
-If $\mathfrak{h}$ is simple but not absolutely simple then it is obviously of type $X_{p}$ with $p\leq n.$ Thus we have to check Lie subalgebras of type I,II and III for $t=n.$ We have
+If $\mathfrak{h}$ is simple but not absolutely simple then it is obviously of type $X_{p}$ with $p\leq n.$ Thus we have to check the Lie subalgebras of type I,II and III for $t=n.$ We have
   • $d(\mathfrak{g})=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab,$ $r=a+b,$
@@
 (III,III) $D=4an+4bn+2a+2b-2a^{2}-2b^{2}-4ab-2a^{2}-a-2b^{2}-b\geq 4ab+a+b>0.$
-Type: $A_{2n+1}$
-Again, by Proposition 7 we see that if $X_{p}\subset A_{2n+1}$ is classical simple non-regular subalgebra or simple but not absolutely simple subalgebra then $p\leq n+1.$
+Type $A_{2n+1}$
+Again, by Proposition 4 we see that if $X_{p}\subset A_{2n+1}$ is a classical simple non-regular subalgebra or simple but not absolutely simple subalgebra then $p\leq n+1.$
 We begin with subalgebras of type $G_{2}.$
@@
 (III’,III’) $D=4an+4bn-2a^{2}-2b^{2}-4ab+4a+4b-a^{2}-2a-b^{2}-2b\geq(a+b)(a+b-2)+2ab>0.$
-Type: $C_{n}$
-In this type we only have to check $\mathfrak{g}=\mathfrak{sp}(a+b,n-a-b).$
-We begin with subalgebras of type $G_{2}.$ We can eliminate (in the case $\mathfrak{h}=\mathfrak{g}_{2(2)}$) $C_{4}$ and $C_{5}$ since the real of $\mathfrak{sp}(2,2)$ and $\mathfrak{sp}(2,3)$ equals $2$ If $\mathfrak{h}:=\mathfrak{g}_{2}(\mathbb{C})\subset A_{6}$ then (since the real rank of $\mathfrak{g}$ has to be greater than 2 and we omit the split case) necessarily $\mathfrak{g}=\mathfrak{sp}(3,3).$ But now $\mathfrak{l}\subset\mathfrak{sp}(3,3)$ has to have real rank equal to 1 and $d=22.$ One easily verifies that this is impossible.
+Type $C_{n}$
+In this case we only have to check $\mathfrak{g}=\mathfrak{sp}(a+b,n-a-b).$
+We begin with subalgebras of type $G_{2}.$ We can eliminate (in the case $\mathfrak{h}=\mathfrak{g}_{2(2)}$) $C_{4}$ and $C_{5}$ since the real rank of $\mathfrak{sp}(2,2)$ and $\mathfrak{sp}(2,3)$ equals $2$ If $\mathfrak{h}:=\mathfrak{g}_{2}(\mathbb{C})\subset A_{6}$ then (since the real rank of $\mathfrak{g}$ has to be greater than 2 and we omit the split case) necessarily $\mathfrak{g}=\mathfrak{sp}(3,3).$ But now $\mathfrak{l}\subset\mathfrak{sp}(3,3)$ has to have real rank equal to 1 and $d=22.$ One easily verifies that this is impossible.
 We can use the same argument for $C_{7}$ (in this case necessarily $\mathfrak{g}=\mathfrak{sp}(3,4)$).
-Regular proper subalgebras of $C_{n}$ are given by $A_{p},C_{p},D_{s},$ $p+1,s\leq n.$ By Proposition 7 to obtain non-regular subalgebra in $C_{n}$ of rank greater than $\frac{n}{2}$ one has to “use” the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $C_{n}.$ By Corollary 6 any non-regular simple subalgebra in $D_{p}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
+Regular proper subalgebras of $C_{n}$ are given by $A_{p},C_{p},D_{s},$ $p+1,s\leq n.$ By Proposition 4 to obtain non-regular subalgebra in $C_{n}$ of rank greater than $\frac{n}{2}$ one has to “use” the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $C_{n}.$ By Proposition 4 any non-regular simple subalgebra in $D_{p}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
 $A_{\frac{n}{2}},B_{n-1},C_{{\frac{n}{2}}},D_{{\frac{n}{2}}}\subset C_{n},$
@@
 (III,III) $D=4an+4bn-4a^{2}-4b^{2}-8ab-2a^{2}-a-2b^{2}-b\geq 2(a^{2}+b^{2})+a(4b-1)+b(4a-1)>0.$
-Type: $B_{n}$
+Type $B_{n}$
 In this type we only have to check $\mathfrak{g}=\mathfrak{so}(a+b,2n+1-a-b).$
-We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then $\mathfrak{g}=\mathfrak{so}(3,6),$ $\mathfrak{so}(4,5).$ If $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then $\mathfrak{g}=\mathfrak{so}(s,13-s),\mathfrak{so}(s,15-s),$ $s=3,4,5,6,7.$ One verifies that there does not exist $\mathfrak{l}$ with adequate $r,d$ with one exception:
+We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then $\mathfrak{g}=\mathfrak{so}(3,6),$ $\mathfrak{so}(4,5).$ If $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then $\mathfrak{g}=\mathfrak{so}(s,13-s),\mathfrak{so}(s,15-s),$ $s=3,4,5,6,7.$ One verifies that there does not exist $\mathfrak{l}$ with appropriate $r,d$, with one exception:
 $(\mathfrak{so}(6,7),\mathfrak{so}(4,7),\mathfrak{g}_{2}(\mathbb{C})),$
 but in this case $\mathfrak{g}$ is split.
-Regular proper subalgebras of $B_{n}$ are given by $A_{p},B_{p},D_{s},$ $p+1,s\leq n.$ By Corollary 6 to obtain non-regular subalgebra in $B_{n}$ of rank greater than $\frac{n}{2}$ one has to consider the chain
+Regular proper subalgebras of $B_{n}$ are given by $A_{p},B_{p},D_{s},$ $p+1,s\leq n.$ By Proposition 4 to obtain non-regular subalgebra in $B_{n}$ of rank greater than $\frac{n}{2}$ one has to consider the chain
 $\mathfrak{h}^{c}\subset B_{p-1}\subset D_{p}\subset B_{n}$
@@
 (III,III) $D=2an+2bn-a^{2}-b^{2}-2ab+a+b-2a^{2}-a-2b^{2}-b\geq(a-b)^{2}.$ Thus $D>0$ unless $a=b=\frac{n}{2}.$ But then $\mathfrak{g}=\mathfrak{so}(n,n+1),$ which is split.
-Type: $D_{n}$
-Notice that if $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then we have to check $D_{6}$ and $D_{7}.$ But if $\mathfrak{g}_{2}(\mathbb{C})$ is contained in Lie algebra of type $D_{p}$ then necessarily $p\geq 8.$
-Regular proper subalgebras of $D_{n}$ are given by $A_{p},D_{p},$ $p\leq n.$ Corollary 6 eliminates the possibility $\operatorname{rank}\mathfrak{h}^{c}>\frac{1}{2}\operatorname{rank}\mathfrak{g}^{c}$.
+Type $D_{n}$
+Notice that if $\mathfrak{h}=\mathfrak{g}_{2}(\mathbb{C})$ then we have to check $D_{6}$ and $D_{7}.$ But if $\mathfrak{g}_{2}(\mathbb{C})$ is contained in the Lie algebra of type $D_{p}$ then necessarily $p\geq 8.$
+Regular proper subalgebras of $D_{n}$ are given by $A_{p},D_{p},$ $p\leq n.$ Proposition 4 eliminates the possibility $\operatorname{rank}\mathfrak{h}^{c}>\frac{1}{2}\operatorname{rank}\mathfrak{g}^{c}$.
 The opposite inequality is analyzed below.
@@
 We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then we have to check $\mathfrak{g}=\mathfrak{so}^{\ast}(4),\mathfrak{so}^{\ast}(5).$ In both cases real rank equals 2 (which equals the real rank of $\mathfrak{g}_{2(2)}$).
-By Corollary 6 any non-regular simple subalgebra in $B_{n}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
+By Proposition 4 any non-regular simple subalgebra in $B_{n}$ of rank greater or equal to $\frac{p}{2}$ is of the type $B.$ Thus we have to check the following embeddings:
 $A_{\frac{n}{2}},B_{n-1},C_{{\frac{n}{2}}},D_{{\frac{n}{2}}}\subset D_{n},$
@@
 We see that one only needs to check I and III.
-Also $n\geq 4,$ $\frac{n-1}{2}\leq a+b\leq\frac{n}{2}$ and so $a^{2}+b^{2}\leq\frac{n^{2}}{4}-2ab.$ (I,I) $D=n^{2}-n-2an+a+a^{2}-2bn+b+b^{2}\geq-n+a^{2}+a+b^{2}+b>0,$ unless $a=b=1$ and $n=4.$ But in this case we have embedding $B_{3}\subset D_{4},$ which is symmetric and we can conclude by [32].
+Also $n\geq 4,$ $\frac{n-1}{2}\leq a+b\leq\frac{n}{2}$ and so $a^{2}+b^{2}\leq\frac{n^{2}}{4}-2ab.$ (I,I) $D=n^{2}-n-2an+a+a^{2}-2bn+b+b^{2}\geq-n+a^{2}+a+b^{2}+b>0,$ unless $a=b=1$ and $n=4.$ But in this case we have embedding $B_{3}\subset D_{4},$ which is symmetric and we can apply [35].
 (I,III) $D=n^{2}-n-2an+a+a^{2}-2b^{2}-b\geq 2bn-n+a+a^{2}-2b^{2}-b\geq 2b^{2}-3b+4ab+a^{2}-a-1=b(2b-3)+2a+a(a-1)+2ab-1>0,$
@@
 Case $\mathfrak{g}=\mathfrak{so}(a+b,2n-a-b)$
-We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then we have to check $\mathfrak{g}=\mathfrak{so}(3,5),$ $\mathfrak{so}(3,7),$ $\mathfrak{so}(4,6).$ One verifies that there does not exist $\mathfrak{l}$ with adequate $r,d.$
-We have the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $D_{n}.$ By Corollary 6 we can eliminate this case.
+We begin with subalgebras of type $G_{2}.$ If $\mathfrak{h}=\mathfrak{g}_{2(2)}$ then we have to check $\mathfrak{g}=\mathfrak{so}(3,5),$ $\mathfrak{so}(3,7),$ $\mathfrak{so}(4,6).$ One verifies that there does not exist $\mathfrak{l}$ with appropriate $r,d.$
+We have the non-regular embedding $B_{p-1}\subset D_{p},$ where $D_{p}$ is regular in $D_{n}.$ By Proposition 4 we can eliminate this case.
 Thus we may assume that each subalgebra of $\mathfrak{so}(a+b,2n-a-b)$ is of rank at most $\frac{n}{2}.$
@@
 For any combination (that is (III.1,III.1), (III.1,III.2), (III.2,III.2)) we have $D\geq(a-b)^{2}+a+b>0.$
-Proposition 9.
-No real form of the triple $(\mathfrak{g}^{c},\mathfrak{h}^{c},\mathfrak{l}^{c})$ with at least one of the pair $(\mathfrak{h}^{c},\mathfrak{l}^{c})$ satisfies the assumptions 1) or 3) of Proposition 5 can yield a standard compact Clifford-Klein form.
-Proof.
-Consider first $\mathfrak{h}^{c}$ or $\mathfrak{l}^{c}$ which are simple, reducible and contained in direct sums of simple classical Lie algebras $\mathfrak{g}^{c}=\mathfrak{g}_{1}^{c}\oplus\mathfrak{g}^{c}_{2}$.
-One can notice that such subalgebra is necessarily isomorphic to a subalgebra in one of the summands, $\mathfrak{g}_{1}^{c}$ or $\mathfrak{g}_{2}^{c}$.
-But since $\dim\mathfrak{g}_{1}^{c}<\dim\mathfrak{g}^{c}$ and $\dim\mathfrak{g}^{c}_{2}<\dim\mathfrak{g}^{c}$, one can apply the previous considerations to the inclusion of reducible $\tilde{\mathfrak{h}}^{c}$ or $\tilde{\mathfrak{l}}^{c}$ (isomorphic to $\mathfrak{h}^{c}$ or to $\mathfrak{l}^{c}$) into one of the $\mathfrak{g}_{1}^{c}$ or $\mathfrak{g}_{2}^{c}$.
-Since the non-compact dimension is the invariant of a Lie algebra itself we get the same inequality for non-compact dimensions of $\mathfrak{g},\mathfrak{h},\mathfrak{l}$ ensuring that these triples cannot generate compact standard Clifford-Klein forms.
-5.1.5 Proof of Theorem 2 for the semisimple case
-Assume that $(\mathfrak{g},\mathfrak{h},\mathfrak{l})$ corresponds to a standard triple and let $\mathfrak{h}=\mathfrak{h}_{1}\oplus\mathfrak{h}_{2}.$ We may also assume that $\mathfrak{g}$ is of classical type and non-split.
-Let
-$\mathfrak{h}_{i}=\mathfrak{k}_{i}\oplus\mathfrak{a}_{i}\oplus\mathfrak{n}_{i},\ \ i=1,2$
-be the compatible Iwasawa decompositions.
-We see that $\mathfrak{h}_{1}$ is in the centralizer of $\mathfrak{a}_{2}$ in $\mathfrak{g}.$ Similarly, $\mathfrak{h}_{2}$ is in the centralizer of $\mathfrak{a}_{1}.$ Therefore, by Theorem 4 all the roots of $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ are contained on two planes: one orthogonal to $\mathfrak{a}_{1},$ the second orthogonal to $\mathfrak{a}_{2}.$ But now, by Lemma 5 for $\Sigma,$ there exists $\beta\in\Delta_{0}^{\pm}$ such that the restriction of $\beta$ to $\mathfrak{a}_{i},$ $i=1,2,$ is non-zero.
-By Lemma 7 every root of $\Delta_{0}^{\pm}$ is a sum of some two roots in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}.$ Thus $\beta=\alpha_{1}+\sigma(\alpha_{2}),$ so that $\alpha_{1}|_{\mathfrak{a}_{2}}=0$ and $\sigma(\alpha_{2})|_{\mathfrak{a}_{1}}=0.$ It follows that $\alpha_{1}$ and $\sigma(\alpha_{2})$ are orthogonal and so the root subsystem generated by those roots is necessarily of type $B_{2}.$ Thus $\mathfrak{g}$ has to by of type $B_{t}$ or $C_{t}.$ We thus have only two possibilities (we may assume that $\alpha_{1}(iX)>0$).
-Case 1.
-$\mathfrak{g}=\mathfrak{so}(n,m),$ $n+m$ is odd.
-In this case for any root $\alpha$ in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ the root $\sigma(\alpha)$ has the same projection on $\mathfrak{a}.$ But in $\mathfrak{so}(n,m)$ any long root of $\Delta$ can correspond only to a 1-dimensional root space of $\mathfrak{g}.$ It follows that $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ consists only of short roots.
-But now $\alpha_{1}+\alpha_{2}$ is a root because a sum of any two (non-colinear) short roots in $B_{t}$ is a root.
-By Theorem 4 we see that
-$S_{h}^{1}=x_{\alpha_{1}}+\sum_{l=1}^{k}a^{i}_{k+l}x_{\alpha_{k+l}}+Q_{1},$
-$S_{h}^{2}=x_{\alpha_{2}}+\sum_{l=1}^{k}a^{i}_{k+l}x_{\alpha_{k+l}}+Q_{2},$
-belong to $\mathfrak{h}^{c}.$ We also see that $[x_{\alpha_{1}},x_{\alpha_{2}}]$ is non-zero and it is a unique factor in $[S_{h}^{1},S_{h}^{2}]$ with maximal value on $ad_{iX}.$ Thus $0\neq Y:=[S_{h}^{1},S_{h}^{2}]\in\mathfrak{h}^{c}$ and
-$ad_{\mathfrak{a}_{i}}(Y)\neq 0,\ \ i=1,2.$
-But this shows that $\mathfrak{h}\neq\mathfrak{h}_{1}\oplus\mathfrak{h}_{2},$ a contradiction.
-Case 2 One can deal with case 2 only by using general structure theory of Lie algebras.
-Indeed, let $\mathfrak{g}=\mathfrak{sp}(n,m),$ where $n\leq m.$ Since $\mathfrak{g}^{c}$ is of type $C_{p}$ therefore we may represent $\Delta$ as
-$\Delta=\{\pm e_{i}\pm e_{j},\pm 2e_{i}\ |\ i\neq j,1\leq i,j\leq p\},$
-where $(e_{i})_{i=1}^{p}$ is a standard basis of $\mathbb{R}^{p}.$ From the general consideration in Section 5.1.5 we know that there must exist roots $\alpha_{1}$ and $\sigma(\alpha_{2})$ in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ such that
-  1. 1.
-    $\alpha_{1}$ and $\sigma(\alpha_{2})$ are orthogonal,
-  2. 2.
-    $\alpha_{1}(iX)=-\sigma(\alpha_{2})(iX)\neq 0,$
-  3. 3.
-    $\beta=\alpha_{1}+\sigma(\alpha_{2})$ is a (long) root of $\Delta$ contained in $\Delta_{0}^{\pm},$
-  4. 4.
-    $\alpha_{1}|_{\mathfrak{a}_{2}}=0$ and $\sigma(\alpha_{2})|_{\mathfrak{a}_{1}}=0.$
-Therefore necessarily
-$\alpha_{1}=e_{i}+e_{j},\ \ \ \sigma(\alpha_{2})=e_{i}-e_{j},$
-for some $i\neq j$ and $e_{i}(iX)=0,$ $e_{j}(iX)>0.$ Consider the root
-$\tilde{\beta}:=\alpha_{1}-\sigma(\alpha_{2})=2e_{j}.$
-We see that
-$\tilde{\beta}(iX)\neq 0,\ \ \tilde{\beta}|_{\mathfrak{a}_{2}}=\sigma(\alpha_{2}),\ \ \tilde{\beta}|_{\mathfrak{a}_{1}}=\alpha_{1}.$
-Thus $\tilde{\beta}$ is in $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ and is non-zero on $\mathfrak{a}_{1}$ and on $\mathfrak{a}_{2}.$ But, as is shown in Section 5.1.5, the roots of $\Delta_{p}^{\pm}\cup\Delta_{n}^{\pm}$ are contained in two planes: one orthogonal to $\mathfrak{a}_{1},$ the second orthogonal to $\mathfrak{a}_{2}.$ A contradiction.
-$\square$.
-Remark 4.
-The case of $\mathfrak{g}=\mathfrak{sp}(n,m),$ $n\leq m.$ can be settled also by case-by-case analysis as in Section 5.1.4.
-One shows that the only possibility is that $\mathfrak{h},\mathfrak{l}$ are of the form
-$\mathfrak{h}=\oplus_{s=1}^{w_{h}}\mathfrak{sp}(n_{s},m_{s}),\ \ \ \mathfrak{l}=\oplus_{s=1}^{w_{l}}\mathfrak{sp}(n_{s},m_{s}).$
-By [21] (Theorem 1.5), for any $s$ we have $m_{s}=m$ (otherwise one could construct a reductive subalgebra $\mathfrak{g}^{\prime}$ in $\mathfrak{g}$ with $d(\mathfrak{g}^{\prime})>d(\mathfrak{g})$ and $\mathfrak{a}_{g^{\prime}}=\mathfrak{a}_{h}$, which is impossible.
-But now we see that $\mathfrak{h}$ and $\mathfrak{l}$ have to be simple.
-A contradiction.
-6 Addendum: tables and explicit calculations
-6.1 Tables
+6 Tables
+Non-compact dimensions
+[…]
+$A_{t}$ $2a(t+1)-a$ $1\leq a\leq\frac{t}{2}$
+$A_{t}$ $2a^{2}$ $t=2a-1$ $a\geq 2$
+$B_{t}$ $a(2t+1-a)$ $1\leq a\leq t$
+$C_{t}$ $4a(t-a)$ $1\leq a\leq\frac{t}{2}$
+$D_{t}$ $a(2t-a)$ $1\leq a\leq t$
+$D_{t}$ $2a(2a-1)$ $t=2a$
+$D_{t}$ $2a(2a+1)$ $t=2a+1$
+Table 4:
 Dimensions of fundamental representations
@@
 Table 5:
-Non-maximal irreducible subalgebras
-Type of inclusion dimension $\omega_{r}$
-$A_{n}\subset A_{l}$ $3\binom{n+2}{4}$ 00
-$A_{n}\subset A_{l}$ $3\binom{n+3}{4}$ 00
-$B_{2n+1}\subset\mathfrak{g}^{c}$ $\prod_{s=1}^{2n}{\binom{k+s-1}{k}/\binom{k+s}{k}}$ $(-1)^{n+1)k}$
+Maximal subalgebras of maximal rank in $\mathfrak{g}^{c}$
+$\mathfrak{g}^{c}$ $\mathfrak{\tilde{h}}^{c}$ $\operatorname{rank}\mathfrak{\tilde{h}}^{c}$
+$\mathfrak{so}(2l+1,\mathbb{C})$ $\mathfrak{so}(2k,\mathbb{C})\oplus\mathfrak{so}(2(l-k)+1,\mathbb{C})$ $l$
+$\mathfrak{sp}(2l,\mathbb{C})$ $\mathfrak{sp}(2k,\mathbb{C}\oplus\mathfrak{sp}(2(l-k)+1,\mathbb{C})$ $l$
+$\mathfrak{so}(2l,\mathbb{C})$ $\mathfrak{so}(2k,\mathbb{C})\oplus\mathfrak{so}(2(l-k)+1),\mathbb{C})$ $l$
+$\mathfrak{sl}(n,\mathbb{C})$ $\mathfrak{so}(2,\mathbb{C})\oplus\mathfrak{so}(2l-1,\mathbb{C})$ $l-1$
+$\mathfrak{so}(2l+2,\mathbb{C})$ $\mathfrak{so}(2,\mathbb{C})\oplus\mathfrak{so}(2l-1),\mathbb{C})$ $l-1$
+$\mathfrak{sp}(2l,\mathbb{C})$ $\mathfrak{gl}(l,\mathbb{C})$ $l-1$
+$\mathfrak{so}(2l,\mathbb{C})$ $\mathfrak{so}(2,\mathbb{C})\oplus\mathfrak{so}(2l-2,\mathbb{C})$ $l-1$
+$\mathfrak{so}(2l,\mathbb{C})$ $\mathfrak{gl}(l,\mathbb{C})$ $l-1$
 Table 6:
-Irreducible triples
-Inclusion 1 Inclusion 2 $\dim$ of the representation
-$C_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n-1+k}{k}$
-$B_{n}\subset A_{2n}$ $A_{2n}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n+1}{k}$
-$D_{n}\subset A_{2n-1}$ $A_{2n-1}\subset\tilde{\mathfrak{g}}^{c}$ $\binom{2n}{k}$
-$A_{n}\subset A_{\frac{(n-1)(n+2)}{2}}$ $A_{\frac{(n-1)(n+2)}{2}}\subset\tilde{\mathfrak{g}}^{c}$ $\frac{1}{8}(n-1)(n+1)(n+2)$
-$A_{n}\subset A_{\frac{n(n+3)}{2}}$ $A_{\frac{n(n+3)}{2}}\subset\tilde{\mathfrak{g}}^{c}$ $\frac{1}{8}(n+1)(n+2)(n+3)$
-$B_{n}\subset D_{n+1}$ $D_{n+1}\subset\tilde{\mathfrak{g}}^{c}$ $\prod_{s=1}^{n-1}\binom{k+2s-1}{k}/\binom{k+s}{k}$
-Table 4:
-6.2 Explanations to Tables
-In Table 2 $\bar{\omega}_{r}$ is 00, if the representation is linear, but neither symplectic, nor orthogonal, it is $1$ is the representation is orthogonal, and it is $(-1)$ if the representation is symplectic.
-Table 3 reproduces inclusions $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}$ of complex simple Lie algebras when $\mathfrak{h}^{c}$ is not maximal in $\mathfrak{g}^{c}$ from [14], Table 1 (recall that we reproduce only the infinite series).
-Table 4 reproduces inclusions of triples of irreducible inclusions $\mathfrak{h}^{c}\subset\mathfrak{g}^{c}\subset\tilde{\mathfrak{g}}^{c}$ together with the dimensions of $\tilde{\mathfrak{g}}^{c}$.
-Here we need to explain some simplification in Tables 4 and 5.
-In [14] the author is concerned with the classification of subalgebras up to conjugation.
-Then non-conjugate subalgebras differ by integer labels $k$ in the corresponding Dynkin diagrams.
-These numbers yield different values of $\dim\tilde{\mathfrak{g}}^{c}$.
-However, our argument does not use this more rained classification, since we consider all possible real ranks for all possible real forms.
-Therefore, we do not indicate different (non-conjugate) inclusions of the same subalgebra.
-6.3 An estimate
-Lemma 12.
-The following inequality holds for $l\geq 3$
-$\binom{2l}{r}-\binom{2l}{r-2}>l^{2},2\leq r\leq l.$ (4)
-Proof.
-For $r=2$ one obtains
-$\binom{2l}{2}-1=\frac{(2l-1)2l}{2}-1=2l^{2}-l-1>l^{2},\,\text{for}\,l>1.$
-For $r=3$ one obtains
-$\binom{2l}{3}-\binom{2l}{1}=\frac{1}{3}l(4l^{2}-2l-\frac{4}{3}),\,\text{for}\,l\geq 3.$
-For $r=4$ one obtains
-$\binom{2l}{4}-\binom{2l}{2}=\frac{1}{12}l(2l-1)(4l^{2}-10l+6),\,\text{for}\,l\geq 4.$
-For $r\geq 5$ we use the formula
-$\binom{n}{k}=\prod_{i=1}^{k}\frac{n+1-i}{i}.$
-In our case
-$\binom{2l}{r}-\binom{2l}{r-2}=\prod_{i=1}^{r}\frac{2l+1-i}{i}-\prod_{i=1}^{r-2}\frac{2l+1-i}{i}=\binom{2l}{r-2}\cdot(\frac{(2l+2-r)(2l+1-r)}{(r-1)r}-1)=$
-$=\binom{2l}{r-2}\cdot\frac{4l^{2}-4lr+6l-2r+2}{r(r-1)}\geq\binom{2l}{3}\cdot\frac{4l^{2}-4lr+6l-2r+2}{r(r-1)}=(*)$
-Notice that (*)(*)( * ) obtains the minimal value for $r=l$ and so
-(*$(*)\geq\binom{2l}{3}\cdot\frac{4l+2}{l(l-1)}>2l^{2}-3l+1>l^{2},\,\text{for}\,l>4.$